### v2

# Entanglement measures for detectability

Abstract
We propose new entanglement measures as the detection performance based on the hypothesis testing setting. We clarify how our measures work for detecting an entangled state by extending the quantum Sanov theorem. Our analysis covers the finite-length setting. Exploiting this entanglement measure, we present how to derive entanglement witness to detect the given entangled state by using the geometrical structure of this measure. We derive their calculation formulas for maximally correlated states, and propose their algorithms that work for general entangled states. In addition, we investigate how our algorithm works for solving the membership problem for separability.
I Introduction
Entanglement is a valuable resource for quantum information processing. It is important to distinguish a given entangled state from separable states, i.e., non-entangled states. However, detecting an entanglement state has two meanings, the physical detection and the mathematical detection. The physical detection means the experimental detection of a generated entangled state. The mathematical detection means identifying whether a given density matrix has a separable form or not. Since the mathematical detection is often called the membership problem of separability, in this paper, we consider that entanglement detection means the physical one while this paper address both problems.
Several practical methods to achieve entanglement detection have been proposed [1, 2, 3, 4, 5, 6, 7], but they did not discuss the optimal performance for the detection. On the other hand, to characterize the amount of entanglement, many entangled measures have been proposed. However, these proposed entangled measures have less relation with the experimental detectability of entanglement. In fact, when Vedral et al [8, 9] introduced entanglement of relative entropy, they pointed out the relation between the hypothesis testing and the relative entropy. They employed the minimum relative entropy between the given entangled state and a separable state. However, this measure expresses the performance of the discrimination between the given entangled state and the closest separable state in the sense of relative entropy. When one hypothesis is chosen to be the set of all separable states, this measure does not express the performance of this discrimination [10]. Until now, no study discussed the optimal detection efficiency for entanglement verification.
In this paper, to resolve this problem, we propose a new entanglement measure that expresses the optimal performance of the discrimination between the given entangled state and the set of separable states. In this characterization, in the first step, we employ the quantum Sanov theorem [11, 12] for hypothesis testing, and consider the same minimization as entanglement of relative entropy by exchanging the first and second inputs in the relative entropy. The proposed quantity satisfies the entanglement monotone. To enhance this characterization, we derive a much simpler upper bound of the error probability in the finite-length setting of this kind of hypothesis testing, which can be considered another proof of the quantum Sanov theorem. In addition, we derive the strong converse exponent of the same problem setting as the quantum Sanov theorem. However, when we simply apply the quantum Sanov theorem, we assume that the separable states are limited to the iid states of separable states. To remove this assumption, in the second step, we derive the optimal performance of this discrimination when possible separable states are allowed to take a form of a general separable state.
This entanglement measure is useful for finding an entanglement witness to detect the target entangled state. This relation is characterized by information geometry [13]. Also. we calculate the proposed entanglement measures when a state is a maximally correlated state. This class contains all pure states. For this calculation, we introduce a notable relation with coherence measures, and several calculation formulas for coherence measures.
In the next step, we propose an efficient algorithm to compute the proposed entanglement measure on the joint system ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. It is quite difficult to directly calculate entanglement measures because the calculations of entanglement measures are directly related to the membership problem of separability, which is known as NP-hard [14, 15, 16, 17]. Similar to the paper [17, Section 3], we employ $\delta$-nets on the second system ${\mathcal{H}}_{B}$. In the following, we assume that the first system ${\mathcal{H}}_{A}$ has a larger dimension. That is, to optimize the choice of a separable state, we fix as many states on ${\mathcal{H}}_{B}$ as possible, and optimize the choice of states on ${\mathcal{H}}_{A}$. Then, we define an approximate value of the proposed measure. Employing the quantum Arimoto-Blahut algorithm proposed by the paper [18], we propose an algorithm for the approximate value. This algorithm is an iterative algorithm, and we show its convergence to the global minimum and its convergence speed.
As the third step, we apply our algorithm to the membership problem of separability. Although it is known that this problem is NP-hard [14, 15, 16, 17], several algorithms have been proposed for this problem. As a typical method, Doherty et al. [19, 20] proposed a method to identify whether a given state is a separable state or not by considering “symmetric extensions” and “PPT-symmetric extensions” in the second system ${\mathcal{H}}_{B}$. This kind of extension is called DPS hierarchy. Later, Navascués et al. [21] evaluated its calculation complexity. DPS hierarchy was discussed in [22, (14)], [23, 24]. The paper [17] proposed another algorithm for this problem by using $\delta$-nets on the second system ${\mathcal{H}}_{B}$. The paper [17] also discussed its calculation complexity. It was shown that the calculation complexity of these preceding algorithms are exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. The papers [21, 17] evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in their algorithms. The paper [21] evaluated the exponential component for the dimension of the first system ${\mathcal{H}}_{A}$, but the paper [17] evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. This paper discussed our calculation complexity, and showed that the calculation complexity of our algorithm is exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. In addition, we evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in our algorithm, and evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. As the comparison with the evaluation by [21], we find that our exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ is larger than that in the analysis [21], but our polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$ is smaller than that in the analysis [21]. This fact shows that our algorithm has an advantage over the algorithm based on “symmetric extensions” and “PPT-symmetric extensions” when the dimension of the second system ${\mathcal{H}}_{B}$ is fixed and only the dimension of the first system ${\mathcal{H}}_{A}$ increases.
Finally, we apply our algorithm to finding a separation hyperplane. When we solve a conic linear programming with a general cone, it is a key step to find a separation hyperplane for an element that does not belong to the cone [25, 26, 27, 28]. Recently, a conic linear programming with the separable cone has been applied to several problems in quantum information [29, 30]. Therefore, it is desired to find a separation hyperplane for the separable cone. In quantum information, a separation hyperplane for the separable cone is called an entanglement witness [31, 32, 33, 34, 35]. In this paper, based on our algorithm, we also propose an algorithm to find an entanglement witness, i.e., a separation hyperplane for an entangled state.
The remainder of this paper is organized as follows. Section II formulates the detection of an entangled state as a problem of hypothesis testing under the iid setting. Section III addresses this problem under the non-iid setting. Section IV explains information geometrical characterization of our proposed measure and its relation to entanglement witness. Section V proposes and studies new coherence measures as the preparation of the calculation of our entanglement measures. Section VI derives several calculation formulas for our entanglement measures for maximally entangled states based on the analysis in Section V. Section VII proposes algorithms for our entanglement measures by using $\delta$-net. Section VIII applies our results to the membership problem for separability and separation hyperplane. Section IX numerically calculates our entanglement measure, and check how our method works for the membership problem. Section IV explains that our method finds an entanglement witness. Section X presents our exponential version of quantum Sanov theorem. Section XI makes conclusions and discussions. Appendix A presents concrete constructions of $\delta$-nets. Appendix B proves the theorem for the relation between our entanglement measures and our proposed coherence measure. Appendix C proves theorems for the convergence for our algorithms. Appendix D proves our exponential version of quantum Sanov theorem. Appendix E proves our exponential evaluation of error probabilities in our problem of hypothesis testing under the non-iid setting.
II Detecting entangled state with iid setting
We consider the joint system ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$, and discuss how efficiently an entangled state $\rho_{AB}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ is distinguished from the set ${\cal S}_{AB}$ of separable states on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. This problem can be formulated as a hypothesis testing problem to distinguish $\rho_{AB}$ from ${\cal S}_{AB}$. For this detection, we assume the $n$-fold independently and identical distributed (iid) condition. That is, we have the following two hypotheses for the possible state as follows.
$H_{0}:$ The state is $\hbox{The state is }\rho_{AB}^{\otimes n}.$ (1)
$H_{1}:$ The state is $\hbox{The state is }\sigma_{AB}^{\otimes n}\hbox{ with }\sigma_{AB}\in{\cal S}_{AB}.$ (2)
To address this problem, we focus on the following quantity.
$\beta_{\epsilon,n}(\rho_{AB})$
::=:= $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}^{\otimes n}\leq\epsilon\Big{\}}.$ (3)
In the above setting, we do not impose the LOCC restriction to our measurement unlike [1, 33]. However, these existing studies did not discuss the optimal detection efficiency for the entanglement verification. Hence, the above quantity expresses the performance of discrimination at the fundamental level.
Here, it is better to explain the relation with the entanglement verification [36, 37, 38, 39, 40, 41]. Entanglement verification is the task of identifying a given entangled state within a certain error. Since entanglement verification is a more difficult task, the detection efficiency of the entanglement verification is worse than the detection efficiency of the entanglement detection. Existing studies for entanglement verification are limited to the verification for several special entangled states. No general theory covers the entanglement verification for general entangled states.
The application of quantum Sanov theorem [11, 12] guarantees the following
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{\epsilon,n}(\rho_{AB})=E^{*}(\rho_{AB})$
::=:= $\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\sigma_{AB}\|\rho_{AB}),$ (4)
where $D(\sigma\|\rho):=\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log\rho)$. This value is different from entanglement of relative entropy $E(\rho_{AB}):=\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\rho_{AB}\|\sigma_{AB})$. This fact shows that our measure $E^{*}(\rho_{AB})$ expresses the limit of the performance to distinguish the entangled state $\rho_{AB}$ from ${\cal S}_{AB}$. In contrast, when the choice of $H_{0}$ and $H_{1}$ is exchanged, $E(\rho_{AB})$ does not have the same meaning [10]. The proof of (4) is given as follows. The part $\leq$ in (4) follows from the strong converse part of quantum Stein’s lemma. The paper [11] showed the part $\geq$ in (4). However, they did not consider the case when $\epsilon$ is exponentially small. To enhance the meaning of the quantity $E^{*}(\rho_{AB})$, it is needed to evaluate $\beta_{\epsilon,n}(\rho_{AB})$ in such a case.
Table 1: List of mathematical symbols for detecting entangled state from Section II onwards.
Symbol Description Eq. number
${\cal S}_{AB}$ Set of separable states on
${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$
$\beta_{\epsilon,n}(\rho_{AB})$ Minimum error probability (3)
with constraint $\epsilon$ when
$H_{1}$ is given as (2)
$E^{*}(\rho_{AB})$ Minimum relative entropy (4)
$E_{\alpha}^{*}(\rho_{AB})$ Minimum relative Rényi (5)
entropy of order $\alpha$
${\cal S}_{AB}^{n}$ Set of separable states on
${\mathcal{H}}_{A}^{\otimes n}\otimes{\mathcal{H}}_{B}^{\otimes n}$
$\overline{\beta}_{\epsilon,n}(\rho_{AB})$ Minimum error probability (3)
with constraint $\epsilon$ when
$H_{1}$ is given as (10)
$\overline{E}^{*}(\rho_{AB})$ Normalized minimum (12)
relative entropy
$\overline{E}_{\alpha}^{*}(\rho_{AB})$ Normalized minimum relative (13)
Rényi entropy of order $\alpha$
$\tilde{E}^{*}(\rho_{AB})$ The above value with (18)
$\alpha=1-0$
$\sigma_{AB}^{*}$ Closest separable state (19)
${\cal S}_{D}$ Divergence sphere (25)
${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ Separation hyperplane (20)
$|\Phi\rangle$ Maximally entangled state
$\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle|j\rangle$
$E_{{\cal D}}^{*}(\rho_{AB})$ Minimum relative entropy (58)
with fixed finite set ${\cal D}$
$\epsilon_{2}({\cal D})$ (59)
$E_{{\alpha,\cal D}}^{*}(\rho_{AB})$ Minimum relative Rényi (70)
entropy of order $\alpha$
with fixed finite set ${\cal D}$
For this aim, using the sandwich relative entropy $D_{\alpha}(\sigma\|\rho):=\frac{1}{\alpha-1}\log\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$, we generalize $E^{*}(\rho_{AB})$ as
$E_{\alpha}^{*}(\rho_{AB}):=\min_{\sigma_{AB}\in{\cal S}_{AB}}D_{\alpha}(\sigma_{AB}\|\rho_{AB}).$ (5)
As presented in Section X, we have the following theorem.
Theorem 1
We have the following relations
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha<1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha},$ (6)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho_{AB})$
$\leq$ $\inf_{\alpha\geq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha}.$ (7)
More precisely, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha<1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha}$
$-\frac{(d_{AB}+2)(d_{AB}-1)}{n}\log(n+1),$ (8)
where $d_{AB}$ is the dimension of ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$.
Under the limit $r\to 0$, the RHS of (6) goes to $E_{1}^{*}(\rho_{AB}):=E^{*}(\rho_{AB})$. Hence, the relation (6) recovers the inequality $\geq$ in (4). When a TP-CP map $\Lambda$ preserves the set ${\cal S}_{AB}$, these quantities satisfy the monotonicity as
$E^{*}(\rho_{AB})\geq E^{*}(\Lambda(\rho_{AB})),\quad E_{\alpha}^{*}(\rho_{AB})\geq E_{\alpha}^{*}(\Lambda(\rho_{AB})).$ (9)
III Detecting entangled state with non-iid setting
The above discussion assumes the iid assumption for the alternative hypothesis. However, several existing studies for entanglement verification and entanglement detection assume that the alternative hypothesis includes general separable states including non-iid separable stets [5, 4, 39, 40, 41], which is called the non-iid setting. Our analysis can be extended to the non-iid setting as follows. We denote the set of separable states on ${\mathcal{H}}_{A}^{\otimes n}\otimes{\mathcal{H}}_{B}^{\otimes n}$ by ${\cal S}_{AB}^{n}$. Then, the alternative hypothesis is changed to
$H_{1}:$ The state is $\hbox{The state is }\sigma\hbox{ with }\sigma\in{\cal S}_{AB}^{n}$ (10)
while the null hypothesis is kept to (1). To address this problem, we focus on the following quantity.
$\overline{\beta}_{\epsilon,n}(\rho_{AB})$
::=:= $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (11)
The relation (4) yields that
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{\epsilon,n}(\rho_{AB})\leq\overline{E}^{*}(\rho_{AB}):=\inf_{n}\frac{1}{n}{E}^{*}(\rho_{AB}^{\otimes n})$
$=$ $\lim_{n\to\infty}\frac{1}{n}{E}^{*}(\rho_{AB}^{\otimes n})=\lim_{n\to\infty}\frac{1}{n}\min_{\sigma\in{\cal S}_{AB}^{n}}D(\sigma\|\rho_{AB}^{\otimes n}).$ (12)
Extending ${E}_{\alpha}^{*}(\rho_{AB})$, we define
$\overline{E}_{\alpha}^{*}(\rho_{AB}):=$ $\inf_{n}\frac{1}{n}{E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})=\lim_{n\to\infty}\frac{1}{n}{E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})$
$=$ $\lim_{n\to\infty}\frac{1}{n}\min_{\sigma_{AB}\in{\cal S}_{AB}^{n}}D_{\alpha}(\sigma_{AB}\|\rho_{AB}^{\otimes n}).$ (13)
As shown in Appendix E, we have the following theorem.
Theorem 2
We have
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha\leq 1}\frac{(1-\alpha)\overline{E}_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha},$ (14)
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{1-e^{-nr},n}(\rho_{AB})$
$=$ $\inf_{\alpha\geq 1}\frac{(1-\alpha)\overline{E}_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha}.$ (15)
More precisely, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha\leq 1}\frac{\frac{1}{n}(1-\alpha)E_{\alpha}^{*}(\rho_{AB}^{\otimes n})-\alpha r}{1-\alpha}$
$-\frac{(d_{AB}+2)(d_{AB}-1)}{n}\log(n+1),$ (16)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho_{AB})$
$\leq$ $\inf_{\alpha\geq 1}\frac{\frac{1}{n}(1-\alpha){E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})-\alpha r}{1-\alpha}.$ (17)
As an inequality opposite to (4), the inequality (14) implies the following inequality.
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{\epsilon,n}(\rho_{AB})\geq\tilde{E}^{*}(\rho_{AB}):=\lim_{\delta\to+0}\overline{E}_{1-\delta}^{*}(\rho_{AB}).$ (18)
While $\overline{E}^{*}(\rho_{AB})\geq\tilde{E}^{*}(\rho_{AB})$, it is not clear whether equality holds in general. When the quantity $E_{\alpha}^{*}(\rho_{AB})$ satisfies the additivity $E_{\alpha}^{*}(\rho_{AB,1}\otimes\rho_{AB,2})=E_{\alpha}^{*}(\rho_{AB,1})+E_{\alpha}^{*}(\rho_{AB,2})$, the equations ${E}^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})=\tilde{E}^{*}(\rho_{AB})$ and ${E}_{\alpha}^{*}(\rho_{AB})=\overline{E}_{\alpha}^{*}(\rho_{AB})$ hold. Therefore, it is crucial to clarify when the above additivity holds.
IV Information-geometrical characterization
Given a non-separable state $\rho_{AB}\notin{\cal S}_{AB}$, we discuss how to find an entanglement witness to distinguish $\rho_{AB}$ and ${\cal S}_{AB}$. For this aim, we employ the minimizer of $E^{*}(\rho_{AB})$, i.e., the state
$\sigma_{AB}^{*}:=\mathop{\rm argmin}\limits_{\sigma_{AB}\in{\cal S}_{AB}}D(\sigma_{AB}\|\rho_{AB}).$ (19)
Then, using the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})$, we define the hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ as (See Fig. 1.)
${\cal M}(\sigma_{AB}^{*},\rho_{AB})$
::=:= $\{\rho^{\prime}_{AB}|\operatorname{\mathrm{Tr}}\rho_{AB}^{\prime}(\log\sigma_{AB}^{*}-\log\rho_{AB})=D(\sigma_{AB}^{*}\|\rho_{AB})\}.$ (20)
Figure 1: Entanglement witness: This figure shows the relation among the entangled state $\rho_{AB}$, the set of separable states ${\cal S}_{AB}$, and the separation hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$.
The following theorem guarantees that ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ separate $\rho_{AB}$ from ${\cal S}_{AB}$. Hence, the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})-D(\sigma_{AB}^{*}\|\rho_{AB})$ works as entanglement witness for $\rho_{AB}$.
Theorem 3
(i) For any separable state $\sigma_{AB}\in{\cal S}$, we have
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}-\log\rho_{AB})\geq D(\sigma_{AB}^{*}\|\rho_{AB}).$ (21)
(ii) Also, we have
$\operatorname{\mathrm{Tr}}\rho_{AB}(\log\sigma_{AB}^{*}-\log\rho_{AB})=-D(\rho_{AB}\|\sigma_{AB}^{*})<0.$ (22)
To grasp the geometrical meaning of the above fact, we employ information geometry. Given a Hermitian matrix $X$, we often focus on the set :=${\cal E}:=\{\rho_{t}\}$, where $\rho_{t}:=\exp(tX+\log\rho_{0})/c_{t}$ with $c_{t}:=\operatorname{\mathrm{Tr}}\exp(tX+\log\rho_{0})$. The set $\{\rho_{t}\}$ is called an exponential family generated by $X$. Also, the set :=${\cal M}:=\{\rho|\operatorname{\mathrm{Tr}}\rho X=\operatorname{\mathrm{Tr}}\rho_{0}X\}$ is called a mixture family generated by $X$. Any elements $\sigma\in{\cal M}$ and $\rho\in{\cal E}$ satisfy
$D(\sigma\|\rho)=D(\sigma\|\rho_{0})+D(\rho_{0}\|\rho),$ (23)
which is called Pythagorean theorem [13], [42, Theorem 2.3]. We apply (23) to the case with $X=-(\log\sigma_{AB}^{*}-\log\rho_{AB})$ and $\rho_{0}=\sigma_{AB}^{*}$. Any element $\sigma_{AB}^{\prime}\in{\cal M}(\sigma_{AB}^{*},\rho_{AB})$ satisfies
$D(\sigma_{AB}^{\prime}\|\rho_{AB})=D(\sigma_{AB}^{\prime}\|\sigma_{AB}^{*})+D(\sigma_{AB}^{*}\|\rho_{AB}).$ (24)
Proof: Since (ii) is trivial, we show only (i) by contradiction. We assume that (i) does not hold. Hence, there exists a separable state $\sigma_{AB}\in{\cal S}$ that does not satisfy (21).
We define the divergence sphere
${\cal S}_{D}:=\{\rho_{AB}^{\prime}|D(\rho_{AB}^{\prime}\|\rho_{AB})=D(\sigma_{AB}^{*}\|\rho_{AB})\}.$ (25)
Then, ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ is the tangent space of the sphere ${\cal S}_{D}$ at $\sigma_{AB}^{*}$. If a separable state $\sigma_{AB}$ satisfies
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}-\log\rho_{AB})<D(\sigma_{AB}^{*}\|\rho_{AB}),$ (26)
there exists $\lambda\in(0,1)$ such that $\sigma_{AB}^{\prime}:=\lambda\sigma_{AB}+(1-\lambda)\sigma_{AB}^{*}$ is inside of the sphere ${\cal S}_{D}$ as Fig. 2. Hence,
$D(\sigma_{AB}^{\prime}\|\rho_{AB})<D(\sigma_{AB}^{*}\|\rho_{AB}).$ (27)
Since $\lambda\sigma_{AB}+(1-\lambda)\sigma_{AB}^{*}$ is also a separable state, the relation (27) contradicts with the choice of $\sigma_{AB}^{*}$.
Figure 2: Divergence sphere: This figure shows the relation among the entangled state $\rho_{AB}$, the divergence sphere ${\cal S}_{D}$, and the separation hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$.
V Coherence measure
For further analysis on our entanglement measures, we study the coherence measures defined below, because, as explained in the next section, our entanglement measures are related to these coherence measures. We consider a state $\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$ on a $d$-dimensional Hilbert ${\mathcal{H}}$ spanned by $\{|j\rangle\}_{j=0}^{d-1}$, and denote the set of incoherent states by ${\cal I}_{c}$, which is written as $\{\sigma(p)\}$ with $\sigma(p):=\sum_{i}p_{i}|i\rangle\langle i|$, where $p=(p_{i})$ is a distribution. Then, we define the following coherence measures.
$C^{*}(\rho):=\min_{\sigma\in{\cal I}_{c}}D(\sigma\|\rho),\quad C_{\alpha}^{*}(\rho):=\min_{\sigma\in{\cal I}_{c}}D_{\alpha}(\sigma\|\rho).$ (28)
Although the papers [43, 44] considered the coherence measures based on Rényi relative entropy, our measures are different from their coherence measures because their measures are based on the quantities with exchanging the two input states from ours. In fact, $C^{*}(\rho)$ has a geometrical characterization similar to Section IV by replacing ${\cal S}_{AB}$ by ${\cal I}_{c}$.
For $\alpha\in(0,1)$ and a pure state $|\phi\rangle\langle\phi|$ with $|\phi\rangle=\sum_{j}e^{\theta_{j}\sqrt{-1}}p_{j}|j\rangle$, we have
$e^{(\alpha-1)C_{\alpha}^{*}(|\phi\rangle\langle\phi|)}=\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(|\phi\rangle\langle\phi|^{\frac{1-\alpha}{2\alpha}}\sigma|\phi\rangle\langle\phi|^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(|\phi\rangle\langle\phi|\sigma|\phi\rangle\langle\phi|)^{\alpha}$
$=$ $\max_{\sigma(q)\in{\cal I}_{c}}(\langle\phi|\sigma(q)|\phi\rangle)^{\alpha}=\max_{q}(\sum_{j}q_{j}p_{j})^{\alpha}=(\max_{i}p_{i})^{\alpha}.$ (29)
Hence, we have
$C_{\alpha}^{*}(|\phi\rangle\langle\phi|)=\frac{\alpha}{\alpha-1}\log(\max_{i}p_{i}).$ (30)
Also, $C^{*}(|\phi\rangle\langle\phi|)$ and $C_{\alpha}^{*}(|\phi\rangle\langle\phi|)$ take infinity value for $\alpha>1$ unless it is an incoherent state.
In the following discussion, we employ the operators $Z:=\sum_{j=0}^{d-1}\omega^{j}|j\rangle\langle j|$ with $\omega:=e^{2\pi\sqrt{-1}/d}$ and $X:=\sum_{j=0}^{d-1}|j+1\rangle\langle j|$. We define $\rho(p):=\sum_{j=0}^{d-1}p_{j}Z^{j}|+\rangle\langle+|Z^{-j}$ with $|+\rangle:=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle$. Then, we have
$C^{*}(\rho(p))$ $=D(p_{\mathop{\rm mix}}\|p)=-\frac{1}{d}\sum_{j=0}^{d-1}\log p_{j}-\log d$ (31)
$C_{\alpha}^{*}(\rho(p))$ $=D_{\alpha}(p_{\mathop{\rm mix}}\|p)$
$=\frac{1}{\alpha-1}\log\sum_{j=0}^{d-1}p_{j}^{1-\alpha}-\frac{\alpha}{\alpha-1}\log d,$ (32)
where $p_{mix}$ is the uniform distribution. Since the above second equations follow from the direct calculation, the following analysis shows the first equations.
We have $X\rho(p)X^{-1}=\rho(p).$ Thus, $D(\sigma(q)\|\rho(p))=D(X^{j}\sigma(q)X^{-j}\|\rho(p))$. Hence, the joint convexity of the relative entropy implies
$D(\sigma(q)\|\rho(p))=\sum_{j=0}^{d-1}\frac{1}{d}D(X^{j}\sigma(q)X^{-j}\|\rho(p))$
$\geq$ $D(\sum_{j=0}^{d-1}\frac{1}{d}X^{j}\sigma(q)X^{-j}\|\rho(p))$
$=$ $D(\rho_{mix}\|\rho(p))=D(\rho(p_{mix})\|\rho(p))=D(p_{mix}\|p),$ (33)
where $\rho_{mix}$ is the completely mixed state. The above relations imply the first equation (31). Similarly, since $x\mapsto x^{\alpha}$ is matrix convex for $\alpha>1$,
$e^{(\alpha-1)D_{\alpha}(\sigma(q)\|\rho(p))}=\sum_{j=0}^{d-1}\frac{1}{d}e^{(\alpha-1)D_{\alpha}(X^{j}\sigma(q)X^{j}\|\rho(p))}$
$\geq$ $e^{(\alpha-1)D_{\alpha}(\sum_{j=0}^{d-1}\frac{1}{d}X^{j}\sigma(q)X^{j}\|\rho(p))}$
$=$ $e^{(\alpha-1)D_{\alpha}(\rho_{mix}\|\rho(p))}=e^{(\alpha-1)D_{\alpha}(p_{mix}\|p)},$ (34)
which implies the first equation (32) for $\alpha>1$. Similarly, since $x\mapsto x^{\alpha}$ is matrix concave for $\alpha\in[0,1)$, an analysis similar to (34) yields the first equation (32) for $\alpha\in[0,1)$.
To calculate $C^{*}(\rho)$ for a general state, we define the state $I_{c}(\rho):=\frac{1}{t(\rho)}\exp((\log\rho)^{\rm diag})$ with $t(\rho):=\operatorname{\mathrm{Tr}}\exp((\log\rho)^{\rm diag})$ and $(\log\rho)^{\rm diag}$ is the diagonal matrix whose diagonal elements are the same as $\log\rho$. Then, $C^{*}(\rho)$ is calculated as follows.
Lemma 4
The relation
$C^{*}(\rho)=D(I_{c}(\rho)\|\rho)$ (35)
holds.
Proof: For $\sigma\in{\cal I}_{c}$, we have
$D(\sigma\|\rho)=\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log\rho)$
$=$ $\operatorname{\mathrm{Tr}}\sigma(\log\sigma-(\log\rho)^{\rm diag}))$
$=$ $\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log I_{c}(\rho)-\log t(\rho)$
$=$ $D(\sigma\|I_{c}(\rho))-\log t(\rho).$ (36)
Hence, the minimization is realized when $\sigma=I_{c}(\rho)$.
Lemma 5
The relation
$C_{\alpha}^{*}(\rho)$
$=$ $\left\{\begin{array}[]{l}\max_{\tau}\frac{\alpha}{\alpha-1}\log\|(\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|\\ \hskip 150.69397pt\hbox{ for }\alpha\in(0,1)\\ \max_{\tau}\frac{-\alpha}{\alpha-1}\log\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|\\ \hskip 150.69397pt\hbox{ for }\alpha>1\end{array}\right.$ (41)
holds.
Proof: The idea of this proof comes from the proof of [48, Lemma 6 and (B4)]. We choose $\beta$ such that $\frac{1}{\beta}+\frac{1}{\alpha}=1$ so that $\beta=\frac{\alpha-1}{\alpha}$. We fix $\alpha\in(0,1)$ so that $\beta<0$. For any state $\tau$ and any incoherent state $\sigma\in{\cal I}_{c}$, the reverse matrix Hölder inequality [51, Theorem 2.7] implies
$\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}\geq(\operatorname{\mathrm{Tr}}\tau)^{\beta}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}$
$=$ $(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (42)
Considering the equality condition, we have
$\min_{\tau}\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}=(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (43)
Hence,
$e^{(\alpha-1)C_{\alpha}^{*}(\rho)}=\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $(\max_{\sigma\in{\cal I}_{c}}\min_{\tau}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $(\min_{\tau}\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
$=$ $\min_{\tau}\|(\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|^{\alpha}.$ (44)
Here, $(a)$ follows from the minimax theorem because $\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma$ is linear for $\sigma$ and is convex for $\tau$. Hence, we obtain the desired equation with $\alpha\in(0,1)$.
Next, we fix $\alpha>1$ so that $01>\beta>$ 0. The matrix Hölder inequality [52],[51, Theorem 2.6] implies
$\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}\leq(\operatorname{\mathrm{Tr}}\tau)^{\beta}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}$
$=$ $(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (45)
Considering the equality condition, we have
$\max_{\tau}\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}=(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (46)
Hence,
$e^{(\alpha-1)C_{\alpha}^{*}(\rho)}=\min_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $(\min_{\sigma\in{\cal I}_{c}}\max_{\tau}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $(\max_{\tau}\min_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
$=$ $(\max_{\tau}\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|^{-1})^{\alpha}$
$=$ $\max_{\tau}\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|^{-\alpha}.$ (47)
Here, $(a)$ follows from the minimax theorem because $\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma$ is linear for $\sigma$ and is concave for $\tau$. Hence, we obtain the desired equation with $\alpha>1$.
Coherence measures satisfy the additivity.
Lemma 6
The relations
$\begin{split}C^{*}(\rho_{1}\otimes\rho_{2})&=C^{*}(\rho_{1})+C^{*}(\rho_{2}),\\ C_{\alpha}^{*}(\rho_{1}\otimes\rho_{2})&=C_{\alpha}^{*}(\rho_{1})+C_{\alpha}^{*}(\rho_{2})\end{split}$ (48)
hold.
Proof: Since the first equation in (48) follows from the second equation in (48) with taking the limit $\alpha\to 1$, we show only the second equation in (48). The idea of this proof is similar to the proof of [48, Lemma 7]. For $\alpha\in(0,1)$, using (41), we have
$C_{\alpha}^{*}(\rho_{1})+C_{\alpha}^{*}(\rho_{2})\geq C_{\alpha}^{*}(\rho_{1}\otimes\rho_{2})$
$=$ $\max_{\tau}\frac{\alpha}{\alpha-1}\log\|((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|$
$\geq$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}\log\|((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}(\tau_{1}\otimes\tau_{2})^{-\frac{1-\alpha}{\alpha}}$
$\hskip 86.11084pt\cdot(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|$
$=$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}(\log\|(\rho_{1}^{\frac{1-\alpha}{2\alpha}}\tau_{1}^{-\frac{1-\alpha}{\alpha}}\rho_{1}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|$
$+\log\|(\rho_{2}^{\frac{1-\alpha}{2\alpha}}\tau_{2}^{-\frac{1-\alpha}{\alpha}}\rho_{2}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|)$
$=$ $C_{\alpha}^{*}(\rho_{1})+C_{\alpha}^{*}(\rho_{2}),$ (49)
which implies the second equation in (48). The case with $\alpha>1$ can be shown in the same way.
Table 2: List of mathematical symbols for coherence.
Symbol Description Eq. number
${\cal I}_{c}$ Set of incoherent states
$C^{*}(\rho)$ Minimum relative entropy (28)
$C_{\alpha}^{*}(\rho)$ Minimum relative Rényi (28)
entropy of order $\alpha$
$I_{c}(\rho)$ Closet incoherent state (35)
$\sigma(p)$ $\sum_{i}p_{i}|i\rangle\langle i|$
$\rho(p)$ $\sum_{j=0}^{d-1}p_{j}Z^{j}|+\rangle\langle+|Z^{-j}$
Remark 1
It is possible to make the same discussion for the coherence as Sections IIIV. Since the coherence measures satisfy the additivity as Lemma 6, the detectability even with the non-iid setting like Section III can be characterized by the quantities $C^{*}(\rho)$ and $C_{\alpha}^{*}(\rho)$.
VI Maximally correlated states
The discussion in Section II shows the operational meaning of $E_{\alpha}^{*}(\rho_{AB})$ including $E^{*}(\rho_{AB})$. We calculate $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ for several special cases. When $\rho_{AB}$ is a pure state $|\phi\rangle\langle\phi|$ $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha>1$ are infinity. However, $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha\in[0,1)$ takes a finite value as follows. When the Schmidt coefficients of $|\phi\rangle\langle\phi|$ are $\sqrt{p_{i}}$ with $i=1,\ldots,d$, $\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma|\phi\rangle=\max_{i}p_{i}$ [44, (40) and (97)]. Hence,
$e^{(\alpha-1)E_{\alpha}^{*}(\rho_{AB})}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}(|\phi\rangle\langle\phi|\sigma_{AB}|\phi\rangle\langle\phi|)^{\alpha}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle^{\alpha}$
$=$ $(\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle)^{\alpha}=(\max_{i}p_{i})^{\alpha}.$ (50)
Hence, we have
$E_{\alpha}^{*}(\rho_{AB})=\frac{\alpha}{\alpha-1}\log(\max_{i}p_{i}).$ (51)
More generally, the calculation of these quantities can be converted to the calculation of the coherence measure for special states $\rho_{AB}$. Similar to the paper [44], we consider maximally correlated states. A state $\rho_{AB}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ is called a maximally correlated state when there exist bases $\{|u_{j,A}\rangle\}_{j}$ and $\{|u_{j,B}\rangle\}_{j}$ of ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ such that the support $\rho_{AB}$ is included in the subspace spanned by $\{|u_{j,A}\rangle|u_{j,B}\rangle\}_{j}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ [49, 50], [42, Section 8.8]. In this case, the state $\rho_{AB}$ is written as $\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|$. In this case, the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ can be simplified as follows. Similar to Theorem 1 in [44], the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ are reduced to coherence measures.
Lemma 7
The relations
$E^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=$ $C^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}$ (52)
$E_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=$ $C_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}$ (53)
hold.
Lemma 7 is shown in Appendix B. Since a pure state is a maximally correlated state, the calculation (51) can be recovered by Lemma 7 and (30). Also, $E^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$ can be calculated by Lemma 4 as
$E^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=D(I_{c}(\rho_{\theta})\|\rho_{\theta}),$ (54)
where $\rho_{\theta}:=\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$. The combination of Lemmas 6 and 7 guarantees that any maximally entangled state $\rho_{AB}$ satisfies
$E^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})=\tilde{E}^{*}(\rho_{AB}),~{}E_{\alpha}^{*}(\rho_{AB})=\overline{E}_{\alpha}^{*}(\rho_{AB}).$ (55)
The iid setting and the non-iid setting have the same bounds.
To find another example, we set ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ to be $d$-dimensional spaces, and define the maximally entangled state $|\Phi\rangle:=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle|j\rangle$. We define the state $\rho_{AB}(p):=\sum_{j=0}^{d-1}p_{j}Z^{j}\otimes I|\Phi\rangle\langle\Phi|Z^{-j}\otimes I$. The combination of (31), (32), Lemma 7 implies
$E^{*}(\rho_{AB}(p))$ $=D(p_{\mathop{\rm mix}}\|p)=-\frac{1}{d}\sum_{j=0}^{d}\log p_{j}-\log d$ (56)
$E_{\alpha}^{*}(\rho_{AB}(p))$ $=D_{\alpha}(p_{\mathop{\rm mix}}\|p)$
$=\frac{1}{\alpha-1}\log\sum_{j=0}^{d-1}p_{j}^{1-\alpha}-\frac{\alpha}{\alpha-1}\log d.$ (57)
However, it is not easy to calculate these quantities for general states. Therefore, it is needed to develop an efficient method for their calculation in the general case.
VII Algorithms
VII.1 Algorithm for $E^{*}(\rho_{AB})$
To calculate the proposed quantities for general states, we introduce algorithms for their calculation that can be applied to general states. Since the calculation of $E^{*}(\rho_{AB})$ is hard, we modify $E^{*}(\rho_{AB})$ as follows. We choose a finite set ${\cal D}$ of pure states on ${\cal H}_{B}$ as a discretization. That is, we denote ${\cal D}$ as $\{\rho_{B|c}^{*}\}_{c\in{\cal X}}$. Then, we define
$E_{{\cal D}}^{*}(\rho_{AB}):=\min_{P(c),\rho_{A|c}}D\Big{(}\sum_{c}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (58)
We define $\epsilon_{2}({\cal D})$ as
$\epsilon_{2}({\cal D}):=2\sqrt{1-\min_{\psi_{1}}\max_{\psi_{2}\in{\cal D}}|\langle\psi_{1}|\psi_{2}\rangle|^{2}}.$ (59)
We have
$\max_{P^{\prime}(c),\rho_{A|c}^{\prime},\rho_{B|c}^{\prime}}\min_{P(c),\rho_{A|c}}\Big{\|}\sum_{c}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*}$
$-\sum_{c}P^{\prime}(c)\rho_{A|c}^{\prime}\otimes\rho_{B|c}^{\prime}\Big{\|}_{1}$
$\leq$ $\max_{\psi_{1}}\min_{\psi_{2}\in{\cal D}}\||\psi_{1}\rangle\langle\psi_{1}|-|\psi_{2}\rangle\langle\psi_{2}|\|_{1}$
$=$ $2\sqrt{1-\min_{\psi_{1}}\max_{\psi_{2}\in{\cal D}}|\langle\psi_{1}|\psi_{2}\rangle|^{2}}=\epsilon_{2}({\cal D}).$ (60)
Therefore, Fannes inequality implies
$\max_{P^{\prime}(c),\rho_{A|c}^{\prime},\rho_{B|c}^{\prime}}\min_{P(c),\rho_{A|c}}\Big{|}D\Big{(}\sum_{x}P(x)\rho_{A|x}\otimes\rho_{B|x}^{*}\Big{\|}\rho_{AB}\Big{)}$
$\hskip 86.11084pt-D\Big{(}\sum_{x}P^{\prime}(x)\rho_{A|x}^{\prime}\otimes\rho_{B|x}^{\prime}\Big{\|}\rho_{AB}\Big{)}\Big{|}$
$\leq$ $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D})),$ (61)
where $\eta_{0}$ is defined as
$\eta_{0}(x):=\left\{\begin{array}[]{ll}-x\log x&\hbox{when }0\leq x\leq 1/e\\ 1/e&\hbox{otherwise.}\end{array}\right.$ (64)
Therefore, we have
$E_{{\cal D}}^{*}(\rho_{AB})$
$-\Big{(}\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))\Big{)}$
$\leq$ $E^{*}(\rho_{AB})\leq E_{{\cal D}}^{*}(\rho_{AB}).$ (65)
We rewrite $P(c)\rho_{A|c}$ to $X_{c}$. Using the idea of the algorithm by [18], we propose an iterative algorithm to calculate $E_{{\cal D}}^{*}(\rho_{AB})$, in which matrices $X_{A|c}$ is updated as follows. Given the set $\{X_{c}^{(t)}\}_{c}$, we define $V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$ and $\kappa[\{X_{A|c}^{(t)}\}_{c}]$ as
$V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$
::=:= $\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}^{(t)}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$\hskip 86.11084pt-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)},$ (66)
$\kappa[\{X_{A|c}^{(t)}\}_{c}]:=\sum_{c}\operatorname{\mathrm{Tr}}\exp(X_{A|c}^{(t)}-V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (67)
Then, we choose the updated matrix ${X}_{A|c}^{(t+1)}$
${X}_{A|c}^{(t+1)}:=$ $\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (68)
This algorithm is summarized as Algorithm 1.
Algorithm 1 Minimization of $D\Big{(}\sum_{c}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$
1:Choose the initial value $\{X_{A|c}^{(t)}\}_{c}$;
2:repeat
3:     Calculate ${X}_{A|c}^{(t+1)}:=\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}])$;
4:until convergence.
Then, we have the following theorem.
Theorem 8
We choose ${X}_{A|c}^{(1)}$ such that $\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes{X}_{A|c}^{(1)}$ is the completely mixed state $\rho_{AC,\mathop{\rm mix}}$. Then, we have
$D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}-E_{{\cal D}}^{*}(\rho_{AB})\leq\frac{\log d_{A}|{\cal D}|}{t}.$ (69)
The above theorem is shown in Appendix C by using [18, Theorem 3.3].
VII.2 Algorithm for $E_{\alpha}^{*}(\rho_{AB})$
Next, we consider the algorithm for $E_{\alpha}^{*}(\rho_{AB})$. Similarly, instead of $E_{\alpha}^{*}(\rho_{AB})$, we calculate
$E_{\alpha,{\cal D}}^{*}(\rho_{AB}):=\min_{X_{A|c}}D_{\alpha}\Big{(}\sum_{c}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (70)
The following algorithm minimizes $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$. For this aim, we iteratively choose matrices $X_{A|c}$ as follows. Given the set $\{X_{c}^{(t)}\}_{c}$, we define $V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$ and $\kappa_{\alpha}[\{X_{A|c}^{(t)}\}_{c}]$ as
$V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$
::=:= $\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}^{(t)}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
$\cdot\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*}),$ (71)
$\kappa_{\alpha}[\{X_{A|c}^{(t)}\}_{c}]$
::=:= $\sum_{c}\operatorname{\mathrm{Tr}}\exp(X_{A|c}^{(t)}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (72)
Then, we choose the updated matrix ${X}_{A|c}^{(t+1)}$
${X}_{A|c}^{(t+1)}$
::=:= $\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (73)
Algorithm 2 Minimization of $\mathop{\rm sgn}(\alpha-1)e^{(\alpha-1)D_{\alpha}\Big{(}\sum_{c}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}}$
1:Choose the initial value $\{X_{A|c}^{(t)}\}_{c}$;
2:repeat
3:     Calculate ${X}_{A|c}^{(t+1)}:=\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}])$;
4:until convergence.
Then, we have the following theorem.
Theorem 9
We choose ${X}_{A|c}^{(1)}$ as $\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes{X}_{A|c}^{(1)}$ as the completely mixed state $\rho_{AC,\mathop{\rm mix}}$. For $\alpha\in(0,1)$, and a sufficiently large real number $\gamma>0$, we have
$\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}^{(t+1)}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$
$-\mathop{\rm sgn}(\alpha-1)e^{(\alpha-1)E_{\alpha,{\cal D}}(\rho_{AB})}\leq\frac{\gamma\log d_{A}|{\cal D}|}{t}.$ (74)
The above theorem is shown in Appendix C by using [18, Theorem 3.3].
VIII Application of our algorithm
VIII.1 Membership problem
Next, we apply our algorithm to the membership problem of the set ${\cal S}_{AB}$. That is, we employ the quantity $E^{*}(\rho_{AB})$ for the membership problem. This measure will be used for constructing the separation hyperplane. In the following, we consider how to distinguish the following two cases.
$H_{0}:$ $E^{*}(\rho_{AB})=0$ (75)
$H_{1}:$ $E^{*}(\rho_{AB})\geq\epsilon_{1}.$ (76)
Due to (65), when $E^{*}(\rho_{AB})=0$, we have
$E_{{\cal D}}^{*}(\rho_{AB})\leq\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D})).$ (77)
Also, when $E^{*}(\rho_{AB})>\epsilon$, we have
$E_{{\cal D}}^{*}(\rho_{AB})>\epsilon.$ (78)
Therefore, if the following two cases are distinguished, $H_{0}$ and $H_{1}$ are distinguished.
$H_{0}^{\prime}:E_{{\cal D}}^{*}(\rho_{AB})\leq$ $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})$
$+\eta_{0}(\epsilon_{2}({\cal D})),$ (79)
$H_{1}^{\prime}:E_{{\cal D}}^{*}(\rho_{AB})>$ $\epsilon_{1}$ (80)
under the assumption that $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))\leq\epsilon_{1}$.
We choose the initial state according to the assumption of Theorem 8. Then, when $H_{0}^{\prime}$ holds, i.e., $E_{{\cal D}}^{*}(\rho_{AB})\leq\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))$, we have
$D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$
$\leq$ $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))+\frac{\log d_{A}|{\cal D}|}{t}.$ (81)
Also, when $E_{{\cal D}}^{*}(\rho_{AB})>\epsilon$, we have
$D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}>\epsilon.$ (82)
Therefore, if the following two cases are distinguished, $H_{0}^{\prime}$ and $H_{1}^{\prime}$, i.e. $H_{0}$ and $H_{1}$ are distinguished.
$H_{0}^{\prime\prime}:$ $D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$
$\leq\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))$
$\quad+\frac{\log d_{A}|{\cal D}|}{t},$ (83)
$H_{1}^{\prime\prime}:$ $\operatorname{\mathrm{Tr}}\rho_{AC}^{(t+1)}\Omega_{F}[\rho_{AC}^{(t+1)}]>\epsilon_{1}$ (84)
under the assumption that $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))+\frac{\log d_{A}|{\cal D}|}{t}\leq\epsilon_{1}$. That is, the number $t$ of iterations needs to satisfy $t\geq\log d_{A}|{\cal D}|/\Big{(}\epsilon_{1}-(\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D})))\Big{)}$.
Next, we evaluate the computation complexity to distinguish $H_{0}$ and $H_{1}$ under the above method when we choose ${\cal D}_{n,d_{B}}$ as ${\cal D}$ in the way as Appendix A and we fix $\epsilon_{1}$ and move $d_{A},d_{B}$ with $d_{A}\geq d_{B}$.
In this case, we have $|{\cal D}_{n,d_{B}}|\leq(\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}})^{d_{B}-1}$ with $\epsilon_{2}({\cal D}_{n,d_{B}})=\epsilon_{2}$. The above condition is rewritten as
$t\geq\frac{\log d_{A}+(d_{B}-1)\log\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}}}{\epsilon_{1}-(\epsilon_{2}(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}))}$ (85)
Since $\epsilon_{1}$ is fixed, we need to choose $\epsilon_{2}$ to be $O(\frac{\epsilon_{1}}{\log d_{A}d_{B}})$, which realizes $\epsilon_{1}-(\epsilon_{2}(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}))=O(\epsilon_{1})$. The above condition is rewritten as
$t\geq$ $\frac{\log d_{A}+(d_{B}-1)\log\frac{2d_{B}^{1/2}\pi}{O(\frac{\epsilon_{1}}{\log d_{A}d_{B}})}}{O(\epsilon_{1})}$
$=$ $O\Big{(}\frac{\log d_{A}+(d_{B}-1)\log\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}}}{\epsilon_{1}}\Big{)}$
$=$ $O\Big{(}\frac{\log d_{A}+d_{B}\log d_{B}+d_{B}(\log\log d_{A}-\log\epsilon_{1})}{\epsilon_{1}}\Big{)}.$ (86)
However, each step has various calculations. In each step, we need to calculate ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$. For this calculation, we need to calculate $\log\rho_{AC}$ and $\Omega_{R,D}[\rho_{AC}]$. $\rho_{AC}$ is a block diagonal matrix. That is, it is sufficient to calculate every component. Every component has a matrix function on a $d_{A}d_{B}$-dimensional space, and a factor whose complexity is linear to $|{\cal D}_{n,d_{B}}|$. But, the latter part is common for all $c$-components. The farmer can be done by diagonalization. The complexity of the diagonalization on $d$-dimensional matrix is $O(d^{3})$ with error $\epsilon$ [45, 46, 47]. Therefore, the calculation complexity of ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$ is $O(|{\cal D}_{n,d_{B}}|(d_{A}^{3}d_{B}^{3}))\leq O((\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))=O((\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))$. The total complexity is
$O\Big{(}$ $\frac{\log d_{A}+d_{B}\log d_{B}+d_{B}(\log\log d_{A}-\log\epsilon_{1})}{\epsilon_{1}}$
$\cdot(\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3})\Big{)}.$ (87)
On the other hand, the method by [21] has the following calculation complexity
$O(d_{A}^{6}\epsilon^{-4d_{B}})$ (88)
In the above evaluation, $\epsilon$ takes the same role as $\epsilon_{1}$. For $d_{B}$, our bound has the factor $d_{B}^{(d_{B}-1)/2}$, which is larger than $\frac{1}{\epsilon^{4d_{B}}}$. However, when $d_{B}$ is also fixed and only $d_{A}$ increases, our bound is simplified to $O(\max(\log d_{A},-\log\epsilon_{1})(\log d_{A})^{d_{B}-1}d_{A}^{3}\epsilon_{1}^{-d_{B}})$. When $\epsilon_{1}$ and $\epsilon$ are fixed, our bound is simplified to $O((\log d_{A})^{d_{B}}d_{A}^{3})$, and their bound is simplified to $O(d_{A}^{6})$. In this case, our bound is smaller than their bound.
VIII.2 Separation hyperplane
Next, we discuss how to obtain separation hyperplane by using the above approximation and the geometrical characterization by Theorem 3. This is because it is a key step for conic linear programming with a general cone to find a separation hyperplane for an element that does not belong to the cone [25, 26, 27, 28]. For this aim, we employ the minimizer of $E_{{\cal D}}^{*}(\rho_{AB})$, i.e., the state
$\sigma_{AB}^{*}({\cal D}):=\mathop{\rm argmin}\limits_{\sigma_{AB}\in{\cal S}({\cal D})}D(\sigma_{AB}\|\rho_{AB}).$ (89)
Then, the following theorem means that the operator $(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})$ gives a separation hyperplane to distinguish $\rho_{AB}$ and ${\cal S}$.
Theorem 10
(i) For any separable state $\sigma_{AB}\in{\cal S}$, we have
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})$
$\geq$ $D(\sigma_{AB}^{*}({\cal D})\|\rho_{AB})-\|\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB}\|\epsilon_{2}({\cal D}).$ (90)
(ii) Also, we have
$\operatorname{\mathrm{Tr}}\rho_{AB}(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})=-D(\rho_{AB}\|\sigma_{AB}^{*}({\cal D}))<0.$ (91)
Proof: Since (ii) is trivial, we show only (i) by contradiction. Since the set ${\cal S}({\cal D})$ is the convex set, we have the following in the same way as Theorem 3. (i) For any separable state $\sigma_{AB}\in{\cal S}({\cal D})$, we have
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})\geq D(\sigma_{AB}^{*}({\cal D})\|\rho_{AB}).$ (92)
Due to the combination of (92) and (60), any separable state $\sigma_{AB}\in{\cal S}$ satisfies (90).
IX Numerical analysis
For a typical example, we focus on the case when ${\mathcal{H}}_{A}$ is two qubits and ${\mathcal{H}}_{B}$ is one qubit. We parametrize a qubit pure state as $|\phi_{\theta}\rangle:=\frac{1}{\sqrt{2}}(|0\rangle+e^{\sqrt{-1}\theta}|1\rangle)$. Consider the entangled state $|\Psi_{p}\rangle:=\sqrt{p}|\phi_{0}\rangle|\phi_{0}\rangle|0\rangle+\sqrt{1-p}|\phi_{\pi/2}\rangle|\phi_{\pi/2}\rangle|1\rangle)$. Given a real number $\lambda$ with $10<\lambda<$ 1, we focus on the phase-damping channel ${\cal E}_{\lambda}$ defined as
${\cal E}_{\lambda}\left(\left(\begin{array}[]{cc}a_{0,0}&a_{0,1}\\ a_{1,0}&a_{1,1}\end{array}\right)\right):=\left(\begin{array}[]{cc}a_{0,0}&\lambda a_{0,1}\\ {\lambda}a_{1,0}&a_{1,1}\end{array}\right).$ (97)
This channel acts on the state $|\phi_{\theta}\rangle$ as
${\cal E}_{\lambda}(\rho)=\frac{1+\lambda}{2}\rho+\frac{1-\lambda}{2}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\rho\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right).$ (102)
That is, the phase-damping channel is replaced by the above over the state $|\Psi_{p}\rangle$. Then, we define the state with two parameters $\lambda$ and $\delta$ as
$\rho_{p,\lambda,\delta}:=(1-\delta)({\cal E}_{\lambda}\otimes{\cal E}_{\lambda}\otimes Id)(|\Psi_{p}\rangle\langle\Psi_{p}|)+\delta\rho_{mix},$ (103)
where $\delta\rho_{mix}$ is the completely mixed state on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$.
Applying the unitary matrix $V=\frac{1}{2}\left(\begin{array}[]{cc}1+\sqrt{-1}&1-\sqrt{-1}\\ 1-\sqrt{-1}&1+\sqrt{-1}\end{array}\right)$ on both qubits in ${\mathcal{H}}_{A}$, the above model is converted to the following model. We define $|\tilde{\Psi}_{p}\rangle:=\sqrt{p}|\phi_{\pi}\rangle|\phi_{\pi}\rangle|0\rangle+\sqrt{1-p}|0\rangle|0\rangle|1\rangle)$. We define the channel ~$\tilde{\cal E}_{\lambda}$ as
~$\tilde{\cal E}_{\lambda}(\rho):=\frac{1+\lambda}{2}\rho+\frac{1-\lambda}{2}\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)\rho\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right).$ (108)
Then, we define the state with two parameters $\lambda$ and $\delta$ as
$\tilde{\rho}_{p,\lambda,\delta}:=(1-\delta)(\tilde{\cal E}_{\lambda}\otimes\tilde{\cal E}_{\lambda}\otimes Id)(|\tilde{\Psi}_{p}\rangle\langle\tilde{\Psi}_{p}|)+\delta\rho_{mix}.$ (109)
That is, we have $(V^{\otimes 2}\otimes I){\rho}_{p,\lambda,\delta}(V^{\otimes 2}\otimes I)^{\dagger}=\tilde{\rho}_{p,\lambda,\delta}$. Hence, this model is essentially written with real matrix elements.
We make numerical experiments when $n$ is set to be $16$. That is, $|{\cal D}_{n,d_{B}}|$ is $16^{2}$. By setting $p=1/2$, Fig. 3 plots the minimum relative entropy $E_{{\cal D}}^{*}({\rho}_{1/2,\lambda,\delta})$ with $\delta=10^{-1},10^{-2},10^{-4}$. When $\lambda$ is close to zero, the state ${\rho}_{p,\lambda,\delta}$ approaches to the set ${\cal S}_{AB}$ of separable states.
Figure 3: Minimum relative entropy $E_{{\cal D}}^{*}({\rho}_{1/2,\lambda,\delta})$: The vertical axis shows the minimum relative entropy. The horizontal axis shows the parameter $\lambda$. Blue (orange, green) curve presents the case with $\delta=10^{-1}$, ($10^{-2},10^{-4}$).
Next, we apply our algorithm to the membership problem of separability by choosing $p=0.1$. Figure 4 plots the boundary to satisfy the condition $E_{{\cal D}}^{*}({\rho}_{0.1,\lambda,\delta})=10^{-8}$. To compare our method with the partial transpose negative test, this figure plots the boundary to satisfy the condition that the minimum eigenvalue of the partial transposed state is $-10^{-8}$. These comparisons show that our method can detect entangled states that cannot be detected by the partial transpose negative test even when all the matrix components of the density matrix of a $4\times 2$ system are written in real numbers.
Figure 4: Comparison between our method and the partial transpose negative test: The vertical axis shows the parameter $\lambda$. The horizontal axis shows the parameter $\delta$. Blue (Orange) curve shows our method (the partial transpose negative test). The upper regions of both curves are composed of states that are guaranteed to be entangled by the respective methods.
X Exponential quantum Sanov theorem
We consider a $d$-dimensional Hilbert space ${\cal H}$ and a state $\rho$ on ${\mathcal{H}}$. We assume the $n$-fold independently and identical distributed (iid) condition and consider the following two hypotheses.
$H_{0}:$ The state is $\hbox{The state is }\rho^{\otimes n},$ (110)
$H_{1}:$ The state is $\hbox{The state is }\sigma^{\otimes n}\hbox{ with }\sigma\in S.$ (111)
To address this problem, we focus on the following quantity.
$\beta_{\epsilon,n}(\rho):=\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho^{\otimes n}\Big{|}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}(I-T)\sigma^{\otimes n}\leq\epsilon\Big{\}}.$ (112)
The relation
$\lim_{n\to\infty}\frac{1}{n}-\log\beta_{\epsilon,n}(\rho)=E^{*}(\rho):=\min_{\sigma\in S}D(\sigma\|\rho).$ (113)
is known as quantum Sanov theorem [11, 12]. To get an exponential evaluation, as the extension of $E^{*}(\rho)$, we define the minimum sandwich relative entropy as
$E_{\alpha}^{*}(\rho):=\min_{\sigma\in S}D_{\alpha}(\sigma\|\rho).$ (114)
As an exponential version of quantum Sanov theorem, the following theorem holds.
Theorem 11
We have
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$ $\geq\sup_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}$ (115)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)$ $\leq\inf_{\alpha\geq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (116)
More precisely, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$
$\geq$ $-\frac{(d+2)(d-1)}{n}\log(n+1)$
$+\max_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (117)
Our bound is the optimal when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. Although the paper [12] derived exponential evaluation for both error probabilities, its exponent does not equal to the optimal bound even when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. In this sense, our evaluation is better than the upper bound by [12].
The proof is given in Appendix D. Theorem 11 proves Theorem 1. Under the limit $r\to 0$, the RHS of (6) goes to $E_{1}^{*}(\rho)=E^{*}(\rho)$. Hence, the relation (6) recovers the inequality $\geq$ in (4) while the opposite inequality in (4) follows from the converse part of quantum Stein’s lemma.
XI Discussion
We have proposed new entanglement measures based on the distinguishability of the given entangled state from the set of separable states with two kinds of formulations by using quantum relative entropy and sandwich Rényi relative entropy. That is, we have proposed two kinds of measures based on quantum relative entropy. This measure is different from the entanglement of relative entropy [8, 9] because the choice of two input states are opposite to each other.
In the first formulation, the possible separable states are limited to tensor product states in this discrimination problem. This case can be considered as a special case of quantum Sanov theorem [11, 12]. Our analysis realizes an exponential evaluation for both kinds of error probabilities, which is better than the existing results [12]. In the second formulation, the possible separable states are not limited to tensor product states. This problem setting cannot be considered as a special case of quantum Sanov theorem. Even in this setting, we have derived an exponential evaluation for both kinds of error probabilities. Based on these two formulations, we have proposed two kinds of entanglement measures. The former measure is not smaller than the latter measure.
Also, the proposed quantity is useful for finding entanglement witness based on its information-geometrical structure. In the next step, we have derived several useful calculation formulas for these measures. In particular, we have shown that these two kinds of measures coincide with each other when the entangled state is maximally correlated. In particular, the entanglement measure based on the quantum relative entropy has been concretely derived. However, it is not easy to calculate these measures for general entangled states. To resolve this problem, we have proposed algorithms to calculate these measures. In the first step of this calculation, we have approximated these values by introducing the $\delta$-net on the second quantum system, which was introduced to derive an entanglement witness [17, Section 3]. Then, we apply the quantum Arimoto-Blahut algorithm [18] to these approximating values. Then, we have obtained an efficient algorithm for these approximating values. These algorithms are shown to converge to the global minimum, and their errors are evaluated in a simple form.
Finally, we have applied our algorithm to the membership problem of separability. Although it is known to be NP-hard when the dimensions of both Hilbert spaces increase [14, 15, 16, 17], the calculation complexity is not so large when the dimension of the second Hilbert space is fixed [21]. In this case, our algorithm has a smaller calculation complexity than the algorithm by [21]. Also, we have applied our algorithm to find an entanglement witness.
As a future problem, we can consider the following algorithm. We choose a $\delta$ net on the second system ${\mathcal{H}}_{B}$, and apply our algorithm. Then, by fixing the obtained states on the system ${\mathcal{H}}_{A}$, our algorithm exchanges the roles of ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ so that we derive states on ${\mathcal{H}}_{B}$. In this way, we iteratively improve the states on ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$. We might expect that this algorithm works well even when the $\delta$ is not so small. However, this algorithm solves the membership problem of separability, which is known to be NP-hard when the dimensions of both Hilbert spaces increase. Hence, it is impossible to choose so small $\delta$ that the calculation time complexity is polynomial for the local dimension. It is an interesting problem whether such an iterative algorithm can improve our proposed algorithm.
Acknowledgement
MH was supported in part by the National Natural Science Foundation of China under Grant 62171212. MH is thankful to Professor Hiroshi Hirai for helpful discussions on the membership problem for separability and the importance of separation hyperplane, and informing the references [15, 20, 25, 21, 23, 24]. MH is thankful to Professor Marco Tomamichel for informing the reference [22]. MH and YI are grateful for Professor Masaki Owari to a helpful discussion on membership problem of separability and informing the reference [17].
Appendix A Construction of discrete subset
A.1 Real discrete subset
First, similar to [29, Section 5.4], we construct the real discrete subset ${\cal D}_{n,d,R}\subset{\cal S}_{d}$, where ${\cal S}_{d}$ is the $d$-dimensional sphare in $\mathbb{R}^{d+1}$. We define ${\cal D}_{n,1,R}$ as
${\cal D}_{n,1,R}:=\Big{\{}\Big{(}\cos\frac{2\pi j}{n},\sin\frac{2\pi j}{n}\Big{)}\Big{\}}_{j=0}^{n-1}.$ (118)
Then, we inductively define ${\cal D}_{n,d,R}$ as
${\cal D}_{n,d,R}:=\Big{\{}\Big{(}\cos\frac{2\pi j}{n}v,\sin\frac{2\pi j}{n}\Big{)}\Big{|}v\in{\cal D}_{n,d-1,R},j=0,\ldots,n-1\Big{\}}.$ (119)
We have
$F({\cal D}_{n,1,R}):=$ $\min_{\psi_{1}\in{\cal S}_{2}}\max_{\psi_{2}\in{\cal D}_{n,2,R}}|\langle\psi_{1}|\psi_{2}\rangle|=\cos\frac{\pi}{n}.$ (120)
In general, we have
$F({\cal D}_{n,d,R})=$ $\min_{\psi_{1}\in{\cal S}_{d}}\max_{\psi_{2}\in{\cal D}_{n,d,R}}|\langle\psi_{1}|\psi_{2}\rangle|=\cos^{d}\frac{\pi}{n}.$ (121)
This can be inductively shown as follows.
$F({\cal D}_{n,d})=\min_{\psi_{1}\in{\cal S}_{d}}\max_{\psi_{2}\in{\cal D}_{n,d}}|\langle\psi_{1}|\psi_{2}\rangle|$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{j=0,\ldots,n-1}\max_{\psi_{2}\in{\cal D}_{n,d-1}}\Big{|}\langle\psi_{1}|\psi_{2}\rangle\cos\theta\cos\frac{2\pi j}{n}+\sin\theta\sin\frac{2\pi j}{n}\Big{|}$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{j=0,\ldots,n-1}\max_{\psi_{2}\in{\cal D}_{n,d-1}}\Big{|}\langle\psi_{1}|\psi_{2}\rangle\Big{(}\cos\theta\cos\frac{2\pi j}{n}+\frac{1}{\langle\psi_{1}|\psi_{2}\rangle}\sin\theta\sin\frac{2\pi j}{n}\Big{)}\Big{|}$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|\Big{(}\max_{j=0,\ldots,n-1}\Big{|}\cos\theta\cos\frac{2\pi j}{n}+\frac{1}{\langle\psi_{1}|\psi_{2}\rangle}\sin\theta\sin\frac{2\pi j}{n}\Big{|}\Big{)}$
$\geq$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|$
$\cdot\Big{(}\max_{j=0,\ldots,n-1}\max\Big{(}\Big{|}\cos\theta\cos\frac{2\pi j}{n}+\sin\theta\sin\frac{2\pi j}{n}\Big{|},\Big{|}\cos\theta\cos\frac{2\pi j}{n}-\sin\theta\sin\frac{2\pi j}{n}\Big{|}\Big{)}\Big{)}$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|\Big{(}\max_{j=0,\ldots,n-1}\max\Big{(}\Big{|}\sin(\frac{2\pi j}{n}+\theta)\Big{|},\Big{|}\sin(\frac{2\pi j}{n}-\theta)\Big{|}\Big{)}\Big{)}$
$=$ $\Big{(}\min_{\theta}\max_{j=0,\ldots,n-1}\max\Big{(}\Big{|}\sin(\frac{2\pi j}{n}+\theta)\Big{|},\Big{|}\sin(\frac{2\pi j}{n}-\theta)\Big{|}\Big{)}\Big{(}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|\Big{)}$
$\geq$ $\cos\frac{\pi}{n}F({\cal D}_{n,d-1})=\cos^{d}\frac{\pi}{n}.$ (122)
A.2 Complex discrete subset
Next, we construct the complex dscrete subset ${\cal D}_{n,d}\subset\mathbb{C}^{d}$, where all elements of ${\cal D}_{n,d}$ are normalized vectors. Using ${\cal D}_{n,2d-2,R}$, we define ${\cal D}_{n,d}$ as
[…]
$x=(x_{1},\ldots,x_{2d-1})\in{\cal D}_{n,2d-2,R}\}.$ (123)
The number of elements $|{\cal D}_{n,d}|$ is $n^{2d-2}$.
Then, we have
$\frac{1}{4}\epsilon_{2}({\cal D}_{n,d})^{2}=1-\min_{\psi_{1}\in\mathbb{C}_{n}^{d}}\max_{\psi_{2}\in{\cal D}_{n,d}}|\langle\psi_{1}|\psi_{2}\rangle|^{2}$
$\leq$ $1-(\min_{\psi_{1}\in\mathbb{C}_{n}^{d}}\max_{\psi_{2}\in{\cal D}_{n,d}}\Re|\langle\psi_{1}|\psi_{2}\rangle|)^{2}$
$=$ $1-(\min_{\psi_{1}\in S_{2d-2}}\max_{\psi_{2}\in{\cal D}_{n,2d-2,R}}\langle\psi_{1}|\psi_{2}\rangle)^{2}$
$=$ $1-(\min_{\psi_{1}\in S_{2d-2}}\max_{\psi_{2}\in{\cal D}_{n,2d-2,R}}\langle\psi_{1}|\psi_{2}\rangle)^{2}$
$\leq$ $1-\cos^{2d}\frac{\pi}{n}=1-(1-\sin^{2}\frac{\pi}{n})^{d}\leq d\sin^{2}\frac{\pi}{n}\leq d\frac{\pi^{2}}{n^{2}}.$ (124)
When we choose $n$ as $4d\frac{\pi^{2}}{n^{2}}=\epsilon_{2}^{2}$, i.e., $n=\sqrt{\frac{4d\pi^{2}}{\epsilon_{2}^{2}}}$, the number of elements $|{\cal D}_{n,d}|$ is $(\frac{2d^{1/2}\pi}{\epsilon_{2}})^{d-1}$.
Appendix B Proof of Lemma 7
Since the relation (52) follows from (53) with taking the limit $\alpha\to 1$, we show only (53). Define the projection $P:=\sum_{j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|$ and the TP-CP map
$\Lambda^{*}(\rho^{\prime}):=\sum_{j}P\rho^{\prime}P+(I-P)\rho^{\prime}(I-P).$ (125)
We choose a product pure state $|v_{A}\rangle|v_{B}\rangle$. The relation $\|(I-P)|v_{A}\rangle|v_{B}\rangle\|=0$ holds if and only if the relations
|$\begin{split}&|\{j:|\langle u_{j,A}|v_{A}\rangle|\neq 0\}|=1,\\ &\{j:|\langle u_{j,A}|v_{A}\rangle|\neq 0\}=\{j:|\langle u_{j,B}|v_{B}\rangle|\neq 0\}\end{split}$ (126)
hold. When the above condition holds, the relation $P|v_{A}\rangle|v_{B}\rangle=|u_{j,A}\rangle|u_{j,B}\rangle$ holds for $j\in\{j:|\langle u_{j,A}|v_{A}\rangle|\neq 0\}\cap\{j:|\langle u_{j,B}|v_{B}\rangle|\neq 0\}$.
Also, we have $P|v_{A}\rangle|u_{j,B}\rangle=(\langle u_{j,A}|v_{A}\rangle)|u_{j,A}\rangle|u_{j,B}\rangle$. Hence, we have
$P(\sigma_{A}\otimes I_{B})P=\sum_{j}P(\sigma_{A}\otimes|u_{j,B}\rangle\langle u_{j,B}|)P$
$=$ $\sum_{j}\langle u_{j,A}|\sigma_{A}|u_{j,A}\rangle)|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|.$ (127)
Therefore,
$C_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}$
$=$ $\min_{p}D_{\alpha}\Big{(}\sum_{j}p_{j}|j\rangle\langle j|\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{j,B}|\Big{)}$
$\geq$ $E_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{j,B}|\Big{)}.$ (128)
Now, we fix $\alpha\in(0,1]$. For any separable state $\sigma_{AB}\in{\cal S}_{AB}$, we have
$\sigma_{AB}=\sum_{j}p_{j}\rho_{A,j}\otimes\rho_{B,j}\leq\sum_{j}p_{j}\rho_{A,j}\otimes I_{B}=\sigma_{A}\otimes I_{B},$ (129)
where $\sigma_{A}:=\operatorname{\mathrm{Tr}}_{B}\sigma_{AB}$. Hence, we have
$\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\leq\rho_{AB}^{\frac{1-\alpha}{2\alpha}}(\sigma_{A}\otimes I_{B})\rho_{AB}^{\frac{1-\alpha}{2\alpha}}.$ (130)
Since $x\mapsto x^{\alpha}$ is matrix monotone, we have
$\operatorname{\mathrm{Tr}}(\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho_{AB}^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leq\operatorname{\mathrm{Tr}}(\rho_{AB}^{\frac{1-\alpha}{2\alpha}}(\sigma_{A}\otimes I_{B})\rho_{AB}^{\frac{1-\alpha}{2\alpha}})^{\alpha}.$ (131)
Thus,
$E_{\alpha}^{*}(\rho_{AB})\geq\min_{\sigma_{A}}D_{\alpha}(\sigma_{A}\otimes I_{B}\|\rho_{AB}).$ (132)
We have
$D_{\alpha}\Big{(}\sigma_{A}\otimes I_{B}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$\geq$ $D_{\alpha}\Big{(}\Lambda^{*}(\sigma_{A}\otimes I_{B})\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $D_{\alpha}\Big{(}P(\sigma_{A}\otimes I_{B})P\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $D_{\alpha}\Big{(}\sum_{j}\langle u_{j,A}|\sigma_{A}|u_{j,A}\rangle)|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|$
$\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $D_{\alpha}\Big{(}\sum_{j}\langle u_{j,A}|\sigma_{A}|u_{j,A}\rangle)|j\rangle\langle j|\Big{\|}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}.$ (133)
Taking the minimum for $\sigma_{A}$, we have
$\min_{\sigma_{A}}D_{\alpha}\Big{(}\sigma_{A}\otimes I_{B}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$\geq$ $C_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}.$ (134)
The combination of (128), (132), and (134) yields (53) with $\alpha\in(0,1]$.
Next, we fix $\alpha>1$. When a separable state $\sigma_{AB}\in{\cal S}_{AB}$ satisfies $(I-P)\sigma_{AB}(I-P)\neq 0$, $D_{\alpha}\Big{(}\sigma_{AB}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=\infty$. Hence, we have
$E_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $\min\Big{\{}D_{\alpha}\Big{(}\sigma_{AB}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$\hskip 43.05542pt\Big{|}\sigma_{AB}\in{\cal S}_{AB},(I-P)\sigma_{AB}(I-P)=0\Big{\}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\min_{p}D_{\alpha}\Big{(}\sum_{j}p_{j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|$
$\hskip 43.05542pt\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}.$ (135)
The relation $(a)$ follows from the condition (126). The combination of (128) and (135) yields (53) with $\alpha>1$.
Appendix C Proof of Theorems 8 and 9
C.1 Generalized Arimoto-Blahut algorithm
Extending Arimoto-Blahut algorithm [53, 54], the paper [18] proposed a general algorithm. Recently, this general algorithm was generalized to another form [55]. We consider a finite-dimensional Hilbert space ${\cal H}$ and focus on the set ${\cal S}({\cal H})$ of density matrices on ${\cal H}$. In this section, as a preparation, we consider a minimization problem over the set ${\cal S}({\cal H})$. Given a continuous function $\Omega$ from the set ${\cal S}({\cal H})$ to the set of Hermitian matrices on ${\cal H}$, $B({\cal H})$, we consider the minimization $\min_{\rho\in{\cal S}({\cal H})}{\cal G}(\rho)$;
${\cal G}(\rho):=\operatorname{\mathrm{Tr}}\rho\Omega[\rho].$ (136)
In the following, we discuss the calculation of the following two problems;
$\overline{{\cal G}}(a):=\min_{\rho\in{\cal S}({\cal H})}{\cal G}(\rho),\quad\rho_{*}:=\mathop{\rm argmin}\limits_{\rho\in{\cal S}({\cal H})}{\cal G}(\rho).$ (137)
For this aim, with $\gamma>0$, we define the density matrix ${\cal F}[\sigma]$ as ${\cal F}[\sigma]:=\frac{1}{\kappa[\sigma]}\exp(\log\sigma-\frac{1}{\gamma}\Omega[\sigma])$, where $\kappa[\sigma]$ is the normalization factor $\operatorname{\mathrm{Tr}}\exp(\log\sigma-\frac{1}{\gamma}\Omega[\sigma])$. The paper [18] proposed Algorithm 3.
Algorithm 3 Minimization of ${\cal G}(\rho)$
1:Choose the initial value $\rho^{(1)}\in\mathcal{M}$;
2:repeat
3:     Calculate $\rho^{(t+1)}:={\cal F}[\rho^{(t)}]$;
4:until convergence.
For any two densities $\rho$ and $\sigma$, we define
$D_{\Omega}(\rho\|\sigma):=\operatorname{\mathrm{Tr}}\rho(\Omega[\rho]-\Omega[\sigma]).$ (138)
The paper [18, Theorem 3.3] showed the following theorem, which discusses the convergence to the global minimum and the convergence speed.
Theorem 12
When any two densities $\rho^{(t)}$ and $\rho^{(t+1)}$ in satisfy the condition
$D_{\Omega}(\rho^{(t+1)}\|\rho^{(t)})\leq\gamma D(\rho^{(t+1)}\|\rho^{(t)}),$ (139)
and a state $\rho_{*}$ satisfies
$0\leq D_{\Omega}(\rho_{*}\|\rho)$ (140)
with any state $\rho$, Algorithm 3 satisfies the condition
${\cal G}(\rho^{(t_{0}+1)})-{\cal G}(\rho_{*})\leq\frac{\gamma D(\rho_{*}\|\rho^{(1)})}{t_{0}}$ (141)
with any initial state $\rho^{(1)}$.
C.2 Proof of Theorem 8
We choose a classical-quantum state $\rho_{AC}$ as
$\rho_{AC}:=\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes X_{A|c}.$ (142)
We choose the function $\Omega_{R,{\cal D}}$ as
$\Omega_{{\cal D}}[\rho_{AC}]:=\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (143)
Then, we have
$\operatorname{\mathrm{Tr}}\rho_{AC}\Omega_{\cal D}[\rho_{AC}]=\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AC}(\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}})\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{A}\sum_{c\in{\cal C}}X_{A|c}\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}=D\Big{(}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (144)
That is, we can apply Algorithm 3. In this case, when we set $\sigma_{AC}[\{X_{A|c}\}_{c}]$ as
$\sigma_{AC}[\{X_{A|c}\}_{c}]:=\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes X_{A|c},$ (145)
${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]$ is written as
${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]=$ $\frac{1}{\kappa[\{X_{A|c}\}_{c}]}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\exp(X_{A|c}-V_{c}[\{X_{A|c^{\prime}}\}_{c^{\prime}}]).$ (146)
Algorithm 1 is the same as the special case of Algorithm 3 with the above choice and $\gamma=1$.
Also, we have
$D_{\Omega_{{\cal D}}}(\rho_{AC}\|\rho_{AC}^{\prime})$
$=$ $\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}(\sum_{c\in{\cal C}}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*})\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}$
$=$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}$ (147)
$\leq$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}$
$=$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}}^{\prime}\Big{)}=D(\rho_{AC}\|\rho_{AC}^{\prime}).$ (148)
Also, (147) guarantees that the relation $D_{\Omega_{R,{\cal D}}}(\rho_{AC}\|\rho_{AC}^{\prime})\geq 0$ holds. Hence, the condition of Theorem 12 holds. That is, Theorem 8 holds.
C.3 Proof of Theorem 9
We choose the function $\Omega_{\alpha,{\cal D}}$ as
$\Omega_{\alpha,{\cal D}}[\rho_{AC}]:=\mathop{\rm sgn}(\alpha-1)\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (149)
Then, we have
$\operatorname{\mathrm{Tr}}\rho_{AC}\Omega_{\alpha,\cal D}[\rho_{AC}]$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}_{AC}(\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}})\sum_{c\in{\cal C}}|c\rangle\langle c|$
$\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}(\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*})\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}\Big{)}.$ (150)
That is, we can apply Algorithm 3. In this case, when we set $\sigma_{AC}[\{X_{A|c}\}_{c}]$ as (145), ${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]$ is written as $\frac{1}{\kappa[\{X_{A|c}\}_{c}]}\exp(X_{A|c}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}])$. Algorithm 2 is the same as the special case of Algorithm 3 with the above choice.
We denote the minimizer by $\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]$. For $\alpha\in[0,1]$, We choose $Z_{*}=\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$ and $Z=\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$. Since $\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]$ is the minimizer, we have
$\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}\geq\operatorname{\mathrm{Tr}}Z^{\alpha}.$ (151)
Since $Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})$ and $(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})$ are density matrices, we have
$\operatorname{\mathrm{Tr}}(Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}\geq 1.$ (152)
Thus, we have
$D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])$
$=$ $-\operatorname{\mathrm{Tr}}_{AC}\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$-\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $-\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Bigg{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
$-\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\Bigg{)}$
$=$ $\operatorname{\mathrm{Tr}}Z_{*}Z^{\alpha-1}-\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}$
$=$ $(\operatorname{\mathrm{Tr}}Z_{*}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})+\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}(\operatorname{\mathrm{Tr}}(Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1)$
$\geq$ $0$ (153)
Hence, the condition (140) holds.
We choose $Z^{\prime}=\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{\prime}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$. Then, we have
$D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])$ (154)
$=$ $(\operatorname{\mathrm{Tr}}Z^{\prime}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})+\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha}(\operatorname{\mathrm{Tr}}({Z^{\prime}}^{\alpha}/\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1).$ (155)
When $Z^{\prime}$ is close to $Z$, $\operatorname{\mathrm{Tr}}({Z^{\prime}}^{\alpha}/\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1=O(\|Z^{\prime}-Z\|^{2})$ and $(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}}=O(\|Z^{\prime}-Z\|)$. However, when ${Z^{\prime}}$ and $Z$ are close to $X_{*}$, the first order derivative of $\operatorname{\mathrm{Tr}}{Z}^{\alpha}$ is zero. Hence, $(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}}=O(\|Z^{\prime}-Z\|^{2})$. Since $D(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])=O(\|Z^{\prime}-Z\|^{2})$, the value $\frac{D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])}{D(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])}$ takes a finite value when $\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]$ and $\sigma_{AC}[\{X_{A|c}\}_{c}]$ are close to $\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]$. Since $\{\sigma_{AC}[\{X_{A|c}^{(t)}\}_{c}]\}$ converges to an element, $\frac{D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{(t+1)}\}_{c}]\|\sigma_{AC}[\{X_{A|c}^{(t)}\}_{c}])}{D(\sigma_{AC}[\{X_{A|c}^{(t+1)}\}_{c}]\|\sigma_{AC}[\{X_{A|c}^{(t)}\}_{c}])}$ can be bounded by a finite value. That is, when $\gamma>0$ is taken to be a sufficiently large value, the condition (139) holds.
Appendix D Proof of Theorem 11
D.1 Proof of (115)
Similar to the papers [56, 57, 12], we employ Schur duality of ${\mathcal{H}}^{\otimes n}$ as
${\mathcal{H}}^{\otimes n}=\bigoplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}\otimes{\cal V}_{\lambda},$ (156)
${\cal U}_{\lambda}$ expresses the irreducible space of $\mathop{\rm SU}(d)$ and ${\cal V}_{\lambda}$ expresses the irreducible space of the representation $\pi$ of the permutation group $\mathfrak{S}_{n}$. Then, for any state $\rho$, we have
$\rho^{\otimes n}=\bigoplus_{\lambda\in Y_{d}^{n}}\rho_{\lambda}\otimes\rho_{\lambda,mix},$ (157)
where $\rho_{\lambda,mix}$ is the completely mixed state on ${\cal V}_{\lambda}$. Since the system ${\cal V}_{\lambda}$ has no information, it is sufficient to handle the states $\oplus_{\lambda\in Y_{d}^{n}}\rho_{\lambda},\oplus_{\lambda\in Y_{d}^{n}}\sigma_{\lambda}$ on $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$. Here, we define the map $\kappa_{n}$ as
$\kappa_{n}(\rho^{\prime}):=\sum_{\lambda\in Y_{d}^{n}}\operatorname{\mathrm{Tr}}_{{\cal V}_{\lambda}}P({\cal U}_{\lambda}\otimes{\cal V}_{\lambda})\rho^{\prime}P({\cal U}_{\lambda}\otimes{\cal V}_{\lambda})$ (158)
for a general state on ${\mathcal{H}}^{\otimes n}$. We choose a basis $\{|u_{\lambda,j}\rangle\}$ of ${\cal U}_{\lambda}$ such that $|u_{\lambda,j}\rangle$ is a eigenvector of $\rho_{\lambda}$. We define the pinching map $\Gamma_{\rho,n}$ as
$\Gamma_{\rho,n}(\rho^{\prime})$
::=:= $\sum_{\lambda\in Y_{d}^{n},j}(|u_{\lambda,j}\rangle\langle u_{\lambda,j}|\otimes P({\cal V}_{\lambda}))\rho^{\prime}(|u_{\lambda,j}\rangle\langle u_{\lambda,j}|\otimes P({\cal V}_{\lambda}))$ (159)
for a general state $\rho^{\prime}$ on ${\mathcal{H}}^{\otimes n}$. Due to the above mentioned structure, states $\Gamma_{\rho,n}(\sigma^{\otimes n})$ and $\rho^{\otimes n}$ are commutative with each other. We denote the dimension of $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$ by $d_{n}$. Since $\dim{\cal U}_{\lambda}\leq(n+1)^{d(d-1)/2}$ and $|Y_{d}^{n}|\leq(n+1)^{d-1}$ [51, (6.16) and (6.18)], we have
$d_{n}\leq(n+1)^{(d+2)(d-1)/2}.$ (160)
We have the pinching inequality:
$\rho^{\prime}\leq d_{n}\Gamma_{\rho,n}(\rho^{\prime}).$ (161)
We choose
$T_{n}(\sigma,R):=\{\Gamma_{\rho,n}(\sigma^{\otimes n})\geq e^{nR}\rho^{\otimes n}\}.$ (162)
Given $\alpha\in[0,1]$, we choose $R_{n,\alpha,r}:=\frac{-\max_{\sigma\in S}\log\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}-nr}{n(1-\alpha)}$ so that $-nr=\max_{\sigma\in S}\log\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}+(1-\alpha)nR_{n,\alpha,r}$. Here, we address several projections that are commutative with each other. For two commutative projections $P_{1}$ and $P_{2}$, we define
$P_{1}\cup P_{2}:=I-(I-P_{1})(I-P_{2}).$ (163)
Since it is natural to define $(I-P_{1})\cap(I-P_{2}):=(I-P_{1})(I-P_{2})$, the above definition for $P_{1}\cup P_{2}$ is also natural. When $\{|u_{j}\rangle\}_{j}$ is a common eigenvector system of $P_{1}$ and $P_{2}$, they can be written as $P_{k}=\sum_{j\in J_{k}}|u_{j}\rangle\langle u_{j}|$. Then, we have $P_{1}\cup P_{2}=\sum_{j\in J_{1}\cup J_{2}}|u_{j}\rangle\langle u_{j}|$. Based in this definition, we define the operator $T_{n,\alpha,r}$ as
$T_{n,\alpha,r}:=\bigcup_{\sigma\in S}T_{n}(\sigma,R_{n,\alpha,r}).$ (164)
Although $T_{n}(\sigma,R_{n,\alpha,r})$ is an operator on ${\mathcal{H}}^{\otimes n}$, it can be written by using a projection operator $\oplus_{\lambda\in Y_{d}^{n}}P_{\sigma,n,\lambda}$ on $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$ as
$T_{n}(\sigma,R_{n,\alpha,r})=\bigoplus_{\lambda\in Y_{d}^{n}}P_{\sigma,n,\lambda}\otimes P({\cal V}_{\lambda}).$ (165)
Hence, we have
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}(\bigcup_{\sigma\in S}P_{\sigma,n,\lambda})\otimes P({\cal V}_{\lambda}).$ (166)
Since all $P_{\sigma,n,\lambda}$ are commutative with each other, we can choose $\dim{\cal U}_{\lambda}$ elements $\sigma_{\lambda,j}$ in $S$ such that
$\bigcup_{\sigma\in S}P_{\sigma,n,\lambda}=\bigcup_{j}P_{\sigma_{\lambda,j},n,\lambda}.$ (167)
Therefore, we have
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}(\bigcup_{j}P_{\sigma_{\lambda,j},n,\lambda})\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j}\bigoplus_{\lambda\in Y_{d}^{n}}(P_{\sigma_{\lambda^{\prime},j},n,\lambda})\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r}).$ (168)
We have
$\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}\leq\operatorname{\mathrm{Tr}}\rho^{\otimes n}\sum_{\lambda^{\prime}\in Y_{d}^{n},j}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$\leq$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j}\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$\leq$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$\leq$ $d_{n}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$=$ $d_{n}\big{(}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}\big{)}$
$\cdot\big{(}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{nr}\big{)}^{\frac{\alpha}{1-\alpha}}$
$=$ $d_{n}(\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha})^{\frac{1}{1-\alpha}}e^{n\frac{\alpha r}{1-\alpha}}.$ (169)
For $\sigma\in S$, we have
$\operatorname{\mathrm{Tr}}\sigma^{\otimes n}(I-T_{n,\alpha,r})=\operatorname{\mathrm{Tr}}\sigma^{\otimes n}\Gamma_{\rho,n}(I-T_{n,\alpha,r})$
$=$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})(I-T_{n,\alpha,r})$
$\leq$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})(I-T_{n}(\sigma,R_{n,\alpha,r}))$
$\leq$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{n(1-\alpha)R_{n,\alpha,r}}$
$\leq$ $\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{n(1-\alpha)R_{n,\alpha,r}}=e^{-nr}.$ (170)
Thus, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)\geq-\frac{1}{n}\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}.$ (171)
In fact, applying the matrix monotone function to (161), we have
$d_{n}^{\alpha-1}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\rho,n}(\sigma^{\otimes n})(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $\Big{(}d_{n}\Gamma_{\rho,n}\Big{(}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
$\leq$ $\Big{(}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$ (172)
Thus,
$\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}$
$=$ $\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\rho,n}(\sigma^{\otimes n})(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha}$
$=$ $\operatorname{\mathrm{Tr}}\big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\big{)}$
$\cdot\big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}\big{)}$
$=$ $\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}$
$\leq$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $d_{n}^{1-\alpha}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n}.$ (173)
Combining (169) and (173), we have
$\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}\leq d_{n}^{2}(\max_{\sigma\in S}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{n}{1-\alpha}}e^{n\frac{\alpha r}{1-\alpha}}$
$=$ $d_{n}^{2}\exp\Big{(}n\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}\Big{)}.$ (174)
Since arbitrary $\alpha\in[0,1]$ satisfies (171), (174) yields
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$
$\geq$ $-\frac{1}{n}\log d_{n}^{2}+\max_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (175)
Hence, using (160), we obtain (117) and (115).
D.2 Proof of (116)
Here, we employ the idea invented in [58], whose detail is explained in [42, Section 3.8]. We choose a test $T_{n}$ such that
$\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n}\geq e^{-nr}$ (176)
for $\sigma\in S$. For $\alpha\geq 1$ and $\sigma\in S$, the information processing inequality implies
$D_{\alpha}(\sigma^{\otimes n}\|\rho^{\otimes n})$
$\geq$ $\frac{1}{\alpha-1}\log\Big{(}(\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n})^{\alpha}(\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n})^{1-\alpha}$
$+(\operatorname{\mathrm{Tr}}\sigma^{\otimes n}(I-T_{n}))^{\alpha}(\operatorname{\mathrm{Tr}}\rho^{\otimes n}(I-T_{n}))^{1-\alpha}\Big{)}$
$\geq$ $\frac{1}{\alpha-1}\log\Big{(}(\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n})^{\alpha}(\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n})^{1-\alpha}\Big{)}$
$=$ $\frac{\alpha}{\alpha-1}\log\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n}-\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}$
$\geq$ $-\frac{\alpha}{\alpha-1}nr-\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}.$ (177)
Thus, we have
$-\frac{1}{n}\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}\leq\frac{1}{n}D_{\alpha}(\sigma^{\otimes n}\|\rho^{\otimes n})+\frac{\alpha}{\alpha-1}r.$ (178)
Since any $T_{n}$ satisfying (176) satisfies (178), we have
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)\leq\frac{1}{n}D_{\alpha}(\sigma^{\otimes n}\|\rho^{\otimes n})+\frac{\alpha r}{\alpha-1}$
$=$ $D_{\alpha}(\sigma\|\rho)+\frac{\alpha r}{\alpha-1}.$ (179)
Since any $\sigma\in S$ satisfies (179), we have
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)\leq\min_{\sigma\in S}D_{\alpha}(\sigma\|\rho)+\frac{\alpha r}{\alpha-1}$
$=$ $E_{\alpha}^{*}(\rho)+\frac{\alpha r}{\alpha-1},$ (180)
which implies (116).
Appendix E Proof of Theorem 2
E.1 Proof of (14)
We employ notations and results in Appendix D when ${\mathcal{H}}={\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ and $\rho=\rho_{AB}$. We define $\Gamma_{\rho_{AB},n}$ and $d_{n}$ in the same way.
We define the map ${\cal T}_{n}$ as ${\cal T}_{n}(\sigma):=\sum_{g\in\mathfrak{S}_{n}}\frac{1}{|\mathfrak{S}_{n}|}\pi(g)\sigma\pi(g)^{\dagger}$ for a state on ${\mathcal{H}}^{\otimes n}$. We define the set of permutation-invariant separable states as
${\cal S}_{AB}^{n,sym}:=\{\sigma\in{\cal S}_{AB}^{n}|\pi(g)\sigma\pi(g)^{\dagger}=\sigma,\quad g\in\mathfrak{S}_{n}\}.$ (181)
Then, we have the following lemma.
Lemma 13
$\overline{\beta}_{\epsilon,n}(\rho_{AB})$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (182)
Proof: Since ${\cal S}_{AB}^{n,sym}\subset{\cal S}_{AB}^{n}$, we have
$\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}$
$\geq$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (183)
Hence, it is sufficient to the opposite inequality.
We choose the test $T_{\epsilon}$ such that any element $\sigma_{AB}^{\prime}\in{\cal S}_{AB}^{n,sym}$ satisfies
$\operatorname{\mathrm{Tr}}(I-T_{\epsilon})\sigma_{AB}^{\prime}\leq\epsilon,$ (184)
$\operatorname{\mathrm{Tr}}T_{\epsilon}\rho_{AB}^{\otimes n}$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (185)
Due to the symmetric property, any element $\sigma_{AB}^{\prime}\in{\cal S}_{AB}^{n,sym}$ satisfies
$\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\sigma_{AB}^{\prime}\leq\epsilon,$ (186)
$\operatorname{\mathrm{Tr}}{\cal T}_{n}(T_{\epsilon})\rho_{AB}^{\otimes n}$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (187)
For any element $\sigma_{AB}\in{\cal S}_{AB}^{n}$, we have $\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\sigma_{AB}=\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\pi(g)\sigma_{AB}\pi(g)^{\dagger}$. Since ${\cal T}_{n}(\sigma_{AB})\in{\cal S}_{AB}^{n,sym}$, we have $\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\sigma_{AB}=\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon})){\cal T}_{n}(\sigma_{AB})\leq\epsilon$. Thus,
$\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}$
$\leq$ $\operatorname{\mathrm{Tr}}{\cal T}_{n}(T_{\epsilon})\rho_{AB}^{\otimes n}=\operatorname{\mathrm{Tr}}T_{\epsilon}\rho_{AB}^{\otimes n}$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (188)
Since any element $\sigma\in{\cal S}_{AB}^{n,sym}$ is written as the form
$\sigma=\bigoplus_{\lambda\in Y_{d}^{n}}\rho_{\lambda}\otimes\rho_{\lambda,mix}$ (189)
in the same way as (157), replacing the set $\{\sigma^{\otimes n}|\sigma\in S\}$ of states by ${\cal S}_{AB}^{n,sym}$, given $r>0$, we define $T_{n,R}$ and $R_{n,\alpha,r}$. In the same way as (169), we have
$\operatorname{\mathrm{Tr}}\rho_{AB}^{\otimes n}T_{n,\alpha,r}$
$\leq$ $d_{n}(\max_{\sigma\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}\Gamma_{\rho_{AB},n}(\sigma)^{\alpha}(\rho_{AB}^{\otimes n})^{1-\alpha})^{\frac{1}{1-\alpha}}e^{n\frac{\alpha r}{1-\alpha}}.$ (190)
for $\alpha\in[0,1]$. For $\sigma\in{\cal S}_{AB}^{n,sym}$, in the same way as (170) we have
$\operatorname{\mathrm{Tr}}\sigma(I-T_{n,\alpha,r})\leq e^{-nr}.$ (191)
Thus, we have
$-\frac{1}{n}\log\overline{\beta}_{e^{-nr},n}(\rho_{AB})\geq-\frac{1}{n}\log\operatorname{\mathrm{Tr}}\rho_{AB}^{\otimes n}T_{n,\alpha,r}.$ (192)
In the same way as (173), we have
$\operatorname{\mathrm{Tr}}\Gamma_{\rho_{AB},n}(\sigma)^{\alpha}(\rho_{AB}^{\otimes n})^{1-\alpha}$
$\leq$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho_{AB}^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma(\rho_{AB}^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}.$ (193)
Combining (190) and (193), we have
$\operatorname{\mathrm{Tr}}\rho_{AB}^{\otimes n}T_{n,\alpha,r}\leq d_{n}^{2}\exp\Big{(}\frac{(1-\alpha){E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})-n\alpha r}{1-\alpha}\Big{)}.$ (194)
Since an arbitrary real $\alpha\in[0,1]$ satisfies (171), (194) yields
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $-\frac{1}{n}\log d_{n}^{2}+\max_{0\leq\alpha\leq 1}\frac{\frac{1}{n}(1-\alpha)E_{\alpha}^{*}(\rho_{AB}^{\otimes n})-\alpha r}{1-\alpha},$ (195)
using (160), we obtain (14) and (16) by taking the limit.
E.2 Proof of (15)
For $\alpha\geq 1$ and $\sigma\in{\cal S}_{AB}^{n}$, in the same way as (179), we have
$-\frac{1}{n}\log\overline{\beta}_{1-e^{-nr},n}(\rho)\leq\frac{1}{n}D_{\alpha}(\sigma\|\rho^{\otimes n})+\frac{\alpha r}{\alpha-1}$ (196)
for $\sigma\in{\cal S}_{AB}^{n}$. Since any $\sigma\in{\cal S}_{AB}^{n}$ satisfies (196), we have
$-\frac{1}{n}\log\overline{\beta}_{1-e^{-nr},n}(\rho)\leq$ $\min_{\sigma\in{\cal S}_{AB}^{n}}\frac{1}{n}D_{\alpha}(\sigma\|\rho^{\otimes n})+\frac{\alpha r}{\alpha-1}$
$=$ $\frac{1}{n}E_{\alpha}^{*}(\rho^{\otimes n})+\frac{\alpha r}{\alpha-1},$ (197)
which implies (17) and (15) by taking the limit.