### v3

# Entanglement measures for detectability

Abstract
We propose new entanglement measures as the detection performance based on the hypothesis testing setting. We clarify how our measures work for detecting an entangled state by extending the quantum Sanov theorem. Our analysis covers the finite-length setting. Exploiting this entanglement measure, we present how to derive an entanglement witness to detect the given entangled state by using the geometrical structure of this measure. We derive their calculation formulas for maximally correlated states, and propose their algorithms that work for general entangled states. In addition, we investigate how our algorithm works for solving the membership problem for separability. Further, employing this algorithm, we propose a method to find entanglement witness for a given entangled state.
Index Terms:
entanglement measure, entanglement detection, membership problem, hypothesis testing, entanglement witness
I Introduction
Entanglement is a valuable resource for quantum information processing. It is important to distinguish a given entangled state from separable states, i.e., non-entangled states. However, detecting an entanglement state has two meanings, the physical detection and the mathematical detection. The physical detection means the experimental detection of a generated entangled state. Given a density matrix $\rho_{AB}$ being entangled, when a physical system is prepared as the density matrix $\rho_{AB}$, we verify whether the generated physical state is entangled. This problem is a special case of quantum hypothesis testing. In the mathematical detection, given a density matrix $\rho_{AB}$, we study whether $\rho_{AB}$ is an entangled state or a separable state. Since the mathematical detection is often called the membership problem of separability, in this paper, we consider that entanglement detection means the physical one while this paper addresses these two different problems. This paper aims to propose entangled measures, and to present that our entangled measures work for both topics.
Several practical methods to achieve entanglement detection have been proposed [1, 2, 3, 4, 5, 6, 7], but they did not discuss the optimal performance for the detection. On the other hand, to characterize the amount of entanglement, many entangled measures have been proposed. However, these proposed entangled measures have less relation with the experimental detectability of entanglement. In fact, when Vedral et al [8, 9] introduced relative entropy of entanglement, they pointed out the relation between the hypothesis testing and the relative entropy. They employed the minimum relative entropy between the given entangled state and a separable state. However, this measure expresses the performance of the discrimination between the given entangled state and the closest separable state in the sense of relative entropy. When one hypothesis is chosen to be the set of all separable states, this measure does not express the performance of this discrimination [10]. Until now, no study discussed the optimal detection efficiency for entanglement verification.
In this paper, to resolve this problem, we propose a new entanglement measure that expresses the optimal performance of the hypothesis testing, i.e., discrimination between the given entangled state and the set of separable states. The reason is that quantum hypothesis testing is considered as a key topic in quantum Shannon theory because it is related to various topics in quantum Shannon theory [11, 12, 13, 14, 15, 16, 18, 19, 20]. In this characterization, in the first step, we employ the quantum Sanov theorem [21, 22] for hypothesis testing, and consider the same minimization as relative entropy of entanglement by exchanging the first and second inputs in the relative entropy. The proposed quantity is an entanglement monotone. To enhance this characterization, we derive a much simpler upper bound of the error probability in the finite-length setting of this kind of hypothesis testing, which can be considered another proof of the quantum Sanov theorem. In addition, we derive the strong converse exponent of the same problem setting as the quantum Sanov theorem. However, when we simply apply the quantum Sanov theorem, we assume that the separable states are limited to the independent and identical distributed (iid) states of separable states. To remove this assumption, in the second step, we derive the optimal performance of this discrimination when possible separable states are allowed to take the form of a general separable state.
This entanglement measure works for finding an entanglement witness to detect the target entangled state. This relation is characterized by information geometry [23]. Also, we calculate the proposed entanglement measures when a state is a maximally correlated state. This class contains all pure states. For this calculation, we introduce a notable relation with coherence measures, and several calculation formulas for coherence measures.
In the next step, we propose an efficient algorithm to compute the proposed entanglement measure on the joint system ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. It is quite difficult to directly calculate entanglement measures because the calculations of entanglement measures are directly related to the membership problem of separability, which is known as NP-hard [24, 25, 26, 27]. Similar to the paper [27, Section 3], we employ $\delta$-nets on the second system ${\mathcal{H}}_{B}$. In the following, we assume that the first system ${\mathcal{H}}_{A}$ has a larger or equal dimension. That is, to optimize the choice of a separable state, we fix as many states on ${\mathcal{H}}_{B}$ as possible, and optimize the choice of states on ${\mathcal{H}}_{A}$. Then, we define an approximate value of the proposed measure. Employing the quantum Arimoto-Blahut algorithm proposed by the paper [28], we propose an algorithm for the approximate value. This algorithm is an iterative algorithm, and we show its convergence to the global minimum and its convergence speed.
As the third step, we apply our algorithm to the membership problem of separability. Although it is known that this problem is NP-hard [24, 25, 26, 27], several algorithms have been proposed for this problem. As a typical method, Doherty et al. [29, 30] proposed a method to identify whether a given state is a separable state or not by considering “symmetric extensions” and “PPT-symmetric extensions” in the second system ${\mathcal{H}}_{B}$. This kind of extension is called DPS hierarchy. Later, Navascués et al. [31] evaluated its calculation complexity. DPS hierarchy was discussed in [32, (14)], [33, 34]. The paper [27] proposed another algorithm for this problem by using $\delta$-nets on the second system ${\mathcal{H}}_{B}$. The paper [27] also discussed its calculation complexity. It was shown that the calculation complexity of these preceding algorithms are exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. The papers [31, 27] evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in their algorithms. The paper [31, Eq. (38)] evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. This paper discussed our calculation complexity, and showed that the calculation complexity of our algorithm is exponential only for the dimension of the second system ${\mathcal{H}}_{B}$. In addition, we evaluated the exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ in our algorithm, and evaluated the polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$. As the comparison with the evaluation by [31], we find that our exponential component for the dimension of the second system ${\mathcal{H}}_{B}$ is larger than that in the analysis [31], but our polynomial component for the dimension of the first system ${\mathcal{H}}_{A}$ is smaller than that in the analysis [31]. This fact shows that our algorithm has an advantage over the algorithm based on “symmetric extensions” and “PPT-symmetric extensions” when the dimension of the second system ${\mathcal{H}}_{B}$ is fixed and only the dimension of the first system ${\mathcal{H}}_{A}$ increases.
Finally, we apply our algorithm to finding a separation hyperplane. When we solve a conic linear programming with a general cone, it is a key step to find a separation hyperplane for an element that does not belong to the cone [35, 36, 37, 38]. Recently, a conic linear programming with the separable cone has been applied to several problems in quantum information [39, 40]. Therefore, it is desired to find a separation hyperplane for the separable cone. In quantum information, a separation hyperplane for the separable cone is called an entanglement witness [41, 42, 43, 44, 45]. From the viewpoint of entanglement theory, it is strongly desired to establish a method to find an entanglement witness for a given state, but such a method has been established only for the limited cases [46, 47]. In this paper, based on our algorithm, we also propose an algorithm to find an entanglement witness, i.e., a separation hyperplane for an entangled state. Although the method based on DPS hierarchy does not directly yield an entanglement witness, our method directly directly yields it, which is an advantage of our method.
Overall, in addition to the proposal of the new entangled measures, the contributions of this paper are summarized as follows.
[Statistical hypothesis testing] We have shown that the proposed measures present the performance of physical detection in the sense of hypothesis testing, which has not been pointed out before. (Sections II and III)
[Exponential quantum Sanov theorem] To support the above fact, we have derived the exponential version of the quantum Sanov theorem [21, 22]. This extended version clarifies the relation between the hypothesis testing with composite hypothesis and the minimum sandwiched relative Rényi entropy, which was not discussed in the existing studies [21, 22]. Thanks to this improved evaluation, we could derive the performance of the hypothesis testing when $H_{0}$ hypothesis is the set of all separable states over $n$ bipartite systems, as opposed to an IID $n$-fold tensor product of a separable state. (Sections X and III)
[Maximally correlated states] We have calculated the proposed measures when the entangled state is maximally correlated. For this calculation, we have calculated the similar value for the coherence. (Sections V and VI)
[Algorithm & membership problem] We have shown that the proposed measures can be calculated by using the generalized Arimoto-Blahut algorithm due to the good shape of the proposed measure, which has not been pointed out before. This fact shows that the proposed measures can be used for solving the membership problem for separable states. The superiority between this method and the existing method [29, 30, 31] for the membership problem depends on the parameters of this problem. In addition, we have evaluated the relation between the calculation complexity and the precision of one of the proposed algorithms. We have made a numerical demonstration. (Sections VII, VIII, and IX)
[Finding entanglement witness] We have revealed the geometrical structure of the proposed measure by using the viewpoint of information geometry. Using this geometrical structure, we have proposed a method to find an entanglement witness by using the proposed measure. This fact shows that the calculation of the proposed measure can be used for finding an entanglement witness of a given state. Applying the algorithm stated in (A4), we have proposed an algorithm to find an entanglement witness and such a method has been strongly desired from the viewpoint of entanglement theory [46, 47]. In addition, information-geometric characterization enables us to simplify the proof of the additivity of the proposed measure while the additivity was shown by [55]. (Sections IV and VIII)
The remainder of this paper is organized as follows. Section II formulates the detection of an entangled state as a problem of hypothesis testing under the iid setting. Section III addresses this problem under the non-iid setting. Section IV explains information geometrical characterization of our proposed measure and its relation to an entanglement witness. Section V proposes and studies new coherence measures as the preparation of the calculation of our entanglement measures. Section VI derives several calculation formulas for our entanglement measures for maximally entangled states based on the analysis in Section V. Section VII proposes algorithms for our entanglement measures by using $\delta$-net. Section VIII applies our results to the membership problem for separability and separation hyperplane. Section IX numerically calculates our entanglement measure, and check how our method works for the membership problem. Section IV explains that our method finds an entanglement witness. Section X presents our exponential version of quantum Sanov theorem. Section XI makes conclusions and discussions. Appendix A presents concrete constructions of $\delta$-nets. Appendix B proves the theorem for the relation between our entanglement measures and our proposed coherence measure. Appendix C proves theorems for the convergence for our algorithms. Appendix D proves our exponential version of quantum Sanov theorem. Appendix E proves our exponential evaluation of error probabilities in our problem of hypothesis testing under the non-iid setting.
II Detecting entangled state with iid setting
We consider the joint system ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$, and discuss how efficiently an entangled state $\rho_{AB}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ is distinguished from the set ${\cal S}_{AB}$ of separable states on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$. This problem can be formulated as a hypothesis testing problem to distinguish $\rho_{AB}$ from ${\cal S}_{AB}$. For this detection, we assume the $n$-fold independently and identical distributed (iid) condition. That is, we have the following two hypotheses for the possible state as follows.
$H_{0}:$ The state is $\hbox{The state is }\rho_{AB}^{\otimes n}.$ (1)
$H_{1}:$ The state is $\hbox{The state is }\sigma_{AB}^{\otimes n}\hbox{ with }\sigma_{AB}\in{\cal S}_{AB}.$ (2)
To address this problem, we focus on the following quantity.
$\beta_{\epsilon,n}(\rho_{AB})$
::=:= $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}^{\otimes n}\leq\epsilon\Big{\}}.$ (3)
In the above setting, we do not impose the LOCC restriction to our measurement unlike [1, 43]. However, these existing studies did not discuss the optimal detection efficiency for the entanglement verification. Hence, the above quantity expresses the performance of discrimination at the fundamental level.
Here, it is better to explain the relation with the entanglement verification [48, 49, 50, 51, 52, 53]. Entanglement verification is the task of identifying a given entangled state within a certain error. Since entanglement verification is a more difficult task, the detection efficiency of the entanglement verification is worse than the detection efficiency of the entanglement detection. Existing studies for entanglement verification are limited to the verification for several special entangled states. No general theory covers the entanglement verification for general entangled states.
The application of quantum Sanov theorem [21, 22] guarantees the following
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{\epsilon,n}(\rho_{AB})=E^{*}(\rho_{AB})$
::=:= $\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\sigma_{AB}\|\rho_{AB}),$ (4)
where $D(\sigma\|\rho):=\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log\rho)$. This value is different from relative entropy of entanglement $E(\rho_{AB}):=\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\rho_{AB}\|\sigma_{AB})$. This fact shows that our measure $E^{*}(\rho_{AB})$ expresses the limit of the performance to distinguish the entangled state $\rho_{AB}$ from ${\cal S}_{AB}$. In contrast, when the choice of $H_{0}$ and $H_{1}$ is exchanged, $E(\rho_{AB})$ does not have the same meaning [10]. The proof of (4) is given as follows. The part $\leq$ in (4) follows from the strong converse part of quantum Stein’s lemma. The paper [21] showed the part $\geq$ in (4). However, they did not consider the case when $\epsilon$ is exponentially small. To enhance the meaning of the quantity $E^{*}(\rho_{AB})$, it is needed to evaluate $\beta_{\epsilon,n}(\rho_{AB})$ in such a case.
TABLE I: List of mathematical symbols for detecting entangled state from Section II onwards.
Symbol Description Eq. number
${\cal S}_{AB}$ Set of separable states on
${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$
$\beta_{\epsilon,n}(\rho_{AB})$ Minimum error probability (3)
with constraint $\epsilon$ when
$H_{1}$ is given as (2)
$E^{*}(\rho_{AB})$ Minimum relative entropy (4)
$E_{\alpha}^{*}(\rho_{AB})$ Minimum relative Rényi (5)
entropy of order $\alpha$
${\cal S}_{AB}^{n}$ Set of separable states on
${\mathcal{H}}_{A}^{\otimes n}\otimes{\mathcal{H}}_{B}^{\otimes n}$
$\overline{\beta}_{\epsilon,n}(\rho_{AB})$ Minimum error probability (3)
with constraint $\epsilon$ when
$H_{1}$ is given as (10)
$\overline{E}^{*}(\rho_{AB})$ Regularized minimum (12)
relative entropy
$\overline{E}_{\alpha}^{*}(\rho_{AB})$ Regularized minimum relative (13)
Rényi entropy of order $\alpha$
$\tilde{E}^{*}(\rho_{AB})$ The above value with (18)
$\alpha=1-0$
$\sigma_{AB}^{*}$ Closest separable state (21)
${\cal S}_{D}$ Divergence sphere (27)
${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ Separation hyperplane (22)
$|\Phi\rangle$ Maximally entangled state
$\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle|j\rangle$
$E_{{\cal D}}^{*}(\rho_{AB})$ Minimum relative entropy (66)
with fixed finite set ${\cal D}$
$\epsilon_{2}({\cal D})$ (67)
$E_{{\alpha,\cal D}}^{*}(\rho_{AB})$ Minimum relative Rényi (78)
entropy of order $\alpha$
with fixed finite set ${\cal D}$
For this aim, using the sandwiched relative Rényi entropy $D_{\alpha}(\sigma\|\rho):=\frac{1}{\alpha-1}\log\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$, we generalize $E^{*}(\rho_{AB})$ as
$E_{\alpha}^{*}(\rho_{AB}):=\min_{\sigma_{AB}\in{\cal S}_{AB}}D_{\alpha}(\sigma_{AB}\|\rho_{AB}).$ (5)
As presented in Section X, we have the following strengthened version of (4).
Theorem 1
We have the following relations
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha<1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha},$ (6)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho_{AB})$
$\leq$ $\inf_{\alpha\geq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha}.$ (7)
More precisely, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha<1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha}$
$-\frac{(d_{AB}+2)(d_{AB}-1)}{n}\log(n+1),$ (8)
where $d_{AB}$ is the dimension of ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$.
Inequality (6) presents an attainable exponential rate of the failure probability for detecting the entangled state $\rho_{AB}$ when we impose an exponential constraint for the incorrect detection of any separable state under the iid condition. Under the limit $r\to 0$, the RHS of (6) goes to $E_{1}^{*}(\rho_{AB}):=E^{*}(\rho_{AB})$. Hence, the relation (6) recovers the inequality $\geq$ in (4). Inequality (7) presents an upper bound for the exponential rate of the failure probability to detect the entangled state $\rho_{AB}$ when we relax the constraints for the incorrect detection of any separable state under the iid condition. Due to this inequality, when the entangled state $\rho_{AB}$ can detected with a failure probability of a strictly larger exponential rate than $e^{-nE^{*}(\rho_{AB})}$, a separable state is incorrectly detected as an entangled state with a high probability.
When a TP-CP map $\Lambda$ preserves the set ${\cal S}_{AB}$, these quantities satisfy the monotonicity as
$E^{*}(\rho_{AB})\geq E^{*}(\Lambda(\rho_{AB})),\quad E_{\alpha}^{*}(\rho_{AB})\geq E_{\alpha}^{*}(\Lambda(\rho_{AB})).$ (9)
Our new measure $E^{*}(\rho_{AB})$ has a completely different behavior from conventional measures of entanglement. Most of conventional measures of entanglement of a pure state $\rho_{AB}$ take the von Neumann entropy of its reduced density. However, $E^{*}(\rho_{AB})$ becomes infinity when $\rho_{AB}$ is a pure entangled state because the support of no separable state is included in the support of $\rho_{AB}$. Although the log-negativity of a pure state $\rho_{AB}$ does not equal the von Neumann entropy of its reduced density, it takes a finite value.
Remark 1
When $H_{0}$ and $H_{1}$ are exchanged, $E(\rho_{AB})$ does not have the same meaning as $E^{*}(\rho_{AB})$. This is because quantum Sanov theorem works only when $H_{1}$ is composed of several states [10]. That is, the choice of the hypothesis composed of several states is crucial in this scenario. Instead, to prove that $E(\rho_{AB})$ upper bounds the entanglement distillation rate, the paper [54] employs the simple hypotheses that $H_{0}$ is $\sigma_{AB}^{\otimes n}$ and $H_{1}$ is $\rho_{AB}^{\otimes n}$, where $\sigma_{AB}$ is an arbitrary chosen element of ${\cal S}_{AB}$. More precisely, an application of the strong converse of Stein’s lemma to this hypothesis testing yields the strong converse for the entanglement distillation rate [51, Proof of Theorem 8.7].
III Detecting entangled state with non-iid setting
The above discussion assumes the iid assumption for the alternative hypothesis. However, several existing studies for entanglement verification and entanglement detection assume that the alternative hypothesis includes general separable states including non-iid separable stets [5, 4, 51, 52, 53], which is called the non-iid setting. Our analysis can be extended to the non-iid setting as follows. We denote the set of separable states on ${\mathcal{H}}_{A}^{\otimes n}\otimes{\mathcal{H}}_{B}^{\otimes n}$ by ${\cal S}_{AB}^{n}$. Then, the alternative hypothesis is changed to
$H_{1}:$ The state is $\hbox{The state is }\sigma\hbox{ with }\sigma\in{\cal S}_{AB}^{n}$ (10)
while the null hypothesis is kept to (1). To address this problem, we focus on the following quantity.
$\overline{\beta}_{\epsilon,n}(\rho_{AB})$
::=:= $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (11)
The relation (4) yields that
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{\epsilon,n}(\rho_{AB})\leq\overline{E}^{*}(\rho_{AB}):=\inf_{n}\frac{1}{n}{E}^{*}(\rho_{AB}^{\otimes n})$
$=$ $\lim_{n\to\infty}\frac{1}{n}{E}^{*}(\rho_{AB}^{\otimes n})=\lim_{n\to\infty}\frac{1}{n}\min_{\sigma\in{\cal S}_{AB}^{n}}D(\sigma\|\rho_{AB}^{\otimes n}).$ (12)
Extending ${E}_{\alpha}^{*}(\rho_{AB})$, we define
$\overline{E}_{\alpha}^{*}(\rho_{AB}):=$ $\inf_{n}\frac{1}{n}{E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})=\lim_{n\to\infty}\frac{1}{n}{E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})$
$=$ $\lim_{n\to\infty}\frac{1}{n}\min_{\sigma_{AB}\in{\cal S}_{AB}^{n}}D_{\alpha}(\sigma_{AB}\|\rho_{AB}^{\otimes n}).$ (13)
As shown in Appendix E, we have the following theorem, which is an enhanced version of Theorem 1, which strengthens (13).
Theorem 2
We have
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha\leq 1}\frac{(1-\alpha)\overline{E}_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha},$ (14)
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{1-e^{-nr},n}(\rho_{AB})$
$=$ $\inf_{\alpha\geq 1}\frac{(1-\alpha)\overline{E}_{\alpha}^{*}(\rho_{AB})-\alpha r}{1-\alpha}.$ (15)
More precisely, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $\sup_{0\leq\alpha\leq 1}\frac{\frac{1}{n}(1-\alpha)E_{\alpha}^{*}(\rho_{AB}^{\otimes n})-\alpha r}{1-\alpha}$
$-\frac{(d_{AB}+2)(d_{AB}-1)}{n}\log(n+1),$ (16)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho_{AB})$
$\leq$ $\inf_{\alpha\geq 1}\frac{\frac{1}{n}(1-\alpha){E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})-\alpha r}{1-\alpha}.$ (17)
As an extended version of (6), Inequality (14) presents an attainable exponential rate of the failure probability for detecting the entangled state $\rho_{AB}$ when we impose an exponential constraint for the incorrect detection of any separable state without the iid condition. Under the limit $r\to 0$, the RHS of (14) goes to $\lim_{\delta\to+0}\overline{E}_{1-\delta}^{*}(\rho_{AB})$. Inequality (14) implies the following inequality.
$\lim_{n\to\infty}-\frac{1}{n}\log\overline{\beta}_{\epsilon,n}(\rho_{AB})\geq\tilde{E}^{*}(\rho_{AB}):=\lim_{\delta\to+0}\overline{E}_{1-\delta}^{*}(\rho_{AB}).$ (18)
While $\overline{E}^{*}(\rho_{AB})\geq\tilde{E}^{*}(\rho_{AB})$, it is not clear whether equality holds in general. The reference [55, Proposition 2] showed that the additivity $E^{*}(\rho_{AB,1}\otimes\rho_{AB,2})=E^{*}(\rho_{AB,1})+E^{*}(\rho_{AB,2})$, which implies ${E}^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})$ while its simple proof is given in Section IV-B.
When the quantity $E_{\alpha}^{*}(\rho_{AB})$ satisfies the additivity
$E_{\alpha}^{*}(\rho_{AB,1}\otimes\rho_{AB,2})=E_{\alpha}^{*}(\rho_{AB,1})+E_{\alpha}^{*}(\rho_{AB,2})$ (19)
for $\alpha<1$, the equations ${E}^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})=\tilde{E}^{*}(\rho_{AB})$ and ${E}_{\alpha}^{*}(\rho_{AB})=\overline{E}_{\alpha}^{*}(\rho_{AB})$ hold. Therefore, it is crucial to clarify when the above additivity holds. For example, as shown in (63) of Section VI, the maximally correlated state satisfies the above additivity. The reference [56, Section 5] showed that the additivity (19) with $\alpha\in[1/2,1)$ does not hold in general by considering the bipartite antisymmetric (Werner) state [57, 58] [^1] Since our sandwiched relative entropy of order $\alpha^{\prime}$ corresponds to $\tilde{D}_{1-\alpha^{\prime}}$ of [56, (10)], i.e., the case with $(\alpha,z)=(1-\alpha^{\prime},\alpha^{\prime})$. Hence, the case with $21>\alpha^{\prime}\geq 1/$ 2 belongs to the set ${\cal D}$ defined in [56, Section 2]. Hence, the discussion [56, Section 5] covers the violation of (19) the additivity (19) with $\alpha\in[1/2,1)$..
Inequality (15) presents an upper bound for the exponential rate of the failure probability to detect the entangled state $\rho_{AB}$ when we relax the constraints for the incorrect detection of any separable state without the iid condition.
IV Information-geometrical characterization: Finding entanglement witness
IV-A Single system
Our proposed measure is useful not only for hypothesis testing of the above hypotheses, but also for finding entanglement witness. That is, given a non-separable state $\rho_{AB}\notin{\cal S}_{AB}$, it is important to find an entanglement witness to distinguish $\rho_{AB}$ and ${\cal S}_{AB}$. To see this merit, we discuss its information-geometrical characterization, which yields entanglement witness.
The papers [46, 47] proposed methods to find an entanglement witness for special states. As another method, the references [59], [60, Section 16.3.3] proposed a method to use the tangent space of the set of separable states based on the Hilbert Schmidt norm $\|\cdot\|_{HS}$. The key point of the above method is the fact that the Hilbert-Schmidt norm forms the Euclidean geometry over the space of Hermitian matrices. The above method needs to find the closest separable state in the sense of the Hilbert Schmidt norm $\|\cdot\|_{HS}$. However, no method efficiently minimizes the Hilbert Schmidt norm $\|\cdot\|_{HS}$ between a given state and the set of separable states, i.e., efficiently finds
$\mathop{\rm argmin}\limits_{\sigma_{AB}\in{\cal S}_{AB}}\|\sigma_{AB}-\rho_{AB}\|_{HS}$ (20)
when the system dimensions are large. Hence, it is not easy to find an entanglement witness by using the Hilbert-Schmidt norm under large Hilbert spaces.
Fortunately, this section shows how the minimization $\min_{\sigma_{AB}\in{\cal S}_{AB}}D(\sigma_{AB}\|\rho_{AB})$ works for finding entanglement witness by employing the minimizer of $E^{*}(\rho_{AB})$, i.e., the state
$\sigma_{AB}^{*}:=\mathop{\rm argmin}\limits_{\sigma_{AB}\in{\cal S}_{AB}}D(\sigma_{AB}\|\rho_{AB}).$ (21)
That is, the minimizer of $E^{*}(\rho_{AB})$ produces entanglement witness. In addition, we propose an algorithm to calculate our entanglement measure $E^{*}(\rho_{AB})$ in Section VII. As explained in Section VIII-B, the combination between the algorithm for $\sigma_{AB}^{*}$ and a procedure to find an entanglement witness based on quantum relative entropy enables us to derive an algorithm to find an entanglement witness. Hence, the relation between entanglement witness and the minimizer of $E^{*}(\rho_{AB})$ is quite important.
Although quantum relative entropy is a two-input quantity, it does not form an Euclidean geometry because it is not symmetric with respect to exchanging two input states. Hence, when we employ our entanglement measure, we need to employ a geometric structure different from the Euclidean geometry.
Using the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})$, we define the hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ as (See Fig. 1.)
${\cal M}(\sigma_{AB}^{*},\rho_{AB})$
::=:= $\{\rho^{\prime}_{AB}|\operatorname{\mathrm{Tr}}\rho_{AB}^{\prime}(\log\sigma_{AB}^{*}-\log\rho_{AB})=D(\sigma_{AB}^{*}\|\rho_{AB})\}.$ (22)
Figure 1: Entanglement witness: This figure shows the relation among the entangled state $\rho_{AB}$, the set of separable states ${\cal S}_{AB}$, and the separation hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$. Any other element $\tilde{\sigma}_{AB}$ on the tangent space has larger relative entropy to $\rho_{AB}$ than the projected element ${\sigma}_{AB}^{*}$ due to the Pythagorean theorem (25).
The following theorem guarantees that ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ separates $\rho_{AB}$ from ${\cal S}_{AB}$. Hence, the operator $(\log\sigma_{AB}^{*}-\log\rho_{AB})-D(\sigma_{AB}^{*}\|\rho_{AB})$ works as an entanglement witness for $\rho_{AB}$. In fact, as discussed in [61, Eq. (4)], any entanglement witness can be constructed by a combination of local measurement because any matrix on the composite system ${\cal H}_{A}\otimes{\cal H}_{B}$ can be written as the combination of at most $(\dim{\cal H}_{A})^{2}(\dim{\cal H}_{B})^{2}$ matrices with a product form.
Theorem 3
(i) For any separable state $\sigma_{AB}\in{\cal S}_{AB}$, we have
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}-\log\rho_{AB})\geq D(\sigma_{AB}^{*}\|\rho_{AB}).$ (23)
(ii) Also, we have
$\operatorname{\mathrm{Tr}}\rho_{AB}(\log\sigma_{AB}^{*}-\log\rho_{AB})=-D(\rho_{AB}\|\sigma_{AB}^{*})<0.$ (24)
To grasp the geometrical meaning of the above fact, we employ information geometry. Given a Hermitian matrix $X$ and a state $\tilde{\rho}_{0}$, we often focus on the set :=${\cal E}:=\{\tilde{\rho}_{t}\}$, where $\tilde{\rho}_{t}:=\exp(tX+\log\tilde{\rho}_{0})/c_{t}$ with $c_{t}:=\operatorname{\mathrm{Tr}}\exp(tX+\log\tilde{\rho}_{0})$. The set =${\cal E}=\{\tilde{\rho}_{t}\}$ is called an exponential family generated by $X$. Also, the set :=${\cal M}:=\{\rho|\operatorname{\mathrm{Tr}}\rho X=\operatorname{\mathrm{Tr}}\tilde{\rho}_{0}X\}$ is called a mixture family generated by $X$.
Since the intersection ${\cal M}\cap{\cal E}$ is composed of only one element $\tilde{\rho}_{0}$, any elements $\sigma\in{\cal M}$ and $\rho\in{\cal E}$ satisfy
$D(\sigma\|\rho)=D(\sigma\|\tilde{\rho}_{0})+D(\tilde{\rho}_{0}\|\rho),$ (25)
which is called Pythagorean theorem [23], [12, Theorem 2.3].
We apply (25) to the case with $X=-(\log\sigma_{AB}^{*}-\log\rho_{AB})$ and $\tilde{\rho}_{0}=\sigma_{AB}^{*}$. Any element $\sigma_{AB}^{\prime}\in{\cal M}(\sigma_{AB}^{*},\rho_{AB})$ satisfies
$D(\sigma_{AB}^{\prime}\|\rho_{AB})=D(\sigma_{AB}^{\prime}\|\sigma_{AB}^{*})+D(\sigma_{AB}^{*}\|\rho_{AB}).$ (26)
Proof: Since (ii) is trivial, we show only (i) by contradiction. We assume that (i) does not hold. Hence, there exists a separable state $\sigma_{AB}\in{\cal S}$ that does not satisfy (23).
We define the divergence sphere
${\cal S}_{D}:=\{\rho_{AB}^{\prime}|D(\rho_{AB}^{\prime}\|\rho_{AB})=D(\sigma_{AB}^{*}\|\rho_{AB})\}.$ (27)
Since any state $\sigma_{AB}^{\prime}$ in ${\cal M}(\sigma_{AB}^{*},\rho_{AB})\setminus\{\sigma_{AB}^{*}\}$ satisfies $D(\sigma_{AB}^{\prime}\|\rho_{AB})=D(\sigma_{AB}^{\prime}\|\sigma_{AB}^{*})+D(\sigma_{AB}^{*}\|\rho_{AB})>D(\sigma_{AB}^{*}\|\rho_{AB})$ due to (25), any state $\sigma_{AB}^{\prime\prime}$ in the sphere ${\cal S}_{D}$ satisfies the condition $\operatorname{\mathrm{Tr}}\sigma_{AB}^{\prime\prime}(\log\sigma_{AB}^{*}-\log\rho_{AB})\leq D(\sigma_{AB}^{*}\|\rho_{AB})$. Hence, the hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$ is the tangent space of the sphere ${\cal S}_{D}$ at $\sigma_{AB}^{*}$.
If a separable state $\sigma_{AB}$ satisfies
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}-\log\rho_{AB})<D(\sigma_{AB}^{*}\|\rho_{AB}),$ (28)
there exists $\lambda\in(0,1)$ such that $\sigma_{AB}^{\prime}:=\lambda\sigma_{AB}+(1-\lambda)\sigma_{AB}^{*}$ is inside of the sphere ${\cal S}_{D}$ as Fig. 2. Hence,
$D(\sigma_{AB}^{\prime}\|\rho_{AB})<D(\sigma_{AB}^{*}\|\rho_{AB}).$ (29)
Since $\lambda\sigma_{AB}+(1-\lambda)\sigma_{AB}^{*}$ is also a separable state, the relation (29) contradicts with the choice of $\sigma_{AB}^{*}$.
Figure 2: Divergence sphere: This figure shows the relation among the entangled state $\rho_{AB}$, the divergence sphere ${\cal S}_{D}$, and the separation hyperplane ${\cal M}(\sigma_{AB}^{*},\rho_{AB})$.
IV-B Simple proof of additivity
Next, we consider the composite system of two systems ${\cal H}_{A,1}\otimes{\cal H}_{B,1}$ and ${\cal H}_{A,2}\otimes{\cal H}_{B,2}$. We consider the states $\rho_{1}$ and $\rho_{2}$ on the respective systems. The following proposition is known.
Propositon 4 ([55, Proposition 2])
$E^{*}(\rho_{1}\otimes\rho_{2})=E^{*}(\rho_{1})+E^{*}(\rho_{2}).$ (30)
This proposition can be easily shown by using information geometry as follows.
We denote the sets of separable states on ${\cal H}_{A,j}\otimes{\cal H}_{B,j}$ and $({\cal H}_{A,1}\otimes{\cal H}_{A,2})\otimes({\cal H}_{B,1}\otimes{\cal H}_{B,2})$ by ${\cal S}_{AB,j}$ and ${\cal S}_{AB,12}$, respectively, for $j=1,2$. We choose $\sigma_{j}^{*}:=\mathop{\rm argmin}\limits_{\sigma_{j}\in{\cal S}_{AB,j}}D(\sigma_{j}\|\rho_{j})$. Then, $X_{j}:=(\log\sigma_{j}^{*}-\log\rho_{j})-D(\sigma_{j}^{*}\|\rho_{j})$ is an entanglement witness for $\rho_{j}$. That is, due to Theorem 3, any separable state $\sigma_{j}\in{\cal S}_{AB,j}$ satisfies
$\operatorname{\mathrm{Tr}}\sigma_{j}X_{j}\geq 0.$ (31)
Lemma 5
$X_{1}\otimes I+I\otimes X_{2}$ is an entanglement witness for $\rho_{1}\otimes\rho_{2}$.
Proof: For any element $\sigma\in{\cal S}_{AB,1,2}$, we have
$\operatorname{\mathrm{Tr}}\sigma(X_{1}\otimes I+I\otimes X_{2})$
$=$ $\operatorname{\mathrm{Tr}}(\operatorname{\mathrm{Tr}}_{2}\sigma)X_{1}+(\operatorname{\mathrm{Tr}}_{1}\sigma)X_{2}\geq 0,$ (32)
where the inequality follows from (31).
Proof of Proposition 4:   Now, we consider the mixture family ${\cal M}_{12}:=\{\rho|\operatorname{\mathrm{Tr}}\rho(X_{1}\otimes I+I\otimes X_{2})=0\}$ generated by the matrix $(X_{1}\otimes I+I\otimes X_{2})$. We often focus on the exponential family ${\cal E}_{12}:=\{\rho_{12,t}\}$ generated by $X$, where $\rho_{12,t}:=\exp(t(X_{1}\otimes I+I\otimes X_{2})+\log(\rho_{1}\otimes\rho_{2}))/c_{t}$ with $c_{12,t}:=\operatorname{\mathrm{Tr}}\exp(t(X_{1}\otimes I+I\otimes X_{2})+\log(\rho_{1}\otimes\rho_{2}))$.
Since the intersection ${\cal M}_{12}\cap{\cal E}_{12}$ is composed of only one element $\rho_{12,1}=\sigma_{1}^{*}\otimes\sigma_{2}^{*}$, Eq. (25) (Pythagorean theorem [23], [12, Theorem 2.3]) guarantees that any element $\sigma\in{\cal M}_{12}$ satisfies
$D(\sigma\|\rho_{1}\otimes\rho_{2})=$ $D(\sigma\|\sigma_{1}^{*}\otimes\sigma_{2}^{*})+D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2})$
$\geq$ $D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2}).$ (33)
Since any separable state $\sigma^{\prime}\in{\cal S}_{AB,12}$ is located in the opposite side of $\rho_{1}\otimes\rho_{2}$ with respect to hyperplane ${\cal M}_{12}$, the relation (33) yields
$D(\sigma^{\prime}\|\rho_{1}\otimes\rho_{2})\geq D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2}),$ (34)
which implies
$\min_{\sigma^{\prime}\in{\cal S}_{AB,12}}D(\sigma^{\prime}\|\rho_{1}\otimes\rho_{2})=D(\sigma_{1}^{*}\otimes\sigma_{2}^{*}\|\rho_{1}\otimes\rho_{2}).$ (35)
Therefore, we obtain (30). ∎
V Coherence measure
Next, we consider how to calculate the proposed measure for maximally correlated states, which are special cases. For its preparation, we study the coherence measures defined below, because, as explained in the next section, our entanglement measures are related to these coherence measures. In particular, the analysis of maximally correlated states can be reduced to the analysis of the coherence measures. We consider a state $\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$ on a $d$-dimensional Hilbert space ${\mathcal{H}}$ spanned by $\{|j\rangle\}_{j=0}^{d-1}$, and denote the set of incoherent states by ${\cal I}_{c}$, which is written as $\{\sigma(p)\}$ with $\sigma(p):=\sum_{i}p_{i}|i\rangle\langle i|$, where $p=(p_{i})$ is a distribution. Then, we define the following coherence measures.
$C^{*}(\rho):=\min_{\sigma\in{\cal I}_{c}}D(\sigma\|\rho),\quad C_{\alpha}^{*}(\rho):=\min_{\sigma\in{\cal I}_{c}}D_{\alpha}(\sigma\|\rho).$ (36)
Although the papers [62, 63] considered the coherence measures based on Rényi relative entropy, our measures are different from their coherence measures because their measures are based on the quantities with exchanging the two input states from ours. In fact, $C^{*}(\rho)$ has a geometrical characterization similar to Section IV by replacing ${\cal S}_{AB}$ by ${\cal I}_{c}$.
For $\alpha\in(0,1)$ and a pure state $|\phi\rangle\langle\phi|$ with $|\phi\rangle=\sum_{j}e^{\theta_{j}\sqrt{-1}}p_{j}|j\rangle$, we have
$e^{(\alpha-1)C_{\alpha}^{*}(|\phi\rangle\langle\phi|)}=\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(|\phi\rangle\langle\phi|^{\frac{1-\alpha}{2\alpha}}\sigma|\phi\rangle\langle\phi|^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(|\phi\rangle\langle\phi|\sigma|\phi\rangle\langle\phi|)^{\alpha}$
$=$ $\max_{\sigma(q)\in{\cal I}_{c}}(\langle\phi|\sigma(q)|\phi\rangle)^{\alpha}=\max_{q}(\sum_{j}q_{j}p_{j})^{\alpha}=(\max_{i}p_{i})^{\alpha}.$ (37)
Hence, we have
$C_{\alpha}^{*}(|\phi\rangle\langle\phi|)=\frac{\alpha}{\alpha-1}\log(\max_{i}p_{i}).$ (38)
Also, $C^{*}(|\phi\rangle\langle\phi|)$ and $C_{\alpha}^{*}(|\phi\rangle\langle\phi|)$ take infinity value for $\alpha>1$ unless it is an incoherent state.
In the following discussion, we employ the operators $Z:=\sum_{j=0}^{d-1}\omega^{j}|j\rangle\langle j|$ with $\omega:=e^{2\pi\sqrt{-1}/d}$ and $X:=\sum_{j=0}^{d-1}|j+1\rangle\langle j|$. We define $\rho(p):=\sum_{j=0}^{d-1}p_{j}Z^{j}|+\rangle\langle+|Z^{-j}$ with $|+\rangle:=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle$. Then, we have the following lemma.
Lemma 6
Using the uniform distribution $p_{mix}$, we have
$C^{*}(\rho(p))$ $=D(p_{\mathop{\rm mix}}\|p)=-\frac{1}{d}\sum_{j=0}^{d-1}\log p_{j}-\log d$ (39)
$C_{\alpha}^{*}(\rho(p))$ $=D_{\alpha}(p_{\mathop{\rm mix}}\|p)$
$=\frac{1}{\alpha-1}\log\sum_{j=0}^{d-1}p_{j}^{1-\alpha}-\frac{\alpha}{\alpha-1}\log d,$ (40)
for $\alpha\in(0,1)\cup(1,2]$.
Proof: Since the above second equations follow from the direct calculation, the following analysis shows the first equations.
We have $X\rho(p)X^{-1}=\rho(p).$ Thus, $D(\sigma(q)\|\rho(p))=D(X^{j}\sigma(q)X^{-j}\|\rho(p))$. Hence, the joint convexity of the relative entropy implies
$D(\sigma(q)\|\rho(p))=\sum_{j=0}^{d-1}\frac{1}{d}D(X^{j}\sigma(q)X^{-j}\|\rho(p))$
$\geq$ $D(\sum_{j=0}^{d-1}\frac{1}{d}X^{j}\sigma(q)X^{-j}\|\rho(p))$
$=$ $D(\rho_{mix}\|\rho(p))=D(\rho(p_{mix})\|\rho(p))=D(p_{mix}\|p),$ (41)
where $\rho_{mix}$ is the completely mixed state. The above relations imply the first equation (39). Similarly, since $x\mapsto x^{\alpha}$ is matrix convex for $\alpha\in[1,2]$,
$e^{(\alpha-1)D_{\alpha}(\sigma(q)\|\rho(p))}=\sum_{j=0}^{d-1}\frac{1}{d}e^{(\alpha-1)D_{\alpha}(X^{j}\sigma(q)X^{j}\|\rho(p))}$
$\geq$ $e^{(\alpha-1)D_{\alpha}(\sum_{j=0}^{d-1}\frac{1}{d}X^{j}\sigma(q)X^{j}\|\rho(p))}$
$=$ $e^{(\alpha-1)D_{\alpha}(\rho_{mix}\|\rho(p))}=e^{(\alpha-1)D_{\alpha}(p_{mix}\|p)},$ (42)
which implies the first equation (40) for $\alpha\in(1,2]$. Similarly, since $x\mapsto x^{\alpha}$ is matrix concave for $\alpha\in[0,1)$, an analysis similar to (42) yields the first equation (40) for $\alpha\in[0,1)$.
To calculate $C^{*}(\rho)$ for a general state, we define the state $I_{c}(\rho):=\frac{1}{t(\rho)}\exp((\log\rho)^{\rm diag})$ with $t(\rho):=\operatorname{\mathrm{Tr}}\exp((\log\rho)^{\rm diag})$ and $(\log\rho)^{\rm diag}$ is the diagonal matrix whose diagonal elements are the same as $\log\rho$. Then, $C^{*}(\rho)$ is calculated as follows.
Lemma 7
The relation
$C^{*}(\rho)=D(I_{c}(\rho)\|\rho)$ (43)
holds.
Proof: For $\sigma\in{\cal I}_{c}$, we have
$D(\sigma\|\rho)=\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log\rho)$
$=$ $\operatorname{\mathrm{Tr}}\sigma(\log\sigma-(\log\rho)^{\rm diag}))$
$=$ $\operatorname{\mathrm{Tr}}\sigma(\log\sigma-\log I_{c}(\rho)-\log t(\rho)$
$=$ $D(\sigma\|I_{c}(\rho))-\log t(\rho).$ (44)
Hence, the minimization is realized when $\sigma=I_{c}(\rho)$.
Lemma 8
The relation
$C_{\alpha}^{*}(\rho)$
$=$ $\left\{\begin{array}[]{l}\max_{\tau}\frac{\alpha}{\alpha-1}\log\|(\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|\\ \hskip 150.69397pt\hbox{ for }\alpha\in[1/2,1)\\ \max_{\tau}\frac{-\alpha}{\alpha-1}\log\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|\\ \hskip 150.69397pt\hbox{ for }\alpha>1\end{array}\right.$ (49)
holds.
Proof: The idea of this proof comes from the proof of [67, Lemma 6 and (B4)]. We choose $\beta$ such that $\frac{1}{\beta}+\frac{1}{\alpha}=1$ so that $\beta=\frac{\alpha-1}{\alpha}$. We fix $\alpha\in[1/2,1)$ so that $\beta\in[-1,0)$. For any state $\tau$ and any incoherent state $\sigma\in{\cal I}_{c}$, the reverse matrix Hölder inequality [70, Theorem 2.7] implies
$\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}\geq(\operatorname{\mathrm{Tr}}\tau)^{\beta}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}$
$=$ $(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (50)
Considering the equality condition, we have
$\min_{\tau}\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}=(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (51)
Hence,
$e^{(\alpha-1)C_{\alpha}^{*}(\rho)}=\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $(\max_{\sigma\in{\cal I}_{c}}\min_{\tau}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $(\min_{\tau}\max_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
$=$ $\min_{\tau}\|(\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\|^{\alpha}.$ (52)
Here, $(a)$ follows from the minimax theorem because $\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma$ is linear for $\sigma$ and is convex for $\tau$, which follows from the matrix convexity of $\tau\mapsto\tau^{\beta}$ with $\beta\in[-1,0)$. Hence, we obtain the desired equation with $\alpha\in(0,1)$.
Next, we fix $\alpha>1$ so that $01>\beta>$ 0. The matrix Hölder inequality [71],[70, Theorem 2.6] implies
$\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}\leq(\operatorname{\mathrm{Tr}}\tau)^{\beta}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}$
$=$ $(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (53)
Considering the equality condition, we have
$\max_{\tau}\operatorname{\mathrm{Tr}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}}=(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{1}{\alpha}}.$ (54)
Hence,
$e^{(\alpha-1)C_{\alpha}^{*}(\rho)}=\min_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $(\min_{\sigma\in{\cal I}_{c}}\max_{\tau}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $(\max_{\tau}\min_{\sigma\in{\cal I}_{c}}\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma)^{\alpha}$
$=$ $(\max_{\tau}\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|^{-1})^{\alpha}$
$=$ $\max_{\tau}\|((\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}})^{\rm diag})^{-1}\|^{-\alpha}.$ (55)
Here, $(a)$ follows from the minimax theorem because $\operatorname{\mathrm{Tr}}\rho^{\frac{1-\alpha}{2\alpha}}\tau^{\beta}\rho^{\frac{1-\alpha}{2\alpha}}\sigma$ is linear for $\sigma$ and is concave for $\tau$. Hence, we obtain the desired equation with $\alpha>1$.
Coherence measures satisfy the additivity.
Lemma 9
The relations
$\begin{split}C^{*}(\rho_{1}\otimes\rho_{2})&=C^{*}(\rho_{1})+C^{*}(\rho_{2}),\\ C_{\alpha}^{*}(\rho_{1}\otimes\rho_{2})&=C_{\alpha}^{*}(\rho_{1})+C_{\alpha}^{*}(\rho_{2})\end{split}$ (56)
hold with $\alpha\in[1/2,1)\cup(1,\infty)$.
Proof: Since the first equation in (56) follows from the second equation in (56) with taking the limit $\alpha\to 1$, we show only the second equation in (56). The idea of this proof is similar to the proof of [67, Lemma 7]. For $\alpha\in(0,1)$, using (49), we have
$C_{\alpha}^{*}(\rho_{1})+C_{\alpha}^{*}(\rho_{2})\geq C_{\alpha}^{*}(\rho_{1}\otimes\rho_{2})$
$=$ $\max_{\tau}\frac{\alpha}{\alpha-1}\log\big{\|}((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}\tau^{-\frac{1-\alpha}{\alpha}}(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}$
$\geq$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}\log\big{\|}((\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}}(\tau_{1}\otimes\tau_{2})^{-\frac{1-\alpha}{\alpha}}$
$\hskip 86.11084pt\cdot(\rho_{1}\otimes\rho_{2})^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}$
$=$ $\max_{\tau_{1},\tau_{2}}\frac{\alpha}{\alpha-1}\big{(}\log\big{\|}(\rho_{1}^{\frac{1-\alpha}{2\alpha}}\tau_{1}^{-\frac{1-\alpha}{\alpha}}\rho_{1}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}$
$+\log\big{\|}(\rho_{2}^{\frac{1-\alpha}{2\alpha}}\tau_{2}^{-\frac{1-\alpha}{\alpha}}\rho_{2}^{\frac{1-\alpha}{2\alpha}})^{\rm diag}\big{\|}\big{)}$
$=$ $C_{\alpha}^{*}(\rho_{1})+C_{\alpha}^{*}(\rho_{2}),$ (57)
which implies the second equation in (56). The case with $\alpha>1$ can be shown in the same way.
TABLE II: List of mathematical symbols for coherence.
Symbol Description Eq. number
${\cal I}_{c}$ Set of incoherent states
$C^{*}(\rho)$ Minimum relative entropy (36)
$C_{\alpha}^{*}(\rho)$ Minimum relative Rényi (36)
entropy of order $\alpha$
$I_{c}(\rho)$ Closest incoherent state (43)
$\sigma(p)$ $\sum_{i}p_{i}|i\rangle\langle i|$
$\rho(p)$ $\sum_{j=0}^{d-1}p_{j}Z^{j}|+\rangle\langle+|Z^{-j}$
Remark 2
Although the above-presented proofs employ several techniques that come from prior works [62, 63], the above proofs are not simple combinations of techniques that come from prior works. Since the two entries of the relative entropy and relative Rényi entropy are exchanged in our measure, the above proofs require new careful treatments for matrix convexity related to the parameter $\alpha$.
Remark 3
It is possible to make the same discussion for the coherence as Sections IIIV. Since the coherence measures satisfy the additivity as Lemma 9, the detectability even with the non-iid setting like Section III can be characterized by the quantities $C^{*}(\rho)$ and $C_{\alpha}^{*}(\rho)$.
VI Maximally correlated states
The discussion in Section II shows the operational meaning of $E_{\alpha}^{*}(\rho_{AB})$ including $E^{*}(\rho_{AB})$. We calculate $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ for several special cases. When $\rho_{AB}$ is a pure state $|\phi\rangle\langle\phi|$ $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha>1$ are infinity. However, $E^{*}_{\alpha}(\rho_{AB})$ with $\alpha\in[0,1)$ takes a finite value as follows. When the Schmidt coefficients of $|\phi\rangle\langle\phi|$ are $\sqrt{p_{i}}$ with $i=1,\ldots,d$, $\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma|\phi\rangle=\max_{i}p_{i}$ [63, (40) and (97)]. Hence,
$e^{(\alpha-1)E_{\alpha}^{*}(\rho_{AB})}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}\big{(}\rho^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho^{\frac{1-\alpha}{2\alpha}}\big{)}^{\alpha}$
$=$ $\max_{\sigma_{AB}\in{\cal S}_{AB}}\operatorname{\mathrm{Tr}}\big{(}|\phi\rangle\langle\phi|\sigma_{AB}|\phi\rangle\langle\phi|\big{)}^{\alpha}=\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle^{\alpha}$
$=$ $\Big{(}\max_{\sigma_{AB}\in{\cal S}_{AB}}\langle\phi|\sigma_{AB}|\phi\rangle\Big{)}^{\alpha}=(\max_{i}p_{i})^{\alpha}.$ (58)
Hence, we have
$E_{\alpha}^{*}(\rho_{AB})=\frac{\alpha}{\alpha-1}\log(\max_{i}p_{i}).$ (59)
More generally, the calculation of these quantities can be converted to the calculation of the coherence measure for special states $\rho_{AB}$. Similar to the paper [63], we consider maximally correlated states. A state $\rho_{AB}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ is called a maximally correlated state when there exist bases $\{|u_{j,A}\rangle\}_{j}$ and $\{|u_{j,B}\rangle\}_{j}$ of ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ such that the support $\rho_{AB}$ is included in the subspace spanned by $\{|u_{j,A}\rangle|u_{j,B}\rangle\}_{j}$ on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ [68, 69], [12, Section 8.8]. In this case, the state $\rho_{AB}$ is written as $\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|$. In this case, the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ can be simplified as follows. Similar to Theorem 1 in [63], the calculation of $E^{*}(\rho_{AB})$ and $E^{*}_{\alpha}(\rho_{AB})$ are reduced to coherence measures.
Lemma 10
The relations
$E^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=$ $C^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}$ (60)
$E_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=$ $C_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}$ (61)
hold.
Lemma 10 is shown in Appendix B. Since a pure state is a maximally correlated state, the calculation (59) can be recovered by Lemma 10 and (38). Also, $E^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$ can be calculated by Lemma 7 as
$E^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=D(I_{c}(\rho_{\theta})\|\rho_{\theta}),$ (62)
where $\rho_{\theta}:=\sum_{i,j}\theta_{i,j}|j\rangle\langle i|$. The combination of Lemmas 9 and 10 guarantees that any maximally entangled state $\rho_{AB}$ satisfies
$E^{*}(\rho_{AB})=\overline{E}^{*}(\rho_{AB})=\tilde{E}^{*}(\rho_{AB}),~{}E_{\alpha}^{*}(\rho_{AB})=\overline{E}_{\alpha}^{*}(\rho_{AB}).$ (63)
The iid setting and the non-iid setting have the same bounds.
To find another example, we set ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ to be $d$-dimensional spaces, and define the maximally entangled state $|\Phi\rangle:=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle|j\rangle$. We define the state $\rho_{AB}(p):=\sum_{j=0}^{d-1}p_{j}Z^{j}\otimes I|\Phi\rangle\langle\Phi|Z^{-j}\otimes I$. The combination of (39), (40), Lemma 10 implies
$E^{*}(\rho_{AB}(p))$ $=D(p_{\mathop{\rm mix}}\|p)=-\frac{1}{d}\sum_{j=0}^{d}\log p_{j}-\log d$ (64)
$E_{\alpha}^{*}(\rho_{AB}(p))$ $=D_{\alpha}(p_{\mathop{\rm mix}}\|p)$
$=\frac{1}{\alpha-1}\log\sum_{j=0}^{d-1}p_{j}^{1-\alpha}-\frac{\alpha}{\alpha-1}\log d$ (65)
for $\alpha\in(0,1)\cup(1,2]$. However, it is not easy to calculate these quantities for general states. Therefore, it is needed to develop an efficient method for their calculation in the general case.
VII Algorithms
VII-A Algorithm for $E^{*}(\rho_{AB})$
Although the previous section presents the exact values of the proposed measures of maximally correlated states, it is not so easy to calculate the proposed measures of a general density matrix. To calculate the proposed quantities for general states, we introduce algorithms for their calculation that can be applied to general states. Since the calculation of $E^{*}(\rho_{AB})$ is hard, we modify $E^{*}(\rho_{AB})$ as follows. We choose a finite set ${\cal D}$ of pure states on ${\cal H}_{B}$ as a discretization. That is, we denote ${\cal D}$ as $\{\rho_{B|c}^{*}\}_{c\in{\cal X}}$. Then, we define
$E_{{\cal D}}^{*}(\rho_{AB}):=\min_{P(c),\rho_{A|c}}D\Big{(}\sum_{c}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (66)
We define $\epsilon_{2}({\cal D})$ as
$\epsilon_{2}({\cal D}):=2\sqrt{1-\min_{\psi_{1}}\max_{\psi_{2}\in{\cal D}}|\langle\psi_{1}|\psi_{2}\rangle|^{2}}.$ (67)
We have
$\max_{P^{\prime}(c),\rho_{A|c}^{\prime},\rho_{B|c}^{\prime}}\min_{P(c),\rho_{A|c}}\Big{\|}\sum_{c}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*}$
$-\sum_{c}P^{\prime}(c)\rho_{A|c}^{\prime}\otimes\rho_{B|c}^{\prime}\Big{\|}_{1}$
$\leq$ $\max_{\psi_{1}}\min_{\psi_{2}\in{\cal D}}\||\psi_{1}\rangle\langle\psi_{1}|-|\psi_{2}\rangle\langle\psi_{2}|\|_{1}$
$=$ $2\sqrt{1-\min_{\psi_{1}}\max_{\psi_{2}\in{\cal D}}|\langle\psi_{1}|\psi_{2}\rangle|^{2}}=\epsilon_{2}({\cal D}).$ (68)
Therefore, Fannes inequality implies
$\max_{P^{\prime}(c),\rho_{A|c}^{\prime},\rho_{B|c}^{\prime}}\min_{P(c),\rho_{A|c}}\Big{|}D\Big{(}\sum_{x}P(x)\rho_{A|x}\otimes\rho_{B|x}^{*}\Big{\|}\rho_{AB}\Big{)}$
$\hskip 86.11084pt-D\Big{(}\sum_{x}P^{\prime}(x)\rho_{A|x}^{\prime}\otimes\rho_{B|x}^{\prime}\Big{\|}\rho_{AB}\Big{)}\Big{|}$
$\leq$ $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D})),$ (69)
where $\eta_{0}$ is defined as
$\eta_{0}(x):=\left\{\begin{array}[]{ll}-x\log x&\hbox{when }0\leq x\leq 1/e\\ 1/e&\hbox{otherwise.}\end{array}\right.$ (72)
Therefore, we have
$E_{{\cal D}}^{*}(\rho_{AB})$
$-\Big{(}\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))\Big{)}$
$\leq$ $E^{*}(\rho_{AB})\leq E_{{\cal D}}^{*}(\rho_{AB}).$ (73)
We rewrite $P(c)\rho_{A|c}$ to $X_{c}$. Using the idea of the algorithm by [28], we propose an iterative algorithm to calculate $E_{{\cal D}}^{*}(\rho_{AB})$, in which matrices $X_{A|c}$ is updated as follows. Given the set $\{X_{c}^{(t)}\}_{c}$, we define $V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$ and $\kappa[\{X_{A|c}^{(t)}\}_{c}]$ as
$V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$
::=:= $\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}^{(t)}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$\hskip 86.11084pt-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)},$ (74)
$\kappa[\{X_{A|c}^{(t)}\}_{c}]:=\sum_{c}\operatorname{\mathrm{Tr}}\exp(X_{A|c}^{(t)}-V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (75)
Then, we choose the updated matrix ${X}_{A|c}^{(t+1)}$
${X}_{A|c}^{(t+1)}:=$ $\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (76)
This algorithm is summarized as Algorithm 1.
Algorithm 1 Minimization of $D\Big{(}\sum_{c}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$
1:Choose the initial value $\{X_{A|c}^{(t)}\}_{c}$;
2:repeat
3:     Calculate ${X}_{A|c}^{(t+1)}:=\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-V_{c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}])$;
4:until convergence of $D\Big{(}\sum_{c}X_{A|c}^{(t)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$.
Then, we have the following theorem.
Theorem 11
We choose ${X}_{A|c}^{(1)}$ such that $\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes{X}_{A|c}^{(1)}$ is the completely mixed state $\rho_{AC,\mathop{\rm mix}}$. Then, we have
$D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}-E_{{\cal D}}^{*}(\rho_{AB})\leq\frac{\log d_{A}|{\cal D}|}{t}.$ (77)
The above theorem is shown in Appendix C by using [28, Theorem 3.3].
VII-B Algorithm for $E_{\alpha}^{*}(\rho_{AB})$
Next, we consider the algorithm for $E_{\alpha}^{*}(\rho_{AB})$. Similarly, instead of $E_{\alpha}^{*}(\rho_{AB})$, we calculate
$E_{\alpha,{\cal D}}^{*}(\rho_{AB}):=\min_{X_{A|c}}D_{\alpha}\Big{(}\sum_{c}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (78)
The following algorithm minimizes $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$. For this aim, we iteratively choose matrices $X_{A|c}$ as follows. Given the set $\{X_{c}^{(t)}\}_{c}$, we define $V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$ and $\kappa_{\alpha}[\{X_{A|c}^{(t)}\}_{c}]$ as
$V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]$
::=:= $\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}^{(t)}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
$\cdot\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*}),$ (79)
$\kappa_{\alpha}[\{X_{A|c}^{(t)}\}_{c}]$
::=:= $\sum_{c}\operatorname{\mathrm{Tr}}\exp(X_{A|c}^{(t)}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (80)
Then, we choose the updated matrix ${X}_{A|c}^{(t+1)}$
${X}_{A|c}^{(t+1)}$
::=:= $\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}]).$ (81)
Algorithm 2 Minimization of $\mathop{\rm sgn}(\alpha-1)e^{(\alpha-1)D_{\alpha}\Big{(}\sum_{c}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}}$
1:Choose the initial value $\{X_{A|c}^{(t)}\}_{c}$;
2:repeat
3:     Calculate ${X}_{A|c}^{(t+1)}:=\frac{1}{\kappa[\{X_{A|c}^{(t)}\}_{c}]}\exp(X_{A|c}^{(t)}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}])$;
4:until convergence of $D_{\alpha}\Big{(}\sum_{c}X_{A|c}^{(t)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$.
Then, we have the following theorem.
Theorem 12
We choose ${X}_{A|c}^{(1)}$ as $\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes{X}_{A|c}^{(1)}$ as the completely mixed state $\rho_{AC,\mathop{\rm mix}}$. For $\alpha\in(0,1)$, and a sufficiently large real number $\gamma>0$, we have
$\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}^{(t+1)}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$
$-\mathop{\rm sgn}(\alpha-1)e^{(\alpha-1)E_{\alpha,{\cal D}}(\rho_{AB})}\leq\frac{\gamma\log d_{A}|{\cal D}|}{t}.$ (82)
The above theorem is shown in Appendix C by using [28, Theorem 3.3].
VIII Application of our algorithm
VIII-A Membership problem
The previous section presents algorithms to calculate the proposed measures. The special case of deciding if the entanglement measure is zero can be used for testing separability in a mathematical sense, which is called the membership problem of the set ${\cal S}_{AB}$. The aim of this section is the application of our algorithm to the membership problem of the set ${\cal S}_{AB}$. That is, we employ the quantity $E^{*}(\rho_{AB})$ for the membership problem. This measure will be used for constructing the separation hyperplane. In the following, we consider how to distinguish the following two cases.
$H_{0}:$ $E^{*}(\rho_{AB})=0$ (83)
$H_{1}:$ $E^{*}(\rho_{AB})\geq\epsilon_{1}.$ (84)
Due to (73), when $E^{*}(\rho_{AB})=0$, we have
$E_{{\cal D}}^{*}(\rho_{AB})\leq\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D})).$ (85)
Also, when $E^{*}(\rho_{AB})>\epsilon$, we have
$E_{{\cal D}}^{*}(\rho_{AB})>\epsilon.$ (86)
Therefore, if the following two cases are distinguished, $H_{0}$ and $H_{1}$ are distinguished.
$H_{0}^{\prime}:E_{{\cal D}}^{*}(\rho_{AB})\leq$ $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})$
$+\eta_{0}(\epsilon_{2}({\cal D})),$ (87)
$H_{1}^{\prime}:E_{{\cal D}}^{*}(\rho_{AB})>$ $\epsilon_{1}$ (88)
under the assumption that $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))\leq\epsilon_{1}$.
We choose the initial state according to the assumption of Theorem 11. Then, when $H_{0}^{\prime}$ holds, i.e., $E_{{\cal D}}^{*}(\rho_{AB})\leq\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))$, we have
$D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$
$\leq$ $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))+\frac{\log d_{A}|{\cal D}|}{t}.$ (89)
Also, when $E_{{\cal D}}^{*}(\rho_{AB})>\epsilon$, we have
$D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}>\epsilon.$ (90)
Therefore, if the following two cases are distinguished, $H_{0}^{\prime}$ and $H_{1}^{\prime}$, i.e. $H_{0}$ and $H_{1}$ are distinguished.
$H_{0}^{\prime\prime}:$ $D\Big{(}\sum_{c}X_{A|c}^{(t+1)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$
$\leq\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))$
$\quad+\frac{\log d_{A}|{\cal D}|}{t},$ (91)
$H_{1}^{\prime\prime}:$ $\operatorname{\mathrm{Tr}}\rho_{AC}^{(t+1)}\Omega_{F}[\rho_{AC}^{(t+1)}]>\epsilon_{1}$ (92)
under the assumption that $\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D}))+\frac{\log d_{A}|{\cal D}|}{t}\leq\epsilon_{1}$. That is, the number $t$ of iterations needs to satisfy $t\geq\log d_{A}|{\cal D}|/\Big{(}\epsilon_{1}-(\epsilon_{2}({\cal D})(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}({\cal D})))\Big{)}$.
Next, we evaluate the computation complexity to distinguish $H_{0}$ and $H_{1}$ under the above method when we choose ${\cal D}_{n,d_{B}}$ as ${\cal D}$ in the way as Appendix A and we fix $\epsilon_{1}$ and move $d_{A},d_{B}$ with $d_{A}\geq d_{B}$.
In this case, we have $|{\cal D}_{n,d_{B}}|\leq(\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}})^{d_{B}-1}$ with $\epsilon_{2}({\cal D}_{n,d_{B}})=\epsilon_{2}$. The above condition is rewritten as
$t\geq\frac{\log d_{A}+(d_{B}-1)\log\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}}}{\epsilon_{1}-(\epsilon_{2}(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}))}.$ (93)
Since $\epsilon_{1}$ is fixed, we need to choose $\epsilon_{2}$ to be $O(\frac{\epsilon_{1}}{\log d_{A}d_{B}})$, which realizes $\epsilon_{1}-(\epsilon_{2}(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}))=O(\epsilon_{1})$. The above condition is rewritten as
$t\geq$ $\frac{\log d_{A}+(d_{B}-1)\log\frac{2d_{B}^{1/2}\pi}{O(\frac{\epsilon_{1}}{\log d_{A}d_{B}})}}{O(\epsilon_{1})}$
$=$ $O\Big{(}\frac{\log d_{A}+(d_{B}-1)\log\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}}}{\epsilon_{1}}\Big{)}$
$=$ $O\Big{(}\frac{\log d_{A}+d_{B}\log d_{B}+d_{B}(\log\log d_{A}-\log\epsilon_{1})}{\epsilon_{1}}\Big{)}.$ (94)
However, each step has various calculations. In each step, we need to calculate ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$. For this calculation, we need to calculate $\log\rho_{AC}$ and $\Omega_{R,D}[\rho_{AC}]$. $\rho_{AC}$ is a block diagonal matrix. That is, it is sufficient to calculate every component. Every component has a matrix function on a $d_{A}d_{B}$-dimensional space, and a factor whose complexity is linear in $|{\cal D}_{n,d_{B}}|$. But, the latter part is common for all $c$-components. The former can be done by diagonalization. The complexity of the diagonalization on $d$-dimensional matrix is $O(d^{3})$ with error $\epsilon$ [64, 65, 66]. Therefore, the calculation complexity of ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$ is $O(|{\cal D}_{n,d_{B}}|(d_{A}^{3}d_{B}^{3}))\leq O((\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))=O((\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))$. The total complexity is
$O\Big{(}$ $\frac{\log d_{A}+d_{B}\log d_{B}+d_{B}(\log\log d_{A}-\log\epsilon_{1})}{\epsilon_{1}}$
$\cdot(\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3})\Big{)}.$ (95)
On the other hand, the method by [31, Eq. (38)] has the following calculation complexity
$O(d_{A}^{6}\epsilon^{-4d_{B}}).$ (96)
In the above evaluation, $\epsilon$ takes the same role as $\epsilon_{1}$. For $d_{B}$, our bound has the factor $d_{B}^{(d_{B}-1)/2}$, which is larger than $\frac{1}{\epsilon^{4d_{B}}}$. However, when $d_{B}$ is also fixed and only $d_{A}$ increases, our bound is simplified to $O(\max(\log d_{A},-\log\epsilon_{1})(\log d_{A})^{d_{B}-1}d_{A}^{3}\epsilon_{1}^{-d_{B}})$. When $\epsilon_{1}$ and $\epsilon$ are fixed, our bound is simplified to $O((\log d_{A})^{d_{B}}d_{A}^{3})$, and their bound is simplified to $O(d_{A}^{6})$. In this case, our bound is smaller than their bound.
VIII-B Separation hyperplane, i.e., entanglement witness
A linear function to determine a separation hyperplane is called an entanglement witness in the context of entanglement theory when the cone is the separable cone. Next, we discuss how to obtain a separation hyperplane by using the above approximation and the geometrical characterization by Theorem 3. This is because it is a key step for conic linear programming with a general cone to find a separation hyperplane for an element that does not belong to the cone [35, 36, 37, 38]. For this aim, we employ the minimizer of $E_{{\cal D}}^{*}(\rho_{AB})$, i.e., the state
$\sigma_{AB}^{*}({\cal D}):=\mathop{\rm argmin}\limits_{\sigma_{AB}\in{\cal S}({\cal D})}D(\sigma_{AB}\|\rho_{AB}).$ (97)
Then, the following theorem means that the operator $(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})$ gives a separation hyperplane to distinguish $\rho_{AB}$ and ${\cal S}$.
Theorem 13
(i) For any separable state $\sigma_{AB}\in{\cal S}$, we have
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})$
$\geq$ $D(\sigma_{AB}^{*}({\cal D})\|\rho_{AB})-\|\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB}\|\epsilon_{2}({\cal D}).$ (98)
(ii) Also, we have
$\operatorname{\mathrm{Tr}}\rho_{AB}(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})=-D(\rho_{AB}\|\sigma_{AB}^{*}({\cal D}))<0.$ (99)
Proof: Since (ii) is trivial, we show only (i) by contradiction. Since the set ${\cal S}({\cal D})$ is the convex set, we have the following in the same way as Theorem 3. (i) For any separable state $\sigma_{AB}\in{\cal S}({\cal D})$, we have
$\operatorname{\mathrm{Tr}}\sigma_{AB}(\log\sigma_{AB}^{*}({\cal D})-\log\rho_{AB})\geq D(\sigma_{AB}^{*}({\cal D})\|\rho_{AB}).$ (100)
Due to the combination of (100) and (68), any separable state $\sigma_{AB}\in{\cal S}$ satisfies (98).
IX Numerical analysis
Next, we numerically demonstrate our algorithm when ${\mathcal{H}}_{A}$ is two qubits and ${\mathcal{H}}_{B}$ is one qubit. We parametrize a qubit pure state as $|\phi_{\theta}\rangle:=\frac{1}{\sqrt{2}}(|0\rangle+e^{\sqrt{-1}\theta}|1\rangle)$. Consider the entangled state $|\Psi_{p}\rangle:=\sqrt{p}|\phi_{0}\rangle|\phi_{0}\rangle|0\rangle+\sqrt{1-p}|\phi_{\pi/2}\rangle|\phi_{\pi/2}\rangle|1\rangle)$. Given a real number $\lambda$ with $10<\lambda<$ 1, we focus on the phase-damping channel ${\cal E}_{\lambda}$ defined as
${\cal E}_{\lambda}\left(\left(\begin{array}[]{cc}a_{0,0}&a_{0,1}\\ a_{1,0}&a_{1,1}\end{array}\right)\right):=\left(\begin{array}[]{cc}a_{0,0}&\lambda a_{0,1}\\ {\lambda}a_{1,0}&a_{1,1}\end{array}\right).$ (105)
This channel acts on the state $|\phi_{\theta}\rangle$ as
${\cal E}_{\lambda}(\rho)=\frac{1+\lambda}{2}\rho+\frac{1-\lambda}{2}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\rho\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right).$ (110)
That is, the phase-damping channel is replaced by the above over the state $|\Psi_{p}\rangle$. Then, we define the state with two parameters $\lambda$ and $\delta$ as
$\rho_{p,\lambda,\delta}:=(1-\delta)({\cal E}_{\lambda}\otimes{\cal E}_{\lambda}\otimes Id)(|\Psi_{p}\rangle\langle\Psi_{p}|)+\delta\rho_{mix},$ (111)
where $\delta\rho_{mix}$ is the completely mixed state on ${\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$.
Applying the unitary matrix $V=\frac{1}{2}\left(\begin{array}[]{cc}1+\sqrt{-1}&1-\sqrt{-1}\\ 1-\sqrt{-1}&1+\sqrt{-1}\end{array}\right)$ on both qubits in ${\mathcal{H}}_{A}$, the above model is converted to the following model. We define $|\tilde{\Psi}_{p}\rangle:=\sqrt{p}|\phi_{\pi}\rangle|\phi_{\pi}\rangle|0\rangle+\sqrt{1-p}|0\rangle|0\rangle|1\rangle)$. We define the channel ~$\tilde{\cal E}_{\lambda}$ as
~$\tilde{\cal E}_{\lambda}(\rho):=\frac{1+\lambda}{2}\rho+\frac{1-\lambda}{2}\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)\rho\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right).$ (116)
Then, we define the state with two parameters $\lambda$ and $\delta$ as
$\tilde{\rho}_{p,\lambda,\delta}:=(1-\delta)\big{(}\tilde{\cal E}_{\lambda}\otimes\tilde{\cal E}_{\lambda}\otimes Id\big{)}(|\tilde{\Psi}_{p}\rangle\langle\tilde{\Psi}_{p}|)+\delta\rho_{mix}.$ (117)
That is, we have $(V^{\otimes 2}\otimes I){\rho}_{p,\lambda,\delta}(V^{\otimes 2}\otimes I)^{\dagger}=\tilde{\rho}_{p,\lambda,\delta}$. Hence, this model is essentially written with real matrix elements.
We make numerical experiments when $n$ is set to be $16$. That is, $|{\cal D}_{n,d_{B}}|$ is $16^{2}$. By setting $p=1/2$, Fig. 3 plots the minimum relative entropy $E_{{\cal D}}^{*}({\rho}_{1/2,\lambda,\delta})$ with $\delta=10^{-1},10^{-2},10^{-4}$. When $\lambda$ is close to zero, the state ${\rho}_{p,\lambda,\delta}$ approaches to the set ${\cal S}_{AB}$ of separable states.
Figure 3: Minimum relative entropy $E_{{\cal D}}^{*}({\rho}_{1/2,\lambda,\delta})$: The vertical axis shows the minimum relative entropy. The horizontal axis shows the parameter $\lambda$. Blue (orange, green) curve presents the case with $\delta=10^{-1}$, ($10^{-2},10^{-4}$).
Next, we apply our algorithm to the membership problem of separability by choosing $p=0.1$. For this aim, ideally, it is needed to plot the boundary to satisfy the condition $E_{{\cal D}}^{*}({\rho}_{0.1,\lambda,\delta})=0$. However, due to the limitation of the performance of our computer, we cannot handle a too small number. Hence, instead of the ideal value 00, we set the same small number $10^{-8}$. That is, Figure 4 plots the boundary to satisfy the condition $E_{{\cal D}}^{*}({\rho}_{0.1,\lambda,\delta})=10^{-8}$. To compare our method with the partial transpose negative test. This test also has the same problem to handle the case with zero eigenvalue. Instead of the ideal value 00, we set the same small number $-10^{-8}$ in the same way. That is, this figure plots the boundary to satisfy the condition that the minimum eigenvalue of the partial transposed state is $-10^{-8}$.
These comparisons show that our method can detect entangled states that cannot be detected by the partial transpose negative test even when all the matrix components of the density matrix of a $4\times 2$ system are written in real numbers.
Figure 4: Comparison between our method and the partial transpose negative test: The vertical axis shows the parameter $\lambda$. The horizontal axis shows the parameter $\delta$. Blue (Orange) curve shows our method (the partial transpose negative test). The upper regions of both curves are composed of states that are guaranteed to be entangled by the respective methods.
X Exponential quantum Sanov theorem
Finally, to clarify our contribution on quantum Sanov theorem, we state our obtained result for quantum Sanov theorem in this section. For this aim, we consider a $d$-dimensional Hilbert space ${\cal H}$ and a state $\rho$ on ${\mathcal{H}}$. We assume the $n$-fold independently and identical distributed (iid) condition and consider the following two hypotheses.
$H_{0}:$ The state is $\hbox{The state is }\rho^{\otimes n},$ (118)
$H_{1}:$ The state is $\hbox{The state is }\sigma^{\otimes n}\hbox{ with }\sigma\in S.$ (119)
To address this problem, we focus on the following quantity.
$\beta_{\epsilon,n}(\rho):=\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho^{\otimes n}\Big{|}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}(I-T)\sigma^{\otimes n}\leq\epsilon\Big{\}}.$ (120)
The relation
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{\epsilon,n}(\rho)=E^{*}(\rho):=\min_{\sigma\in S}D(\sigma\|\rho)$ (121)
is known as quantum Sanov theorem [21, 22]. To analyze the case when the value $\epsilon$ goes to zero exponentially, as the extension of $E^{*}(\rho)$, we define the minimum sandwiched relative entropy as
$E_{\alpha}^{*}(\rho):=\min_{\sigma\in S}D_{\alpha}(\sigma\|\rho).$ (122)
As an exponential version of quantum Sanov theorem, the following theorem holds.
Theorem 14
We have
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$ $\geq\max_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}$ (123)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)$ $\leq\min_{\alpha\geq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (124)
More precisely, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$
$\geq$ $-\frac{(d+2)(d-1)}{n}\log(n+1)$
$+\max_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (125)
Our bound is the optimal when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. Although the paper [22] derived an exponential evaluation for both error probabilities, its exponent does not equal to the optimal bound even when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. In this sense, our evaluation is better than the upper bound by [22].
The proof is given in Appendix D. Theorem 14 proves Theorem 1. Under the limit $r\to 0$, the RHS of (6) goes to $E_{1}^{*}(\rho)=E^{*}(\rho)$. Hence, the relation (6) recovers the inequality $\geq$ in (4) while the opposite inequality in (4) follows from the converse part of quantum Stein’s lemma.
XI Discussion
We have proposed new entanglement measures based on the distinguishability of the given entangled state from the set of separable states with two kinds of formulations by using quantum relative entropy and sandwiched Rényi relative entropy. That is, we have proposed two kinds of measures based on quantum relative entropy. This measure is different from the relative entropy of entanglement [8, 9] because the choice of two input states are opposite to each other.
In the first formulation, the possible separable states are limited to tensor product states in this discrimination problem. This case can be considered as a special case of quantum Sanov theorem [21, 22]. Our analysis realizes an exponential evaluation for both kinds of error probabilities, which is better than the existing results [22]. In the second formulation, the possible separable states are not limited to tensor product states. This problem setting cannot be considered as a special case of quantum Sanov theorem. Even in this setting, we have derived an exponential evaluation for both kinds of error probabilities. Based on these two formulations, we have proposed two kinds of entanglement measures. The former measure is not smaller than the latter measure.
Also, the proposed quantity is useful for finding entanglement witness based on its information-geometrical structure. In the next step, we have derived several useful calculation formulas for these measures. In particular, we have shown that these two kinds of measures coincide with each other when the entangled state is maximally correlated. In particular, the entanglement measure based on the quantum relative entropy has been concretely derived. However, it is not easy to calculate these measures for general entangled states. To resolve this problem, we have proposed algorithms to calculate these measures. In the first step of this calculation, we have approximated these values by introducing the $\delta$-net on the second quantum system, which was introduced to derive an entanglement witness [27, Section 3]. Then, we apply the quantum Arimoto-Blahut algorithm [28] to these approximating values. Then, we have obtained an efficient algorithm for these approximating values. These algorithms are shown to converge to the global minimum, and their errors are evaluated in a simple form.
Finally, we have applied our algorithm to the membership problem of separability. Although it is known to be NP-hard when the dimensions of both Hilbert spaces increase [24, 25, 26, 27], the calculation complexity is not so large when the dimension of the second Hilbert space is fixed [31]. In this case, our algorithm has a smaller calculation complexity than the algorithm by [31]. Also, we have applied our algorithm to find an entanglement witness.
As a future problem, we can consider the following algorithm. We choose a $\delta$ net on the second system ${\mathcal{H}}_{B}$, and apply our algorithm. Then, by fixing the obtained states on the system ${\mathcal{H}}_{A}$, our algorithm exchanges the roles of ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$ so that we derive states on ${\mathcal{H}}_{B}$. In this way, we iteratively improve the states on ${\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B}$. We might expect that this algorithm works well even when the $\delta$ is not so small. However, this algorithm solves the membership problem of separability, which is known to be NP-hard when the dimensions of both Hilbert spaces increase. Hence, it is impossible to choose so small $\delta$ that the calculation time complexity is polynomial for the local dimension. It is an interesting problem whether such an iterative algorithm can improve our proposed algorithm.
Acknowledgement
MH was supported in part by the National Natural Science Foundation of China under Grant 62171212. MH is thankful to Professor Hiroshi Hirai for helpful discussions on the membership problem for separability and the importance of separation hyperplane, and informing the references [25, 30, 35, 31, 33, 34]. MH is thankful to Professor Marco Tomamichel and Mr. Roberto Rubboli for informing the references [32], [55], [56]. MH is grateful to Dr. Baichu Yu for informing the references [46, 47]. MH and YI are grateful for Professor Masaki Owari to a helpful discussion on the membership problem of separability and informing the reference [27].
Appendix A Construction of discrete subset
A-A Real discrete subset
First, similar to [39, Section 5.4], we construct the real discrete subset ${\cal D}_{n,d,R}\subset{\cal S}_{d}$, where ${\cal S}_{d}$ is the $d$-dimensional sphere in $\mathbb{R}^{d+1}$. We define ${\cal D}_{n,1,R}$ as
${\cal D}_{n,1,R}:=\Big{\{}\Big{(}\cos\frac{2\pi j}{n},\sin\frac{2\pi j}{n}\Big{)}\Big{\}}_{j=0}^{n-1}.$ (126)
Then, we inductively define ${\cal D}_{n,d,R}$ as
${\cal D}_{n,d,R}$
::=:= $\Big{\{}\Big{(}\cos\frac{2\pi j}{n}v,\sin\frac{2\pi j}{n}\Big{)}\Big{|}v\in{\cal D}_{n,d-1,R},j=0,\ldots,n-1\Big{\}}.$ (127)
We have
$F({\cal D}_{n,1,R}):=$ $\min_{\psi_{1}\in{\cal S}_{2}}\max_{\psi_{2}\in{\cal D}_{n,2,R}}|\langle\psi_{1}|\psi_{2}\rangle|=\cos\frac{\pi}{n}.$ (128)
In general, we have
$F({\cal D}_{n,d,R})=$ $\min_{\psi_{1}\in{\cal S}_{d}}\max_{\psi_{2}\in{\cal D}_{n,d,R}}|\langle\psi_{1}|\psi_{2}\rangle|=\cos^{d}\frac{\pi}{n}.$ (129)
The relation (129) can be inductively shown as (130) in the next page.
$F({\cal D}_{n,d})=\min_{\psi_{1}\in{\cal S}_{d}}\max_{\psi_{2}\in{\cal D}_{n,d}}|\langle\psi_{1}|\psi_{2}\rangle|$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{j=0,\ldots,n-1}\max_{\psi_{2}\in{\cal D}_{n,d-1}}\Big{|}\langle\psi_{1}|\psi_{2}\rangle\cos\theta\cos\frac{2\pi j}{n}+\sin\theta\sin\frac{2\pi j}{n}\Big{|}$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{j=0,\ldots,n-1}\max_{\psi_{2}\in{\cal D}_{n,d-1}}\Big{|}\langle\psi_{1}|\psi_{2}\rangle\Big{(}\cos\theta\cos\frac{2\pi j}{n}+\frac{1}{\langle\psi_{1}|\psi_{2}\rangle}\sin\theta\sin\frac{2\pi j}{n}\Big{)}\Big{|}$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|\Big{(}\max_{j=0,\ldots,n-1}\Big{|}\cos\theta\cos\frac{2\pi j}{n}+\frac{1}{\langle\psi_{1}|\psi_{2}\rangle}\sin\theta\sin\frac{2\pi j}{n}\Big{|}\Big{)}$
$\geq$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|$
$\cdot\Big{(}\max_{j=0,\ldots,n-1}\max\Big{(}\Big{|}\cos\theta\cos\frac{2\pi j}{n}+\sin\theta\sin\frac{2\pi j}{n}\Big{|},\Big{|}\cos\theta\cos\frac{2\pi j}{n}-\sin\theta\sin\frac{2\pi j}{n}\Big{|}\Big{)}\Big{)}$
$=$ $\min_{\theta}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|\Big{(}\max_{j=0,\ldots,n-1}\max\Big{(}\Big{|}\sin(\frac{2\pi j}{n}+\theta)\Big{|},\Big{|}\sin(\frac{2\pi j}{n}-\theta)\Big{|}\Big{)}\Big{)}$
$=$ $\Big{(}\min_{\theta}\max_{j=0,\ldots,n-1}\max\Big{(}\Big{|}\sin(\frac{2\pi j}{n}+\theta)\Big{|},\Big{|}\sin(\frac{2\pi j}{n}-\theta)\Big{|}\Big{)}\Big{(}\min_{\psi_{1}\in{\cal S}_{d-1}}\max_{\psi_{2}\in{\cal D}_{n,d-1}}|\langle\psi_{1}|\psi_{2}\rangle|\Big{)}$
$\geq$ $\cos\frac{\pi}{n}F({\cal D}_{n,d-1})=\cos^{d}\frac{\pi}{n}.$ (130)
A-B Complex discrete subset
Next, we construct the complex discrete subset ${\cal D}_{n,d}\subset\mathbb{C}^{d}$, where all elements of ${\cal D}_{n,d}$ are normalized vectors. Using ${\cal D}_{n,2d-2,R}$, we define ${\cal D}_{n,d}$ as
[…]
$x=(x_{1},\ldots,x_{2d-1})\in{\cal D}_{n,2d-2,R}\}.$ (131)
The number of elements $|{\cal D}_{n,d}|$ is $n^{2d-2}$.
Then, we have
$\frac{1}{4}\epsilon_{2}({\cal D}_{n,d})^{2}=1-\min_{\psi_{1}\in\mathbb{C}_{n}^{d}}\max_{\psi_{2}\in{\cal D}_{n,d}}|\langle\psi_{1}|\psi_{2}\rangle|^{2}$
$\leq$ $1-(\min_{\psi_{1}\in\mathbb{C}_{n}^{d}}\max_{\psi_{2}\in{\cal D}_{n,d}}\Re|\langle\psi_{1}|\psi_{2}\rangle|)^{2}$
$=$ $1-(\min_{\psi_{1}\in S_{2d-2}}\max_{\psi_{2}\in{\cal D}_{n,2d-2,R}}\langle\psi_{1}|\psi_{2}\rangle)^{2}$
$=$ $1-(\min_{\psi_{1}\in S_{2d-2}}\max_{\psi_{2}\in{\cal D}_{n,2d-2,R}}\langle\psi_{1}|\psi_{2}\rangle)^{2}$
$\leq$ $1-\cos^{2d}\frac{\pi}{n}=1-(1-\sin^{2}\frac{\pi}{n})^{d}\leq d\sin^{2}\frac{\pi}{n}\leq d\frac{\pi^{2}}{n^{2}}.$ (132)
When we choose $n$ as $4d\frac{\pi^{2}}{n^{2}}=\epsilon_{2}^{2}$, i.e., $n=\sqrt{\frac{4d\pi^{2}}{\epsilon_{2}^{2}}}$, the number of elements $|{\cal D}_{n,d}|$ is $(\frac{2d^{1/2}\pi}{\epsilon_{2}})^{d-1}$.
Appendix B Proof of Lemma 10
Since the relation (60) follows from (61) with taking the limit $\alpha\to 1$, we show only (61). Define the projection $P:=\sum_{j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|$ and the TP-CP map
$\Lambda^{*}(\rho^{\prime}):=\sum_{j}P\rho^{\prime}P+(I-P)\rho^{\prime}(I-P).$ (133)
We choose a product pure state $|v_{A}\rangle|v_{B}\rangle$. The relation $\|(I-P)|v_{A}\rangle|v_{B}\rangle\|=0$ holds if and only if the relations
|$\begin{split}&|\{j:|\langle u_{j,A}|v_{A}\rangle|\neq 0\}|=1,\\ &\{j:|\langle u_{j,A}|v_{A}\rangle|\neq 0\}=\{j:|\langle u_{j,B}|v_{B}\rangle|\neq 0\}\end{split}$ (134)
hold. When the above condition holds, the relation $P|v_{A}\rangle|v_{B}\rangle=|u_{j,A}\rangle|u_{j,B}\rangle$ holds for $j\in\{j:|\langle u_{j,A}|v_{A}\rangle|\neq 0\}\cap\{j:|\langle u_{j,B}|v_{B}\rangle|\neq 0\}$.
Also, we have $P|v_{A}\rangle|u_{j,B}\rangle=(\langle u_{j,A}|v_{A}\rangle)|u_{j,A}\rangle|u_{j,B}\rangle$. Hence, we have
$P(\sigma_{A}\otimes I_{B})P=\sum_{j}P(\sigma_{A}\otimes|u_{j,B}\rangle\langle u_{j,B}|)P$
$=$ $\sum_{j}\langle u_{j,A}|\sigma_{A}|u_{j,A}\rangle)|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|.$ (135)
Therefore,
$C_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}$
$=$ $\min_{p}D_{\alpha}\Big{(}\sum_{j}p_{j}|j\rangle\langle j|\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{j,B}|\Big{)}$
$\geq$ $E_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{j,B}|\Big{)}.$ (136)
Now, we fix $\alpha\in(0,1]$. For any separable state $\sigma_{AB}\in{\cal S}_{AB}$, we have
$\sigma_{AB}=\sum_{j}p_{j}\rho_{A,j}\otimes\rho_{B,j}\leq\sum_{j}p_{j}\rho_{A,j}\otimes I_{B}=\sigma_{A}\otimes I_{B},$ (137)
where $\sigma_{A}:=\operatorname{\mathrm{Tr}}_{B}\sigma_{AB}$. Hence, we have
$\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\leq\rho_{AB}^{\frac{1-\alpha}{2\alpha}}(\sigma_{A}\otimes I_{B})\rho_{AB}^{\frac{1-\alpha}{2\alpha}}.$ (138)
Since $x\mapsto x^{\alpha}$ is matrix monotone, we have
$\operatorname{\mathrm{Tr}}(\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\sigma_{AB}\rho_{AB}^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leq\operatorname{\mathrm{Tr}}(\rho_{AB}^{\frac{1-\alpha}{2\alpha}}(\sigma_{A}\otimes I_{B})\rho_{AB}^{\frac{1-\alpha}{2\alpha}})^{\alpha}.$ (139)
Thus,
$E_{\alpha}^{*}(\rho_{AB})\geq\min_{\sigma_{A}}D_{\alpha}(\sigma_{A}\otimes I_{B}\|\rho_{AB}).$ (140)
We have
$D_{\alpha}\Big{(}\sigma_{A}\otimes I_{B}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$\geq$ $D_{\alpha}\Big{(}\Lambda^{*}(\sigma_{A}\otimes I_{B})\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $D_{\alpha}\Big{(}P(\sigma_{A}\otimes I_{B})P\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $D_{\alpha}\Big{(}\sum_{j}\langle u_{j,A}|\sigma_{A}|u_{j,A}\rangle)|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|$
$\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $D_{\alpha}\Big{(}\sum_{j}\langle u_{j,A}|\sigma_{A}|u_{j,A}\rangle)|j\rangle\langle j|\Big{\|}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}.$ (141)
Taking the minimum for $\sigma_{A}$, we have
$\min_{\sigma_{A}}D_{\alpha}\Big{(}\sigma_{A}\otimes I_{B}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$\geq$ $C_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|j\rangle\langle i|\Big{)}.$ (142)
The combination of (136), (140), and (142) yields (61) with $\alpha\in(0,1]$.
Next, we fix $\alpha>1$. When a separable state $\sigma_{AB}\in{\cal S}_{AB}$ satisfies $(I-P)\sigma_{AB}(I-P)\neq 0$, $D_{\alpha}\Big{(}\sigma_{AB}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}=\infty$. Hence, we have
$E_{\alpha}^{*}\Big{(}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$=$ $\min\Big{\{}D_{\alpha}\Big{(}\sigma_{AB}\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}$
$\hskip 43.05542pt\Big{|}\sigma_{AB}\in{\cal S}_{AB},(I-P)\sigma_{AB}(I-P)=0\Big{\}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\min_{p}D_{\alpha}\Big{(}\sum_{j}p_{j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{j,A}|\langle u_{j,B}|$
$\hskip 43.05542pt\Big{\|}\sum_{i,j}\theta_{i,j}|u_{j,A}\rangle|u_{j,B}\rangle\langle u_{i,A}|\langle u_{i,B}|\Big{)}.$ (143)
The relation $(a)$ follows from the condition (134). The combination of (136) and (143) yields (61) with $\alpha>1$.
Appendix C Proof of Theorems 11 and 12
C-A Generalized Arimoto-Blahut algorithm
Extending Arimoto-Blahut algorithm [72, 73], the paper [28] proposed a general algorithm. Recently, this general algorithm was generalized to another form [74]. We consider a finite-dimensional Hilbert space ${\cal H}$ and focus on the set ${\cal S}({\cal H})$ of density matrices on ${\cal H}$. In this section, as a preparation, we consider a minimization problem over the set ${\cal S}({\cal H})$. Given a continuous function $\Omega$ from the set ${\cal S}({\cal H})$ to the set of Hermitian matrices on ${\cal H}$, $B({\cal H})$, we consider the minimization $\min_{\rho\in{\cal S}({\cal H})}{\cal G}(\rho)$;
${\cal G}(\rho):=\operatorname{\mathrm{Tr}}\rho\Omega[\rho].$ (144)
In the following, we discuss the calculation of the following two problems;
$\overline{{\cal G}}(a):=\min_{\rho\in{\cal S}({\cal H})}{\cal G}(\rho),\quad\rho_{*}:=\mathop{\rm argmin}\limits_{\rho\in{\cal S}({\cal H})}{\cal G}(\rho).$ (145)
For this aim, with $\gamma>0$, we define the density matrix ${\cal F}[\sigma]$ as ${\cal F}[\sigma]:=\frac{1}{\kappa[\sigma]}\exp(\log\sigma-\frac{1}{\gamma}\Omega[\sigma])$, where $\kappa[\sigma]$ is the normalization factor $\operatorname{\mathrm{Tr}}\exp(\log\sigma-\frac{1}{\gamma}\Omega[\sigma])$. The paper [28] proposed Algorithm 3.
Algorithm 3 Minimization of ${\cal G}(\rho)$
1:Choose the initial value $\rho^{(1)}\in\mathcal{M}$;
2:repeat
3:     Calculate $\rho^{(t+1)}:={\cal F}[\rho^{(t)}]$;
4:until convergence of ${\cal G}(\rho^{(t)})$.
For any two densities $\rho$ and $\sigma$, we define
$D_{\Omega}(\rho\|\sigma):=\operatorname{\mathrm{Tr}}\rho(\Omega[\rho]-\Omega[\sigma]).$ (146)
The paper [28, Theorem 3.3] showed the following theorem, which discusses the convergence to the global minimum and the convergence speed.
Theorem 15
When any two densities $\rho^{(t)}$ and $\rho^{(t+1)}$ in satisfy the condition
$D_{\Omega}(\rho^{(t+1)}\|\rho^{(t)})\leq\gamma D(\rho^{(t+1)}\|\rho^{(t)}),$ (147)
and a state $\rho_{*}$ satisfies
$0\leq D_{\Omega}(\rho_{*}\|\rho)$ (148)
with any state $\rho$, Algorithm 3 satisfies the condition
${\cal G}(\rho^{(t_{0}+1)})-{\cal G}(\rho_{*})\leq\frac{\gamma D(\rho_{*}\|\rho^{(1)})}{t_{0}}$ (149)
with any initial state $\rho^{(1)}$.
C-B Proof of Theorem 11
We choose a classical-quantum state $\rho_{AC}$ as
$\rho_{AC}:=\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes X_{A|c}.$ (150)
We choose the function $\Omega_{R,{\cal D}}$ as
$\Omega_{{\cal D}}[\rho_{AC}]$
::=:= $\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (151)
Then, we have
$\operatorname{\mathrm{Tr}}\rho_{AC}\Omega_{\cal D}[\rho_{AC}]$
$=$ $\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AC}(\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}})\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{A}\sum_{c\in{\cal C}}X_{A|c}\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\log\rho_{AB}\Big{)}$
$=$ $D\Big{(}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (152)
That is, we can apply Algorithm 3. In this case, when we set $\sigma_{AC}[\{X_{A|c}\}_{c}]$ as
$\sigma_{AC}[\{X_{A|c}\}_{c}]:=\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes X_{A|c},$ (153)
${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]$ is written as
${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]$
$=$ $\frac{1}{\kappa[\{X_{A|c}\}_{c}]}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\exp(X_{A|c}-V_{c}[\{X_{A|c^{\prime}}\}_{c^{\prime}}]).$ (154)
Algorithm 1 is the same as the special case of Algorithm 3 with the above choice and $\gamma=1$.
Also, we have
$D_{\Omega_{{\cal D}}}(\rho_{AC}\|\rho_{AC}^{\prime})$
$=$ $\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}(\sum_{c\in{\cal C}}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*})\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}$
$=$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}$ (155)
$\leq$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes$
$X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}$
$=$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}}^{\prime}\Big{)}$
$=$ $D(\rho_{AC}\|\rho_{AC}^{\prime}).$ (156)
Also, (155) guarantees that the relation $D_{\Omega_{R,{\cal D}}}(\rho_{AC}\|\rho_{AC}^{\prime})\geq 0$ holds. Hence, the condition of Theorem 15 holds. That is, Theorem 11 holds.
C-C Proof of Theorem 12
We choose the function $\Omega_{\alpha,{\cal D}}$ as
$\Omega_{\alpha,{\cal D}}[\rho_{AC}]$
::=:= $\mathop{\rm sgn}(\alpha-1)\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (157)
Then, we have
$\operatorname{\mathrm{Tr}}\rho_{AC}\Omega_{\alpha,\cal D}[\rho_{AC}]$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}_{AC}(\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}})\sum_{c\in{\cal C}}|c\rangle\langle c|$
$\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}(\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*})$
$\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}\Big{)}.$ (158)
That is, we can apply Algorithm 3. In this case, when we set $\sigma_{AC}[\{X_{A|c}\}_{c}]$ as (153), ${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]$ is written as $\frac{1}{\kappa[\{X_{A|c}\}_{c}]}\exp(X_{A|c}-\frac{\mathop{\rm sgn}(\alpha-1)}{\gamma}V_{\alpha,c}[\{X_{A|c^{\prime}}^{(t)}\}_{c^{\prime}}])$. Algorithm 2 is the same as the special case of Algorithm 3 with the above choice.
We denote the minimizer by $\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]$. For $\alpha\in[0,1]$, We choose $Z_{*}=\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$ and $Z=\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$. Since $\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]$ is the minimizer, we have
$\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}\geq\operatorname{\mathrm{Tr}}Z^{\alpha}.$ (159)
Since $Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}$ and $(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})$ are density matrices, we have
$\operatorname{\mathrm{Tr}}(Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}\geq 1.$ (160)
Thus, we have
$D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])$
$=$ $-\operatorname{\mathrm{Tr}}_{AC}\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$-\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $-\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$\Bigg{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
$-\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\Bigg{)}$
$=$ $\operatorname{\mathrm{Tr}}Z_{*}Z^{\alpha-1}-\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}$
$=$ $(\operatorname{\mathrm{Tr}}Z_{*}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})$
$+\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}(\operatorname{\mathrm{Tr}}(Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1)$
$\geq$ $0$ (161)
Hence, the condition (148) holds.
We choose $Z^{\prime}=\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{\prime}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$. Then, we have
$D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])$ (162)
$=$ $(\operatorname{\mathrm{Tr}}Z^{\prime}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})$ (163)
$+\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha}(\operatorname{\mathrm{Tr}}({Z^{\prime}}^{\alpha}/\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1).$ (164)
When $Z^{\prime}$ is close to $Z$, $\operatorname{\mathrm{Tr}}({Z^{\prime}}^{\alpha}/\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1=O(\|Z^{\prime}-Z\|^{2})$ and $(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}}=O(\|Z^{\prime}-Z\|)$. However, when ${Z^{\prime}}$ and $Z$ are close to $X_{*}$, the first order derivative of $\operatorname{\mathrm{Tr}}{Z}^{\alpha}$ is zero. Hence, $(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}}=O(\|Z^{\prime}-Z\|^{2})$. Since $D(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])=O(\|Z^{\prime}-Z\|^{2})$, the value $\frac{D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])}{D(\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]\|\sigma_{AC}[\{X_{A|c}\}_{c}])}$ takes a finite value when $\sigma_{AC}[\{X_{A|c}^{\prime}\}_{c}]$ and $\sigma_{AC}[\{X_{A|c}\}_{c}]$ are close to $\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]$. Since $\{\sigma_{AC}[\{X_{A|c}^{(t)}\}_{c}]\}$ converges to an element, $\frac{D_{\Omega_{\alpha,{\cal D}}}(\sigma_{AC}[\{X_{A|c}^{(t+1)}\}_{c}]\|\sigma_{AC}[\{X_{A|c}^{(t)}\}_{c}])}{D(\sigma_{AC}[\{X_{A|c}^{(t+1)}\}_{c}]\|\sigma_{AC}[\{X_{A|c}^{(t)}\}_{c}])}$ can be bounded by a finite value. That is, when $\gamma>0$ is taken to be a sufficiently large value, the condition (147) holds.
Appendix D Proof of Theorem 14
D-A Proof of (123)
Similar to the papers [75, 76, 22], we employ Schur duality of ${\mathcal{H}}^{\otimes n}$ as
${\mathcal{H}}^{\otimes n}=\bigoplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}\otimes{\cal V}_{\lambda},$ (165)
${\cal U}_{\lambda}$ expresses the irreducible space of $\mathop{\rm SU}(d)$ and ${\cal V}_{\lambda}$ expresses the irreducible space of the representation $\pi$ of the permutation group $\mathfrak{S}_{n}$. Then, for any state $\rho$, we have
$\rho^{\otimes n}=\bigoplus_{\lambda\in Y_{d}^{n}}\rho_{\lambda}\otimes\rho_{\lambda,mix},$ (166)
where $\rho_{\lambda,mix}$ is the completely mixed state on ${\cal V}_{\lambda}$. Since the system ${\cal V}_{\lambda}$ has no information, it is sufficient to handle the states $\oplus_{\lambda\in Y_{d}^{n}}\rho_{\lambda},\oplus_{\lambda\in Y_{d}^{n}}\sigma_{\lambda}$ on $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$. Here, we define the map $\kappa_{n}$ as
$\kappa_{n}(\rho^{\prime}):=\sum_{\lambda\in Y_{d}^{n}}\operatorname{\mathrm{Tr}}_{{\cal V}_{\lambda}}P({\cal U}_{\lambda}\otimes{\cal V}_{\lambda})\rho^{\prime}P({\cal U}_{\lambda}\otimes{\cal V}_{\lambda})$ (167)
for a general state on ${\mathcal{H}}^{\otimes n}$, where $P({\cal K})$ expresses the projection to the subspace ${\cal K}$. We choose a basis $\{|u_{\lambda,j}\rangle\}_{j=1}^{d_{\lambda}}$ of ${\cal U}_{\lambda}$ such that $|u_{\lambda,j}\rangle$ is an eigenvector of $\rho_{\lambda}$, where $d_{\lambda}:=\dim{\cal U}_{\lambda}$. We define the pinching map $\Gamma_{\rho,n}$ as
$\Gamma_{\rho,n}(\rho^{\prime})$
::=:= $\sum_{\lambda\in Y_{d}^{n},j}(|u_{\lambda,j}\rangle\langle u_{\lambda,j}|\otimes P({\cal V}_{\lambda}))\rho^{\prime}(|u_{\lambda,j}\rangle\langle u_{\lambda,j}|\otimes P({\cal V}_{\lambda}))$ (168)
for a general state $\rho^{\prime}$ on ${\mathcal{H}}^{\otimes n}$. Due to the above mentioned structure, states $\Gamma_{\rho,n}(\sigma^{\otimes n})$ and $\rho^{\otimes n}$ are commutative with each other. We denote the dimension of $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$ by $d_{n}$. Since $\dim{\cal U}_{\lambda}\leq(n+1)^{d(d-1)/2}$ and $|Y_{d}^{n}|\leq(n+1)^{d-1}$ [70, (6.16) and (6.18)], we have
$d_{n}\leq(n+1)^{(d+2)(d-1)/2}.$ (169)
We have the pinching inequality:
$\rho^{\prime}\leq d_{n}\Gamma_{\rho,n}(\rho^{\prime}).$ (170)
We choose
$T_{n}(\sigma,R):=\{\Gamma_{\rho,n}(\sigma^{\otimes n})\geq e^{nR}\rho^{\otimes n}\}.$ (171)
Given $\alpha\in[0,1]$, we choose $R_{n,\alpha,r}:=\frac{-\max_{\sigma\in S}\log\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}-nr}{n(1-\alpha)}$ so that $-nr=\max_{\sigma\in S}\log\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}+(1-\alpha)nR_{n,\alpha,r}$. Here, we address several projections that are commutative with each other. For two commutative projections $P_{1}$ and $P_{2}$, we define
$P_{1}\cup P_{2}:=I-(I-P_{1})(I-P_{2}).$ (172)
Since it is natural to define $(I-P_{1})\cap(I-P_{2}):=(I-P_{1})(I-P_{2})$, the above definition for $P_{1}\cup P_{2}$ is also natural. When $\{|u_{j}\rangle\}_{j}$ is a common eigenvector system of $P_{1}$ and $P_{2}$, they can be written as $P_{k}=\sum_{j\in J_{k}}|u_{j}\rangle\langle u_{j}|$. Then, we have $P_{1}\cup P_{2}=\sum_{j\in J_{1}\cup J_{2}}|u_{j}\rangle\langle u_{j}|$. Based in the definition (172), we define the operator $T_{n,\alpha,r}$ as
$T_{n,\alpha,r}:=\bigcup_{\sigma\in S}T_{n}(\sigma,R_{n,\alpha,r}).$ (173)
Although $T_{n}(\sigma,R_{n,\alpha,r})$ is an operator on ${\mathcal{H}}^{\otimes n}$, it can be written by using a projection operator $P_{\sigma,n,\lambda}$ on ${\cal U}_{\lambda}$ as
$T_{n}(\sigma,R_{n,\alpha,r})=\bigoplus_{\lambda\in Y_{d}^{n}}P_{\sigma,n,\lambda}\otimes P({\cal V}_{\lambda}).$ (174)
That is, we choose the projection operator $P_{\sigma,n,\lambda}$ as the above. Hence, we have
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}\Big{(}\bigcup_{\sigma\in S}P_{\sigma,n,\lambda}\Big{)}\otimes P({\cal V}_{\lambda}).$ (175)
Since all projections $P_{\sigma,n,\lambda}$ are diagonalized with the basis $\{|u_{\lambda,j}\rangle\}_{j=1}^{d_{\lambda}}$ of ${\cal U}_{\lambda}$, it is sufficient to clarify whether the range of $P_{\sigma,n,\lambda}$ contains $|u_{\lambda,j}\rangle$. If the range of $P_{\sigma,n,\lambda}$ contains $|u_{\lambda,j}\rangle$, there exists an element $\sigma_{\lambda,j}\in S$ such that $P_{\sigma_{\lambda,j},n,\lambda}$ contains $|u_{\lambda,j}\rangle$. If the range of $P_{\sigma,n,\lambda}$ does not contain $|u_{\lambda,j}\rangle$, no element $\sigma\in S$ satisfy the condition $P_{\sigma,n,\lambda}$ contains $|u_{\lambda,j}\rangle$. Therefore, we can choose $d_{\lambda}$ elements $\sigma_{\lambda,1},\ldots,\sigma_{\lambda,d_{\lambda}}$ in $S$ such that
$\bigcup_{\sigma\in S}P_{\sigma,n,\lambda}=\bigcup_{j\in[d_{\lambda}]}P_{\sigma_{\lambda,j},n,\lambda},$ (176)
where $[d_{\lambda}]:=\{1,\ldots,d_{\lambda}\}$. Here, the elements $\sigma_{\lambda,j}$ are not necessarily distinct, in general.
Therefore, we have
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}\Big{(}\bigcup_{j\in[d_{\lambda}]}P_{\sigma_{\lambda,j},n,\lambda}\Big{)}\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}\bigoplus_{\lambda\in Y_{d}^{n}}(P_{\sigma_{\lambda^{\prime},j},n,\lambda})\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r}).$ (177)
We have
$\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}\stackrel{{\scriptstyle(a)}}{{\leq}}\operatorname{\mathrm{Tr}}\rho^{\otimes n}\sum_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$=$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
=$\stackrel{{\scriptstyle(d)}}{{=}}$ $d_{n}\big{(}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}\big{)}$
$\cdot\big{(}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{nr}\big{)}^{\frac{\alpha}{1-\alpha}}$
$=$ $d_{n}(\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha})^{\frac{1}{1-\alpha}}e^{n\frac{\alpha r}{1-\alpha}}.$ (178)
Each step can be shown as follows. Step $(a)$ follows from (177). Step $(b)$ can be shown as follows. Since the condition $\Gamma_{\rho,n}(\sigma^{\otimes n})\geq e^{nR}\rho^{\otimes n}$ implies $\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}\geq e^{\alpha nR}(\rho^{\otimes n})^{\alpha}$, the definition (171) guarantees that the inequality
$T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})\rho^{\otimes n}$
$=$ $T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})(\rho^{\otimes n})^{1-\alpha}(\rho^{\otimes n})^{\alpha}$
$\leq$ $\operatorname{\mathrm{Tr}}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$=$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}.$
Step $(c)$ follows from the relation $\sum_{\lambda\in Y_{d}^{n}}d_{\lambda}=\dim\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}=d_{n}$. Step $(d)$ follows from the definition of $R_{n,\alpha,r}$.
For $\sigma\in S$, we have
$\operatorname{\mathrm{Tr}}\sigma^{\otimes n}(I-T_{n,\alpha,r})=\operatorname{\mathrm{Tr}}\sigma^{\otimes n}\Gamma_{\rho,n}(I-T_{n,\alpha,r})$
$=$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})(I-T_{n,\alpha,r})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})(I-T_{n}(\sigma,R_{n,\alpha,r}))$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{n(1-\alpha)R_{n,\alpha,r}}$
$\leq$ $\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{n(1-\alpha)R_{n,\alpha,r}}=e^{-nr}.$ (179)
Each step can be shown as follows. Step $(a)$ follows from the relation $T_{n}(\sigma,R_{n,\alpha,r})\leq T_{n,\alpha,r}$, which is shown by the relation (175). Step $(b)$ follows from (171) in the same way as Step $(b)$ of (178). Thus, we have
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)\geq-\frac{1}{n}\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}.$ (180)
In fact, applying the matrix monotone function to (170), we have
$d_{n}^{\alpha-1}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\rho,n}(\sigma^{\otimes n})(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $\Big{(}d_{n}\Gamma_{\rho,n}\Big{(}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
$\leq$ $\Big{(}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}.$ (181)
Thus, we have
$\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}$
$=$ $\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\rho,n}(\sigma^{\otimes n})(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha}$
$=$ $\operatorname{\mathrm{Tr}}\big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\big{)}$
$\cdot\Big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
$=$ $\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot\Big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $d_{n}^{1-\alpha}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (182)
where Step $(a)$ follows from (181).
Combining (178) and (182), we have
$\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}\leq d_{n}^{2}(\max_{\sigma\in S}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{\frac{n}{1-\alpha}}e^{n\frac{\alpha r}{1-\alpha}}$
$=$ $d_{n}^{2}\exp\Big{(}n\frac{-(1-\alpha)E_{\alpha}^{*}(\rho)+\alpha r}{1-\alpha}\Big{)}.$ (183)
Since an arbitrary number $\alpha\in[0,1]$ satisfies (180), (183) yields
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$
$\geq$ $-\frac{1}{n}\log d_{n}^{2}+\max_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (184)
Hence, using (169), we obtain (125) and (123).
Remark 4
Although the above presented proof employs several techniques that come from prior works, this proof is not simple combination of techniques that come from prior works. Our test is given in (175), and it has not been introduced in existing references [21, 22]. In fact, the method by [21, 22] cannot be applied to the case when $H_{0}$ hypothesis is the set of all separable states over $n$ bipartite systems, as opposed to an IID $n$-fold tensor product of a separable state. To extend quantum Sanov theorem to this case, our obtained evaluation is essentially needed. To upper bound the error probability, we have invented a bounding method different from the existing bounding method for the simple hypothesis [78] because the existing bounding method does not work in this case and it is related to Petz relative Rényi entropy. In fact, the above evaluation requires new careful handling the parameter $\alpha$ due to the matrix convexity.
D-B Proof of (124)
Here, we employ the idea invented in [77], whose detail is explained in [12, Section 3.8]. We choose a test $T_{n}$ such that
$\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n}\geq e^{-nr}$ (185)
for $\sigma\in S$. For $\alpha\geq 1$ and $\sigma\in S$, the information processing inequality implies
$D_{\alpha}(\sigma^{\otimes n}\|\rho^{\otimes n})$
$\geq$ $\frac{1}{\alpha-1}\log\Big{(}(\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n})^{\alpha}(\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n})^{1-\alpha}$
$+(\operatorname{\mathrm{Tr}}\sigma^{\otimes n}(I-T_{n}))^{\alpha}(\operatorname{\mathrm{Tr}}\rho^{\otimes n}(I-T_{n}))^{1-\alpha}\Big{)}$
$\geq$ $\frac{1}{\alpha-1}\log\Big{(}(\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n})^{\alpha}(\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n})^{1-\alpha}\Big{)}$
$=$ $\frac{\alpha}{\alpha-1}\log\operatorname{\mathrm{Tr}}\sigma^{\otimes n}T_{n}-\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}$
$\geq$ $-\frac{\alpha}{\alpha-1}nr-\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}.$ (186)
This inequality is converted to
$-\frac{1}{n}\log\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}\leq\frac{1}{n}D_{\alpha}(\sigma^{\otimes n}\|\rho^{\otimes n})+\frac{\alpha}{\alpha-1}r.$ (187)
Since any $T_{n}$ satisfying (185) satisfies (187), we have
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)\leq\frac{1}{n}D_{\alpha}(\sigma^{\otimes n}\|\rho^{\otimes n})+\frac{\alpha r}{\alpha-1}$
$=$ $D_{\alpha}(\sigma\|\rho)+\frac{\alpha r}{\alpha-1}.$ (188)
Since any $\sigma\in S$ satisfies (188), we have
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)\leq\min_{\sigma\in S}D_{\alpha}(\sigma\|\rho)+\frac{\alpha r}{\alpha-1}$
$=$ $E_{\alpha}^{*}(\rho)+\frac{\alpha r}{\alpha-1},$ (189)
which implies (124).
Appendix E Proof of Theorem 2
E-A Proof of (14)
We employ notations and results in Appendix D when ${\mathcal{H}}={\mathcal{H}}_{A}\otimes{\mathcal{H}}_{B}$ and $\rho=\rho_{AB}$. We define $\Gamma_{\rho_{AB},n}$ and $d_{n}$ in the same way.
We define the map ${\cal T}_{n}$ as ${\cal T}_{n}(\sigma):=\sum_{g\in\mathfrak{S}_{n}}\frac{1}{|\mathfrak{S}_{n}|}\pi(g)\sigma\pi(g)^{\dagger}$ for a state on ${\mathcal{H}}^{\otimes n}$. We define the set of permutation-invariant separable states as
${\cal S}_{AB}^{n,sym}:=\{\sigma\in{\cal S}_{AB}^{n}|\pi(g)\sigma\pi(g)^{\dagger}=\sigma,\quad g\in\mathfrak{S}_{n}\}.$ (190)
Then, we have the following lemma.
Lemma 16
$\overline{\beta}_{\epsilon,n}(\rho_{AB})$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (191)
Proof: Since ${\cal S}_{AB}^{n,sym}\subset{\cal S}_{AB}^{n}$, we have
$\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}$
$\geq$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (192)
Hence, it is sufficient to prove the opposite inequality.
We choose the test $T_{\epsilon}$ such that any element $\sigma_{AB}^{\prime}\in{\cal S}_{AB}^{n,sym}$ satisfies
$\operatorname{\mathrm{Tr}}(I-T_{\epsilon})\sigma_{AB}^{\prime}\leq\epsilon,$ (193)
$\operatorname{\mathrm{Tr}}T_{\epsilon}\rho_{AB}^{\otimes n}$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (194)
Due to the symmetric property, any element $\sigma_{AB}^{\prime}\in{\cal S}_{AB}^{n,sym}$ satisfies
$\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\sigma_{AB}^{\prime}\leq\epsilon,$ (195)
$\operatorname{\mathrm{Tr}}{\cal T}_{n}(T_{\epsilon})\rho_{AB}^{\otimes n}$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (196)
For any element $\sigma_{AB}\in{\cal S}_{AB}^{n}$, we have $\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\sigma_{AB}=\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\pi(g)\sigma_{AB}\pi(g)^{\dagger}$. Since ${\cal T}_{n}(\sigma_{AB})\in{\cal S}_{AB}^{n,sym}$, we have $\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon}))\sigma_{AB}=\operatorname{\mathrm{Tr}}(I-{\cal T}_{n}(T_{\epsilon})){\cal T}_{n}(\sigma_{AB})\leq\epsilon$. Thus,
$\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}$
$\leq$ $\operatorname{\mathrm{Tr}}{\cal T}_{n}(T_{\epsilon})\rho_{AB}^{\otimes n}=\operatorname{\mathrm{Tr}}T_{\epsilon}\rho_{AB}^{\otimes n}$
$=$ $\min_{0\leq T\leq I}\Big{\{}\operatorname{\mathrm{Tr}}T\rho_{AB}^{\otimes n}\Big{|}\max_{\sigma_{AB}\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}(I-T)\sigma_{AB}\leq\epsilon\Big{\}}.$ (197)
Since any element $\sigma\in{\cal S}_{AB}^{n,sym}$ is written as the form
$\sigma=\bigoplus_{\lambda\in Y_{d}^{n}}\rho_{\lambda}\otimes\rho_{\lambda,mix}$ (198)
in the same way as (166), replacing the set $\{\sigma^{\otimes n}|\sigma\in S\}$ of states by ${\cal S}_{AB}^{n,sym}$, given $r>0$, we define $T_{n,R}$ and $R_{n,\alpha,r}$. In the same way as (178), we have
$\operatorname{\mathrm{Tr}}\rho_{AB}^{\otimes n}T_{n,\alpha,r}$
$\leq$ $d_{n}(\max_{\sigma\in{\cal S}_{AB}^{n,sym}}\operatorname{\mathrm{Tr}}\Gamma_{\rho_{AB},n}(\sigma)^{\alpha}(\rho_{AB}^{\otimes n})^{1-\alpha})^{\frac{1}{1-\alpha}}e^{n\frac{\alpha r}{1-\alpha}}$ (199)
for $\alpha\in[0,1]$. For $\sigma\in{\cal S}_{AB}^{n,sym}$, in the same way as (179), we have
$\operatorname{\mathrm{Tr}}\sigma(I-T_{n,\alpha,r})\leq e^{-nr}.$ (200)
Thus, we have
$-\frac{1}{n}\log\overline{\beta}_{e^{-nr},n}(\rho_{AB})\geq-\frac{1}{n}\log\operatorname{\mathrm{Tr}}\rho_{AB}^{\otimes n}T_{n,\alpha,r}.$ (201)
In the same way as (182), we have
$\operatorname{\mathrm{Tr}}\Gamma_{\rho_{AB},n}(\sigma)^{\alpha}(\rho_{AB}^{\otimes n})^{1-\alpha}$
$\leq$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho_{AB}^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma(\rho_{AB}^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}.$ (202)
Combining (199) and (202), we have
$\operatorname{\mathrm{Tr}}\rho_{AB}^{\otimes n}T_{n,\alpha,r}\leq d_{n}^{2}\exp\Big{(}\frac{-(1-\alpha){E}_{\alpha}^{*}(\rho_{AB}^{\otimes n})+n\alpha r}{1-\alpha}\Big{)}$ (203)
in the same way as (183).
Since an arbitrary real $\alpha\in[0,1]$ satisfies (180), (203) yields
$-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho_{AB})$
$\geq$ $-\frac{1}{n}\log d_{n}^{2}+\max_{0\leq\alpha\leq 1}\frac{\frac{1}{n}(1-\alpha)E_{\alpha}^{*}(\rho_{AB}^{\otimes n})-\alpha r}{1-\alpha},$ (204)
using (169), we obtain (14) and (16) by taking the limit.
E-B Proof of (15)
For $\alpha\geq 1$ and $\sigma\in{\cal S}_{AB}^{n}$, in the same way as (188), we have
$-\frac{1}{n}\log\overline{\beta}_{1-e^{-nr},n}(\rho)\leq\frac{1}{n}D_{\alpha}(\sigma\|\rho^{\otimes n})+\frac{\alpha r}{\alpha-1}$ (205)
for $\sigma\in{\cal S}_{AB}^{n}$. Since any $\sigma\in{\cal S}_{AB}^{n}$ satisfies (205), we have
$-\frac{1}{n}\log\overline{\beta}_{1-e^{-nr},n}(\rho)\leq$ $\min_{\sigma\in{\cal S}_{AB}^{n}}\frac{1}{n}D_{\alpha}(\sigma\|\rho^{\otimes n})+\frac{\alpha r}{\alpha-1}$
$=$ $\frac{1}{n}E_{\alpha}^{*}(\rho^{\otimes n})+\frac{\alpha r}{\alpha-1},$ (206)
which implies (17) and (15) by taking the limit.